### 05 Apr 2026 v1

# April 2026

Abstract
This is the second and final part of a work initiated in [Mor26]. We construct a class of positive-energy relativistic spatial localization observables in Minkowski spacetime within the standard framework of quantum field theory, based on the stress–energy–momentum tensor smeared with suitable test functions. For every fixed timelike direction, the construction yields a family of positive operator-valued measures (POVMs) defined on spacelike hypersurfaces, which are well defined on each $n$-particle sector and satisfy a natural relativistic causality condition ruling out superluminal propagation of detection probabilities. The proposed localization observables arise from local or quasi-local quantum field theoretic quantities, thereby providing a rigorous realization of previously heuristic constructions. In the one-particle sector, the scheme reduces to the observable introduced in [Mor23], and its first moment reproduces the Newton–Wigner position operator under suitable normalization conditions. Since the normally ordered stress–energy–momentum tensor fails to be positive on the full Fock space due to the Reeh–Schlieder theorem, we analyze the role of quantum energy inequalities and establish lower bounds that allow us to control deviations from positivity. This leads to the introduction of regularized families of bounded from below operators approximating the localization effects. We further construct conditional localization observables associated with finite laboratories by means of suitably modified local energy operators and their Friedrichs self-adjoint extensions. In particular, thanks to Haag duality and a result by Kadison on the affiliation of Friedrichs extensions with von Neumann algebras, the resulting conditional POVMs are shown to belong to local von Neumann algebras, and thus to commute when associated with causally separated regions, in agreement with the Araki–Haag–Kastler framework. Our results provide a quantum field theoretic implementation of the idea that commutativity of localization observables should be recovered at the level of conditional measurements in spacetime regions of finite extent.
1 Introduction
1.1 Localization, Causality, Commutativity, and QFT
In [Mor26], we addressed the issue of commutativity of effects for POVMs describing the spatial localization of a quantum system in the rest space of a reference frame in Minkowski spacetime. The problem stems from the analysis of spatial localization by Halvorson and Clifton [HaCl02], who showed that natural assumptions like positive energy, additivity, etc. underlying any notion of localization are incompatible with the commutativity of the effects $A(\Delta)$, $A(\Delta^{\prime})$ associated with spacelike separated regions $\Delta$, $\Delta^{\prime}$, thus appearing to be in tension with the description of microcausality in the Araki–Haag–Kastler (AHK) framework of local quantum physics: $A(\Delta),A(\Delta^{\prime})]\neq 0$ in general. Without entering into the details of specific localization observables, and drawing on an analysis of relativistic causality largely inspired by Busch, we showed in [Mor26] that the requirement of commutativity is in fact not justified for observables describing the position of a relativistic quantum particle. Here, a particle is understood as a quantum system endowed with the propensity to assume a definite position when subjected to a complete detection procedure in which the entire rest space $\Sigma$ of an observer—more generally, a Cauchy surface—is filled with detectors. These detectors are represented by a POVM associated with the Borel sets $\Delta\in\mathscr{B}(\Sigma)$ of the space: $\mathscr{B}(\Sigma)\ni\Delta\mapsto A(\Delta)$. In this setting, there is no compelling reason to require commutativity of the effects associated with causally separated spatial regions because standard arguments based on no-signaling or relativistic consistence are not triggered. In that analysis, we worked within a framework more basic than the AHK one, without assuming that physically localized sets of observables carry any a priori algebraic structure. From the AHK perspective, the effects of the POVMs under consideration are therefore not elements of local operator algebras. When the AHK formalism is assumed, this conclusion can also be justified a posteriori, in particular in light of certain consequences of the Reeh–Schlieder theorem when one imposes that the localization probability of the vacuum state is zero.
On the other hand, in [Mor26] we suggested that more realistic experimental situations can be considered, in which detectors are switched on only within a finite-size laboratory, and one considers conditional localization POVMs $\mathscr{B}(\Delta_{0})\ni\Delta\mapsto B_{\Delta_{0}}(\Delta)$. That is, they measure the probability of detecting a particle in a spatial subregion $\Delta$, given that it is detected somewhere in the finite rest space of the laboratory $\Delta_{0}\supset\Delta$. In these cases [Mor26], commutativity could in principle be recovered in the following sense: two effects $B_{\Delta_{0}}(\Delta)$ and $B_{\Delta^{\prime}_{0}}(\Delta^{\prime})$ associated with regions in causally separated laboratories with respective rest-spaces $\Delta_{0}$, $\Delta^{\prime}_{0}$ may commute:
$[B_{\Delta_{0}}(\Delta),B_{\Delta^{\prime}_{0}}(\Delta^{\prime})]=0\>.$ (1)
More precisely, they could be represented by elements of local operator algebras $B_{\Delta_{0}}(\Delta)\in{\mathfrak{W}}({\cal O})$, $B_{\Delta^{\prime}_{0}}(\Delta^{\prime})\in{\mathfrak{W}}({\cal O}^{\prime})$ with $\Delta\subset{\cal O}$, $\Delta^{\prime}\subset{\cal O}^{\prime}$, in accordance with the AHK framework.
Assume that a localization observable is given in terms of effects $A(\Delta)$ with $\Delta$ any Borel region of any complete rest space $\Sigma$, and the bounded $\Delta_{0}$ defines the spatially finite rest space of a laboratory. Taking advantage of the so-called gentle measurement lemma, it was suggested in [Mor26] that the POVM normalized on $\Delta_{0}$
$B_{\Delta_{0}}(\Delta)=V\frac{1}{\sqrt{A(\Delta_{0})}}A(\Delta)\frac{1}{\sqrt{A(\Delta_{0})}}V^{\dagger}\>,\quad\Delta\subset\Delta_{0}$ (2)
where $V$ is a given unitary operator, has the properties of a conditional localization POVM for states whose probability of finding the system in $\Delta_{0}$ (measured with the effect $VA(\Delta_{0})V^{\dagger}$) is close to $1$. This intepretation makes sense in the general case where the operators $A(\Delta)$ are positive and define a positive-operator valued measure on $\Sigma$, but are not necessarily effects, i.e., bounded by $I$.
The present paper has two main aims. First, we show that positive-energy localization observables can be constructed within the standard formalism of QFT, in particular by exploiting the stress–energy–momentum tensor operator $\thinspace\hat{T}_{\mu\nu}\thinspace:[f]=\int_{{\mathbb{M}}}:\thinspace\hat{T}_{\mu\nu}(x)f(x)d^{4}x$ smeared with test functions $f\in\mathscr{D}({\mathbb{M}})$. This idea is consistent with a previous result in which a positive-energy localization observable was constructed on the one-particle space of a real scalar quantum field by restricting the stress–energy tensor to that space [Mor23], thereby placing on rigorous mathematical grounds, and extending, an idea originally proposed by Terno [Ter14]. That construction was also shown [Mor23, DRM24] to satisfy a basic causality requirement due to Castrigiano [Cas17] who extended and generalized the causality constraints studied by Hegerfeldt in his celebrated works.
For future convenience, observe that, if ${\mathbb{R}}^{4}\ni x\mapsto U_{x}$ is the unitary representation of the translation group of ${\mathbb{M}}^{4}$, then $U_{x}:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]U_{x}^{\dagger}=:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]$ where $f_{x}(y):=f(y-x)$. Concerning the generalization developed in this work, we consider an $n$-particle state $\Psi\in{{\cal H}}^{(n)}$ with $n>0$. For every Borel set $\Delta\subset\Sigma$, where $\Sigma$ is a spacelike $3$-plane of Minkowski spacetime $\mathbb{M}$}, the effects $A_{f}^{u}(\Delta)$ of the notion of spatial localization we construct satisfy
$\langle\Psi|A_{f}^{u}(\Delta)\Psi\rangle=\lim_{\epsilon\to 0^{+}}\left\langle\Psi\left|\frac{1}{\sqrt{H^{u}+\epsilon I}}\int_{\Delta}:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}^{2}]\>u^{\mu}u^{\nu}_{\Sigma}\>d\Sigma(x)\frac{1}{\sqrt{H^{u}+\epsilon I}}\right.\Psi\right\rangle$
for a choice of a unit timelike vector $u$ defining the frame of the detectors, and $f\in\mathscr{D}_{R}({\mathbb{M}})$ such that $\int_{\mathbb{M}}f^{2}d^{4}x=1$. Here, $H^{u}$ is the Hamiltonian operator in the $u$-direction, and the regularization is necessary since the Minkowski vacuum $\Omega$ lies in the kernel of this operator. In the identity above, the effect $A_{f}^{u}(\Delta)$ on the left-hand side is defined independently of the integral operator on the right-hand side, and the identity is valid when considering expectation values: it is generally false for generic off-diagonal matrix elements.
The family $A^{u}_{f}(\Delta)$ is a well-behaved positive-energy localization observable in each space ${{\cal H}}^{(n)}$, in particular for one-particle states. Notably, the integral for $\Delta=\Sigma$ is normalized independently of the choice of $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$, provided $\int_{\mathbb{M}}f^{2}d^{4}x=1$. The causality requirement denoted by CC and already established in [Mor23, DRM24] also holds for the localization observables constructed here. We further analyze the physical meaning of the notion of localization introduced above, showing that it is associated with the center of $u$-energy of the field and that, for one-particle states, it reduces to the observable already introduced in [Mor23]. In the one-particle case, one can say even more: if one takes $f(x^{0},\vec{x})=h^{\prime}(x^{0})h(\vec{x})$, then, in the non-relativistic large-mass limit, $h$ determines the precision of a von Neumann measurement scheme for detecting a single particle, whereas $h^{\prime}$ can be chosen independently.
A problem with the family of operators $\frac{1}{\sqrt{H^{u}+\epsilon I}}\int_{\Delta}:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}^{2}]\>u^{\mu}u^{\nu}_{\Sigma}\>d\Sigma(x)\frac{1}{\sqrt{H^{u}+\epsilon I}}$ is that they do not define effects on the whole Fock space ${\mathfrak{F}}_{s}({{\cal H}}^{(1)})$, since the positivity condition $\thinspace\hat{T}_{\mu\nu}\thinspace:[f]u_{\mu}v_{\nu}\geq 0$, for $u,v$ timelike and future-directed, fails as a consequence of the Reeh-Schlieder theorem (though they are bounded from below). This issue will be carefully analyzed from the perspective of quantum energy inequalities. Relying on known general results by Fewster and collaborators [Few05, FeSm08, Few12], we prove in particular that, given $u\in{\mathsf{T}}_{+}$, for every $\eta>0$, it is possible to modify the temporal part $h^{\prime}$ by enlarging its support in the smearing function $f$ in such a way that there is a constant $c^{u}_{\eta,f}>0$ such that $\left(:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}^{2}]-\eta g_{\mu\nu}I\right)u^{\mu}v^{\nu}\geq c^{u}_{\eta,f}|u\cdot v|I$ for every $v\in{\mathsf{T}}_{+}$ and every $x\in{\mathbb{M}}$. Also note that $J_{\mu}(x):=-\left(:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}^{2}]-\eta g_{\mu\nu}I\right)u^{\mu}$ is a conserved, causal, and future-directed (where it does not vanish) current. (The sign - also in front of $\eta$ is due to the fact that we are adopting the signature ,+,+,+-,+,+,+.)
Based on this result we pass to the second main goal of this work. The paper aims to construct conditional localization observables using a slightly modified (in order to remove negative energies) local energy operator, as indicated above. Focusing on local-energy symmetric operators
${\mathsf{H}}^{u}_{f,\eta}(\Delta):=\int_{\Delta}(:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}^{2}]-\eta g_{\mu\nu}I)u^{\mu}u^{\nu}_{\Sigma}\>d\Sigma(x)$
we consider their Friedrichs selfadjoint extensions $\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta)$, since no essential selfadjointness result is known in the literature to the author. These selfadjoint extensions allow the use of functional calculus, leading to a physically meaningful definition of local conditional POVMs of this type (the bar denoting the closure)
$B_{\Delta_{0}}(\Delta)=\overline{\frac{1}{\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}{\mathsf{H}}^{u}_{f,\eta}(\Delta)\frac{1}{\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}}\>,\quad\Delta\subset\Delta_{0}\>.$
At this juncture, Haag duality, together with a careful analysis of the aforementioned Friedrichs extensions, and a result of Kadison concerning the affiliation of such extensions with von Neumann algebras [Kad89], eventually yields the desired commutativity result (1) when $\Delta_{0}$ and $\Delta_{0}^{\prime}$ are strongly causally separated. More precisely, we prove that the operators $B_{\Delta_{0}}(\Delta)$ belong to local von Neumann algebras ${\mathfrak{W}}({\cal O})$ obtained from local Weyl algebras, in agreement with the AHK perspective.
We also prove a further result. Defining positive operators
${\mathsf{A}}_{f,\epsilon,\eta}^{u}(\Delta):=\frac{1}{\sqrt{H^{u}+\epsilon I}}\int_{\Delta}(:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}^{2}]-\eta g_{\mu\nu}I)u^{\mu}u^{\nu}_{\Sigma}\>d\Sigma(x)\frac{1}{\sqrt{H^{u}+\epsilon I}}$
they allow one to approximate the localization effects $A^{u}_{f}(\Delta)$ with arbitrary precision in a laboratory based on the finite rest space $\Delta_{0}$, since
$|tr(\rho A^{u}_{f}(\Delta))-\lim_{\epsilon\to 0^{+}}tr(\rho{\mathsf{A}}^{u}_{f,\epsilon,\eta}(\Delta))|\leq\eta\frac{|u\cdot u_{\Sigma}||\Delta_{0}|}{m}$
for every $n$-particle state $\rho$ with $n>0$, where $m$ is the mass of the particle, and $\Delta\subset\Delta_{0}$ measurable. We establish the following relation, in agreement with (2),
$B_{\Delta_{0}}(\Delta)^{u}_{f,\eta}=V^{u}_{f,\epsilon,\eta,\Delta_{0}}\frac{1}{\sqrt{{\mathsf{A}}^{u}_{f,\epsilon,\eta}(\Delta_{0})}}{\mathsf{A}}^{u}_{f,\epsilon,\eta}(\Delta)\frac{1}{\sqrt{{\mathsf{A}}^{u}_{f,\epsilon,\eta}(\Delta_{0})}}V^{u\dagger}_{f,\epsilon,\eta,\Delta_{0}}\>,$
valid for some unitaties $V^{u}_{f,\epsilon,\eta,\Delta_{0}}$.
In the recent literature, there have been other analyses aimed at reconciling the notion of locality in QFT with the concept of spatial localization for relativistic quantum systems. We mention in particular two works written in a more physics-oriented style. One of them is the comparative study [FaCo24], where various aspects of the apparent violation of locality and causality are analyzed. The other is [Tu26], where $(1+1)$-dimensional QFT is considered for both bosons and fermions, making use of the reduced density matrix formalism. In particular, it is shown there that scalar particles cannot be localized within any compact region.
The style of this work is deliberately elementary, pedagogical, and, as far as possible, self-contained with respect to free QFT in Minkowski spacetime. The relevant notions, including in particular some basic foundational aspects of the Araki–Haag–Kastler approach likethe Reeh-Schleider property and the Haag duality, are introduced progressively. It is nevertheless assumed that the reader is familiar with the mathematical notions of *-algebra, $C^{*}$-algebra, von Neumann algebra, and general spectral theory. The broader goal is to bring the community working on quantum measurement theory closer to the community working on local quantum field theory.
1.2 Structure of this work
After a subsection devoted to listing the fundamental notions and notation used in this work, Section 2 provides a brief review of the notions introduced in [Mor26] regarding localization observables in terms of POVMs and conditional localization observables. Section 3 introduces the basic notions of QFT in Minkowski spacetime for a free real massive scalar quantum field. Section 4 presents several crucial concepts from mathematical physics, such as the smeared normally ordered stress-energy tensor operator and its fundamental properties concerning locality and positivity. Section 5 is devoted to the construction of a relativistic localization observable from the stress-energy operator and to the analysis of its fundamental mathematical and physical properties. The final section 6, before the conclusions stated in Section 7, focuses on conditional localization POVMs arising from a local notion of the energy operator. In particular, we prove that these POVMs belong to local von Neumann algebras, as expected in the AHK framework, and in particular that they commute when associated with causally separated laboratories. Furthermore, we establish a relation between these POVMs and the relativistic localization observables defined on the whole spacetime introduced in the previous section. The appendix section contains several technical proofs of intermediate statements appearing in the main text.
1.3 Notions and notations
The reader may initially skip this section and come back to it later when necessary.
We assume $c=1$, $\hbar=1$ throughout the rest of this paper, and the notation $A\subset B$ allows the case $A=B$. The Hilbert spaces we consider are complex and symmetric operators are densely defined by definition.
A. Minkowski spacetime. A four-dimensional real affine space whose space of translations $\mathsf{V}$ is equipped with a bilinear, non-degenerate, symmetric form $g$ with signature (,+,+,+)(-,+,+,+) is the Minkowski spacetime $\mathbb{M}$}. The points of $\mathbb{M}$} are called events and $g$ is called the Minkowski metric. We shall make use of the notation $u\cdot v:=g(u,v)$ if $u,v\in\mathsf{V}$. We also use the dot to indicate the standard (positive) scalar product of $3$-vectors $\vec{u}\cdot\vec{v}$, viewing them as spacelike vectors (see below). Upon choosing an origin $o\in{\mathbb{M}}$, the points $p\in{\mathbb{M}}$ are in one-to-one correspondence with vectors of $\mathsf{V}$ through the map ${\mathbb{M}}\ni p\mapsto p-o\in\mathsf{V}$. We shall take advantage of this identification several times in the rest of this paper. If $p\in{\mathbb{M}}$ and $v\in\mathsf{V}$, $q=p+v$ means that qp=vq-p=v.
A vector $v\in\mathsf{V}$ is spacelike if $g(v,v)>0$ or $v=0$. It is causal if $g(v,v)\leq 0$ and $v\neq 0$. A causal vector $v$ is timelike if $g(v,v)<0$, or lightlike if $g(v,v)=0$. Smooth curves are classified analogously according to their tangent vectors.
A set $\Lambda\subset{\mathbb{M}}$ is achronal if pqp-q cannot be timelike for $p,q\in\Lambda$. A maximal achronal set is an achronal set that is not a proper subset of another achronal set. $\Lambda\subset{\mathbb{M}}$ is spacelike if pqp-q is spacelike for $p,q\in\Lambda$.
The set of timelike vectors is an open cone made up of two disjoint open connected halves. A choice of one of them $\mathsf{V}_{+}$ defines a time orientation of $\mathbb{M}$}. The latter is henceforth assumed to be time oriented: $\mathsf{V}_{+}\subset\mathsf{V}$ is the open cone of future-directed timelike vectors. $\overline{\mathsf{V}_{+}}\setminus\{0\}$ is the cone of future-directed causal vectors. Notice that if $v\in\overline{\mathsf{V}_{+}}$ and $u$ is causal, then $u\in\overline{\mathsf{V}_{+}}$ if and only if $g(u,v)\leq 0$. We finally define the set of unit timelike future-directed vectors ${\mathsf{T}}_{+}:=\{u\in\mathsf{V}_{+}\>|\>g(u,u)=-1\}$.
If $A\subset{\mathbb{M}}$, the causal future $J^{+}(A):=\{p\in{\mathbb{M}}\>|\>p-q\in\overline{\mathsf{V}_{+}}\quad\mbox{for some $q\in A$}\}$ represents the events of $\mathbb{M}$} in the future of $A$ that can be physically influenced by $A$. The causal past $J^{-}(A):=\{p\in{\mathbb{M}}\>|\>q-p\in\overline{\mathsf{V}_{+}}\quad\mbox{for some $q\in A$}\}$ is defined symmetrically.
$A,B\subset{\mathbb{M}}$ are said to be causally separated if $(J^{+}(A)\cup J^{-}(A))\cap B=\varnothing$ (which is equivalent to $(J^{+}(B)\cup J^{-}(B))\cap A=\varnothing$).
If $R\subset{\mathbb{M}}$, its causal complement and causal completion are, respectively,
$R^{\perp\!\!\!\!\perp}:={\mathbb{M}}\setminus(J^{+}(R)\cup J^{-}(R))\quad\mbox{and}\quad(R^{{\perp\!\!\!\!\perp}})^{{\perp\!\!\!\!\perp}}\>.$ (3)
It is easy to prove that $R\subset(R^{{\perp\!\!\!\!\perp}})^{{\perp\!\!\!\!\perp}}$. If $R\subset{\mathbb{M}}$, $(R^{{\perp\!\!\!\!\perp}})^{{\perp\!\!\!\!\perp}}$ turns out to consist of the points $p\in{\mathbb{M}}$ such that every causal straight line passing through $p$ meets $R$ somewhere [^1] This is equivalent, in $\mathbb{M}$}, to the set of points $p$ such that every inextensible causal curve passing through $p$ also meets $R$ somewhere. This latter set, in a generic spacetime $M$, is also called the domain of dependence of $R$ when $R$ is achronal..
A Minkowskian reference frame – physically representing an inertial reference frame or an observer – is a unit timelike vector $n\in{\mathsf{T}}_{+}$. A Cartesian coordinate system $\psi:{\mathbb{M}}\ni p\mapsto(x^{0},x^{1},x^{2},x^{3})\equiv(x^{0},\vec{x})\in{\mathbb{R}}\times{\mathbb{R}}^{3}$ with origin $o\in{\mathbb{M}}$ and axes $e_{0},e_{1},e_{2},e_{3}\in\mathsf{V}$, is a Minkowskian coordinate system if the basis $\{e_{0},e_{1},e_{2},e_{3}\}=\{\partial_{x^{0}},\partial_{x^{1}},\partial_{x^{2}},\partial_{x^{3}}\}$ is $g$-orthonormal: $g(e_{a},e_{b})=g(\partial_{x^{a}},\partial_{x^{b}})=g_{ab}$, where $g_{ab}]=\eta:=diag(-1,1,1,1)$ and $e_{0}=\partial_{x^{0}}\in{\mathsf{T}}_{+}$. The Minkowskian coordinate system $\psi$ is adapted to or comoving with $u\in{\mathsf{T}}_{+}$ if $\partial_{x^{0}}=u$.
Given Minkowskian coordinates, vectors are decomposed as $\mathsf{V}\ni v\equiv(v^{0},\vec{v})$ where, according to the Einstein summation convention we adopt henceforth, $v=v^{\mu}e_{\mu}=v^{\mu}\partial_{x^{\mu}}$ so that $g(u,v)=u\cdot v=-u^{0}v^{0}+\vec{u}\cdot\vec{v}\>.$ Here $v^{0}$ is called the temporal component of $v$ and $\vec{v}$ are called the spatial components of $v$ referred to the said Minkowskian reference frame (or Minkowskian coordinate system).
We shall take advantage of tensorial notation and of the raising- and lowering-index procedure, so that, for instance, in Minkowskian coordinates, $p_{\mu}=g_{\mu\nu}p^{\nu}$, ${T^{\mu}}_{\nu}=T^{\mu\alpha}g_{\alpha\nu}$.
A rest space of a Minkowski reference frame $u\in{\mathsf{T}}_{+}$ is an affine $3$-plane $\Sigma\subset{\mathbb{M}}$ $g$-normal to $u$, written $u\perp\Sigma$. If $o\in{\mathbb{M}}$ is a given origin, the family of rest spaces of $u\in{\mathsf{T}}_{+}$ is labeled by the time at which they occur in the said reference frame, $t_{o,\Sigma}=-(p-o)\cdot u$, which does not depend on $p\in\Sigma$. Every spacelike affine $3$-plane is the rest space of a Minkowskian reference frame, indicated by $u_{\Sigma}$ and said to be adapted to $\Sigma$, at some time. $u_{\Sigma}$ is the future-directed unit normal vector to $\Sigma$, which is necessarily timelike.
A rest space meets exactly once every straight line $p(r)=p_{0}+ru$, $r\in{\mathbb{R}}$, parallel to any given causal vector $u\in\overline{\mathsf{V}_{+}}\setminus\{0\}$ and passing through any given $p_{0}\in{\mathbb{M}}$. That is because a rest space is a spacelike smooth Cauchy surface [ONe83] of $\mathbb{M}$}: more generally, it meets exactly once every inextendible smooth causal curve.
$\mathscr{B}(\Sigma)$ denotes the family of Borel subsets of a spacelike $3$-plane $\Sigma$ and $\mathscr{B}_{b}(\Sigma)\subset\mathscr{B}(\Sigma)$ the subfamily of bounded elements, boundedness being equivalently referred to any of the said coordinate systems. $d\Sigma$ denotes the natural translationally invariant Borel measure on the rest space $\Sigma$ of $u_{\Sigma}$, which coincides with the Lebesgue measure on the ${\mathbb{R}}^{3}$ space of the spatial coordinates of any Minkowskian coordinate system $x^{0},x^{1},x^{2},x^{3}$ comoving with $u_{\Sigma}$. It is easy to prove that $d\Sigma$ does not depend on the choice of such a Minkowskian coordinate system. We use the notation $|\Delta|=\int_{\Sigma}\chi_{\Delta}(x)d\Sigma(x)=\int_{{\mathbb{R}}^{3}}\chi_{\Delta}(x)d^{3}x$ for the Lebesgue measure of $\Delta\in\mathscr{B}(\Sigma)$.
The Lie group of metric-preserving affine maps $h:{\mathbb{M}}\to{\mathbb{M}}$ is known as the Poincaré group $IO(1,3)$. The Lie subgroup of affine maps that also preserve the time orientation is called the orthochronous Poincaré group $IO(1,3)_{+}$. The subgroup of $IO(1,3)_{+}$ that leaves fixed an arbitrarily chosen origin [^2] Different choices of $o$ give rise to isomorphic definitions and the component $\Lambda$ of an element of $IO(1,3)_{+}$ does not depend on the choice of $o$. $o\in{\mathbb{M}}$ is the orthochronous Lorentz group $O(1,3)_{+}$. Elements $h\in IO(1,3)_{+}$ are in one-to-one correspondence with the pairs $(\Lambda,v)$ where $v\in\mathsf{V}$ and $\Lambda\in O(1,3)_{+}$, and the action of $h$ on a point $q\in{\mathbb{M}}$ is $(\Lambda,v)q=o+v+\Lambda(q-o)$. In a given Minkowskian coordinate system centered at $o\in{\mathbb{M}}$, the transformations $\Lambda\in O(1,3)_{+}$ are in one-to-one correspondence with the matrices, denoted by the same symbol, $\Lambda\in GL(4,{\mathbb{R}})$ such that $\Lambda^{t}\eta\Lambda=\eta$ and ${\Lambda^{0}}_{0}>0$, where $\eta=diag(-1,1,1,1)$ as above.
The subgroup of $IO(1,3)_{+}$ known as the proper orthochronous Poincaré group $ISO(1,3)_{+}$ is obtained by replacing $O(1,3)_{+}$ with the proper orthochronous Lorentz group $SO(1,3)_{+}$. The latter, representing $O(1,3)_{+}$ in a Minkowskian coordinate system as above, is constructed by restricting to the Lorentz matrices with $\det\Lambda>0$.
B. Quantum observables in Hilbert space. For every set $\cal S$$\subset{{\cal H}}$, the latter being a Hilbert space, the span of $\cal S$}, denoted by $span\>{\cal S}\subset{{\cal H}}$, is the subspace consisting of all complex finite linear combinations of elements of $\cal S$}.
Operators in Hilbert space have their own domains, $A:D(A)\to{{\cal H}}$, where $D(A)\subset{{\cal H}}$ is a subspace. The operations $A+B$, ABAB, aAaA ($a\in{\mathbb{C}}$) are defined on their standard domains: $D(A+B):=D(A)\cap D(B)$, $D(AB):=\{x\in D(B)\>|\>Bx\in D(A)\}$, $D(aA):=D(A)$ unless $a=0$, in which case $D(0A):=D(0):={{\cal H}}$. $A\geq 0$ means that $\langle\psi|A\psi\rangle\geq 0$ if $\psi\in D(A)$. In that case we say that the operator $A$ is positive.
${\mathfrak{B}}({{\cal H}})$ denotes the $C^{*}$-algebra of bounded operators $A:{{\cal H}}\to{{\cal H}}$. If $\in{\mathfrak{B}}({{\cal H}}), ABA\geq B$ (equivalently $B\leq A$) means that $A-B\geq 0$.
[…]
Generally, mixed states $\rho$ are trace-class operators $\rho\in{\mathfrak{B}}_{1}({{\cal H}})$, positive ($\rho\geq 0$), and normalized ($\mathrm{tr}\rho=1$). The convex body of states will be denoted by ${\mathsf{S}}({{\cal H}})$. A special case of states is given by one-dimensional projectors $\rho=|\psi\rangle\langle\psi|$ for unit vectors, $\psi\in{{\cal H}}$. These are pure states, i.e., extremal elements in the space of states if the von Neumann algebra of observables is the whole ${\mathfrak{B}}({{\cal H}})$.
For a state $\rho\in{\mathsf{S}}({{\cal H}})$, $tr(E(B)\rho)=tr(\rho E(B))$ is interpreted as the probability of obtaining an outcome in $B$ when the system is in the state $\rho$.
If $\Sigma({\mathbb{R}})$ is the Borel $\sigma$-algebra $\mathscr{B}({\mathbb{R}})$ and all the effects $E(B)$ are orthogonal projectors, we have a standard Projector-Valued Measure (PVM). As is well known, every PVM is in one-to-one correspondence with a selfadjoint operator $\hat{E}=\int_{{\mathbb{R}}}xE(dx)$ through the spectral theorem. In this sense, a POVM is a generalized observable.
In the special case where $\Sigma(X)$ is the power set of $X:=\{1,\ldots,N\}$, a POVM on $X$ is completely determined by the special effects $E_{j}:=E(\{j\})$ with $j\in\{1,\ldots,N\}$. As a general reference textbook on this mathematical technology applied to physics, we suggest [BLPY16]. References on general spectral theory as applied to physics that we shall use are [Mor18, Mor19]. We assume the reader is familiar with basic properties of von Neumann algebras [Tak02, StZs19]
[…]
One of the fundamental assumptions is relativistic locality: operators belonging to algebras associated with causally separated regions must commute,
$A_{1},A_{2}]=0\quad\mbox{if $A_{1}\in{\mathfrak{A}}({\cal O}_{1})$, $A_{2}\in{\mathfrak{A}}({\cal O}_{2})$ and ${\cal O}_{1}\cap(J^{+}({\cal O}_{2})\cup J^{+}({\cal O}_{2}))=\varnothing$.}$
The net of algebras is assumed to admit a strongly continuous unitary representation of the Abelian translation group $\mathsf{V}$ of $\mathbb{M}$} satisfying the spectral condition: the joint spectrum of the self-adjoint generators must lie in $\mathsf{V}_{+}$. This representation is assumed to extend to a full (strongly continuous) unitary representation of $IO(1,3)_{+}$.
[…]
Most of the features of this approach can be generalized to the case in which $\mathfrak{A}$} and every ${\mathfrak{A}}(\cal O)$ are unital *-algebras, in particular algebras of operators on a given Hilbert space with a common invariant domain. Some further notions and results of the AHK approach will be briefly presented in Sec.6.3.
2 Localization observables and conditional localization
2.1 Localization observables, causal conditions, non-commutativity
The notion of localization used in [Mor26] and in this paper is encapsulated in the following definition of relativistic spatial localization observable. This notion, in a slightly simplified form, was introduced and analyzed in depth for the first time by Castrigiano (see [Cas17] and references therein) under the name of Poincaré covariant POL.
Definition 2.1 (Relativistic Spatial Localization Observable):
A relativistic spatial localization observable is a quadruple $({{\cal H}},{\cal R},A,U)$ where
  • (a)
    ${\cal R}=\cup\{\mathscr{B}(\Sigma)\>|\>\Sigma\subset{\mathbb{M}}\>\mbox{spacelike $3$-plane}\}$ where $\mathscr{B}(\Sigma)$ is the Borel $\sigma$-algebra on $\Sigma$;
  • (b)
[…]
  • (c)
    $U:IO(1,3)_{+}\to{\mathfrak{B}}({{\cal H}})$ is a strongly continuous unitary representation of the orthochronous Poincaré group;
  • (d)
    AA is $U$-covariant, i.e., $U_{g}A^{v}(\Delta)U_{g}^{-1}=A^{gv}(g\Delta)$ for every $g\in{IO(1,3)_{+}}$ and $\Delta\in{\cal R}$, where ${\mathsf{S}}\ni v\mapsto gv\in{\mathsf{S}}$ denotes the action of $IO(1,3)_{+}$} on the tensors in $\mathsf{S}$}.
$({{\cal H}},{\cal R},A,U)$ is of positive-energy type if the selfadjoint generator of spacetime translations ${\mathbb{R}}\ni r\mapsto U_{(I,rv)}$ is positive for $v\in{\mathsf{T}}_{+}$. $\blacksquare$
For future convenience, observe that, if $V\in{\mathfrak{B}}({{\cal H}})$ is a unitary operator and $({{\cal H}},{\cal R},A,U)$ is a relativistic spatial localization observable, then $({{\cal H}},{\cal R},A_{V},U_{V})$ is, where $A^{v}_{V}(\Delta):=VA^{v}(\Delta)V^{\dagger}$ and $(U_{V})_{g}:=VU_{g}V^{\dagger}$ is another relativistic spatial localization observable. It also satisfies the causality condition below and is of positive-energy type if $({{\cal H}},{\cal R},A,U)$ does.
Let us now turn to the causality condition imposed on relativistic localization observables in order to comply with locality constraints at the level of detection probabilities, which cannot evolve superluminally. The condition stated below, due to Castrigiano, is stronger than the original one formulated by Hegerfeldt, which however ruled out all positive-energy relativistic spatial localization observables described by PVMs, as the one associated to the triple of Newton-Wigner operators. A discussion of the relevance of this type of conditions appears in [Cas17, Mor23, Cas25, CDM26].
Definition 2.2:
A relativistic spatial localization observable $({{\cal H}},{\cal R},A,U)$ is causal if it satisfies the following. For every $\Delta\in{\cal R}$, every spacelike $3$-plane $\Sigma$, and every $v\in{\mathsf{S}}$,
  • ($C$)
    $A^{v}(\Delta)\leq A^{v}(\Delta_{\Sigma})\quad\mbox{where $\Delta_{\Sigma}:=\Sigma\cap(J^{+}(\Delta)\cup J^{-}(\Delta))$ provided $\Delta_{\Sigma}\in{\cal R}$.}$ $\blacksquare$
The explicit examples of relativistic spatial localization observables, especially of positive-energy type, presented in [Cas17, Mor23, DRM24, Cas24, Cas25, CDM26] for various types of particles satisfy CC. They therefore show that Hegerfeldt’s causality issue can in fact be regarded as harmless when one considers certain unsharp notions of localization, whereas sharp ones are ruled out.
In [Mor26], after an accurate analysis of relativistic locality from Busch’s perspective, we have asserted that, dealing with relativistic spatial localization observable (also satisfying CC), there is no compelling reason for requiring that $A^{v}(\Delta),A^{v^{\prime}}(\Delta^{\prime})]=0$ when $\Delta$ and $\Delta^{\prime}$ are sharply causally separated, i.e., they are included in respective open regions which are causally separated. This is true for systems like particles which have the propensity to localize in a unique position of a rest space when the rest space is ideally filled with detectors. The found result is in agreement with the celebrated achievement by Halvorson and Clifton [HaCl02] which proves that the above commutativity is actually forbidden if $({{\cal H}},{\cal R},A,U)$ is of positive-energy type [^4] This is even true with a weaker notion of localization observable than the one of Definition 2.2, see the review in [Mor26].. Failure of commutativity has an important consequence. If $\Delta$ and $\Delta^{\prime}$ are bounded regions contained in causally separated open sets $\cal O$} and $\cal O$$^{\prime}$ of $\mathbb{M}$}, one concludes that $A^{v}(\Delta)$ and $A^{v}(\Delta^{\prime})$ cannot belong to the corresponding local algebras of observables in the AHK approach. Indeed, if this were the case, they would necessarily commute. This shows that the effects $A^{v}(\Delta)$ of a relativistic spatial localization observable are not local observables in AHK sense.
2.2 Conditional localization and commutativity
As discussed in [Mor26], realistic localization experiments in $\mathbb{M}$} are performed in laboratories, which are spatially finite regions that are also causally complete in their spacetime description: everything that may happen in such a spacetime region should be determined by physical actions performed within it, at least at a macroscopic level.
If $\Sigma$ is a flat $3$-dimensional plane, a laboratory $L(\Delta_{0})\subset{\mathbb{M}}$, with space given by a bounded set $\Delta_{0}\in\mathscr{B}(\Sigma)$, is defined as the causal completion $L(\Delta_{0}):=(\Delta_{0}^{\perp\!\!\!\!\perp})^{\perp\!\!\!\!\perp}$. If $\Delta_{0}$ is open in the relative topology, the resulting open set $L(\Delta_{0})$ is a globally hyperbolic spacetime in its own right, with $\Delta_{0}$ as a possible smooth spacelike Cauchy surface, as well as other curved smooth spacelike Cauchy surfaces.
[…]
$B^{V}_{\Delta_{0}}(\Delta)=V\frac{1}{\sqrt{A(\Delta_{0})}}A(\Delta)\frac{1}{\sqrt{A(\Delta_{0})}}V^{\dagger}\>,\quad\Delta\in\mathscr{B}(\Delta_{0})$ (4)
where $V\in{\mathfrak{B}}({{\cal H}})$ is a given unitary operator. We assumed above that $A(\Delta_{0})$ is strictly positive; a discussion of this technical point can be found in [Mor26].
As established in [Mor26], for $0\leq\delta<1$ and all $\rho\in{\mathsf{S}}({{\cal H}})$, the so-called gentle measurement lemma entails
$tr(\rho A_{V}(\Delta_{0}))\geq 1-\delta\quad\mbox{implies}\quad\left|tr\left({\rho}B^{V}_{\Delta_{0}}(\Delta)\right)-\frac{tr(\rho A_{V}(\Delta))}{tr(\rho A_{V}(\Delta_{0}))}\right|\leq 2\sqrt{\delta}+\delta\>.$ (5)
In other words, if we start with a state $\rho\in{\mathsf{S}}({{\cal H}})$ whose probability of finding the system in $\Delta_{0}$ is close to $1$, $B^{V}_{\Delta_{0}}(\Delta)$ measures the fraction of detections in $\Delta$ over the total number of detections observed in the laboratory $\Delta_{0}$, when an ensemble of identical systems is prepared in the initial state $\rho$} and the relativistic spatial localization observable with effects ${\cal R}\ni\Delta\mapsto A_{V}(\Delta):=VA(\Delta)V^{\dagger}$ is used to define the spatial localization of the system. In this sense, $B^{V}_{\Delta_{0}}$ represents, at least in the above limit, a conditional localization observable: $B^{V}_{\Delta_{0}}(\Delta)$ gives the probability of finding the system in $\Delta\subset\Delta_{0}$, provided that it is known to be found in the bounded measurable set $\Delta_{0}$. This interpretation becomes increasingly accurate as the initial probability of finding the system in $\Delta_{0}$ approaches unity.
[…]
In [Mor26] we asserted that, in contrast to what happens for the effects of relativistic spatial localization observables, the effects $B^{V}_{\Delta_{0}}(\Delta)$ and $B^{\prime V^{\prime}}_{\Delta^{\prime}_{0}}(\Delta^{\prime})$ may commute if $\Delta_{0}$ and $\Delta^{\prime}_{0}$ are (sharply) causally separated. The present paper aims to construct similar local conditional localization observables for a system described by a free real scalar quantum field. More precisely, the operators $B^{V}_{\Delta_{0}}(\Delta)$ that we shall introduce will be elements of local von Neumann algebras of observables ${\mathfrak{W}}({\cal O})$ generated by the corresponding local Weyl algebras ${{\cal W}}(\cal O)$, where $\cal O$} is a sufficiently large open double cone such that, in particular, ${\cal O}\supset\Delta_{0}$.
3 Elements of free QFT in Minkowski spacetime
The next sections review the elementary rigorous formulation of quantum field theory for a free real scalar field in Minkowski spacetime, in the Fock representation associated with the Minkowski vacuum state. The exposition is deliberately elementary, pedagogical and, as far as possible, self-contained (see [KhMo15] for a more advanced review of QFT in curved spacetime in the algebraic formalism). The relevant notions as in particular, some elementary foundational aspects of the Araki-Haag-Kastler approach, are introduced progressively. It is however throughout assumed that the reader is familiar with the mathematical notions of *-algebra, $C^{*}$-algebra, von Neumann algebra. The goal is to bring closer the community working on quantum measurement theory and the community working on local quantum field theory.
3.1 Bosonic Fock space and particle states
We shall refer here to the mathematical formulation of the free real scalar boson quantum field as presented in Section X.7 of [ReSa75], but using different notation in order to make contact with the results in [Mor23, DRM24]. The complex Schwartz space on ${\mathbb{R}}^{n}$ is denoted by $\mathscr{S}({\mathbb{R}}^{n})$ and $\mathscr{D}({\mathbb{R}}^{n}):=C_{c}^{\infty}({\mathbb{R}}^{n})\subset\mathscr{S}({\mathbb{R}}^{n})$ denotes the space of complex test functions. The respective real subspaces of real-valued functions are denoted by $\mathscr{S}_{\mathbb{R}}({\mathbb{R}}^{n})$ and $\mathscr{D}_{\mathbb{R}}({\mathbb{R}}^{n})$.
[…]
${{\cal H}}^{(1)}:={{\cal H}}^{1}:={{\cal H}}\>,\quad{{\cal H}}^{n}:={{\cal H}}\otimes\cdots\mbox{\scriptsize(n times)}\cdots\otimes{{\cal H}}\quad\mbox{and}\quad{{\cal H}}^{(n)}=S_{n}{{\cal H}}^{n}\quad\mbox{if $n>1$}\>,$
where $S_{n}:{{\cal H}}^{n}\to{{\cal H}}^{n}$ is the orthogonal projector onto the completely symmetric subspace of ${{\cal H}}^{(n)}$ under the action of the unitary representation of the permutation group of $n$ elements. The (separable) Hilbert space in which we develop our theory will be the bosonic Fock space ${\mathfrak{F}}_{s}({{\cal H}}^{(1)})$ of scalar particles of mass $m>0$
${\mathfrak{F}}_{s}({{\cal H}}):=\bigoplus_{n=0}^{+\infty}{{\cal H}}^{(n)}\quad\mbox{where in our case}\quad{{\cal H}}={{\cal H}}^{(1)}:={{\cal H}}_{m}:=L^{2}(\mathsf{V}_{m,+},d\mu_{m}(p))\>.$ (6)
In (6), the symbol $\oplus$ denotes the Hilbert direct orthogonal sum of Hilbert spaces. The scalar product on the vacuum subspace ${{\cal H}}^{0}:={{\cal H}}^{(0)}:={\mathbb{C}}$ is standard multiplication. The future mass shell $\mathsf{V}_{m,+}$, with $m>0$, and the $O(1,3)_{+}$-invariant measure on it are
$\mathsf{V}_{m,+}:=\{p\in V_{+}\>|\>p\cdot p=-m^{2}\}\>,\quad d\mu_{m}(p):=\frac{d^{3}p}{E(p)}$
The latter identity is valid in every Minkowskian coordinate frame, where one identifies ${\mathbb{R}}^{3}$ with the $3$-space of the spatial part of the four-momenta, and
$p\equiv(p^{0},\vec{p})\quad\mbox{with}\quad p^{0}=-p_{0}:=E(p):=\sqrt{m^{2}+\vec{p}^{2}}\>.$ (7)
Some relevant terminology is listed below where, from now on, if $\Psi\in{\mathfrak{F}}_{s}({{\cal H}}^{(1)})$, $\Psi_{n}$ denotes its component in ${{\cal H}}^{(n)}$: $\Psi=\oplus_{n=0}^{+\infty}\Psi_{n}$
  • (a)
    Each normalized $\Psi=\Psi_{n}\in{{\cal H}}^{(n)}$ is an $n$-particle vector state;
  • (b)
    $\Omega=1\in{\mathbb{C}}={{\cal H}}^{(0)}$ represents the Minkowski vacuum state;
  • (c)
[…]
  • (d)
    A relevant dense subspace is the finite-particle subspace,
[…]
As the above terminology suggests, unit vectors $\Psi\in{\mathfrak{F}}_{s}({{\cal H}}^{(1)})$, when written in terms of states $|\Psi\rangle\langle\Psi|\in{\mathsf{S}}({\mathfrak{F}}_{s}({{\cal H}}^{(1)}))$, represent pure quantum states of a quantum system of quantum particles whose mass is $m$ and whose elementary properties, such as the four-momentum, are described in the one-particle Hilbert space ${{\cal H}}^{(1)}={{\cal H}}_{m}$. The system of particles is associated with a quantum field of real scalar bosonic type, which we shall introduce shortly.
All the notions presented are Poincaré invariant under the unitary strongly continuous representation of the orthochronous Poincaré group defined on the Fock space:
$U:IO(1,3)_{+}\ni g\mapsto U_{g}:=U_{g}^{(0)}\oplus\bigoplus_{n=1}^{+\infty}U_{g}^{(1)}\otimes\cdots\mbox{\scriptsize(n times)}\cdots\otimes U^{(1)}_{g}\in{\mathfrak{B}}({\mathfrak{F}}_{s}({{\cal H}}_{m}))$ (9)
where, if $\psi\in{{\cal H}}^{(1)}={{\cal H}}_{m}$ and $(\Lambda,a)\equiv g\in IO(1,3)_{+}$,
$(U^{(1)}_{(\Lambda,a)}\psi)(p)=e^{-ia\cdot p}\psi\left(\Lambda^{-1}p\right)\>,$ (10)
and $U^{(0)}$ is the trivial representation of $IO(1,3)_{+}$ on ${{\cal H}}^{(0)}={\mathbb{C}}$. In particular, the Minkowski vacuum state represented by $\Omega$ is Poincaré invariant by construction.
From now on, $\mathsf{V}_{m,+}^{n}:=\times_{j=1}^{n}\mathsf{V}_{m,+}$ and $\mathscr{S}(\mathsf{V}^{n}_{m,+})$ is the space of maps $\psi:\mathsf{V}_{m,+}^{n}\to{\mathbb{C}}$ which are in $\mathscr{S}({\mathbb{R}}^{3n})$ in a given Minkowskian coordinate system where $\mathsf{V}_{m,+}$ is identified with ${\mathbb{R}}^{3}$ made of the spatial components $\vec{p}$ of the four-momenta $p\equiv(E(p),\vec{p})$. We therefore write $\mathscr{S}(\mathsf{V}^{n}_{m,+})\equiv\mathscr{S}({\mathbb{R}}^{3n})$ in Minkowskian coordinates. An analogous definition is given for $\mathscr{D}(\mathsf{V}^{n}_{m,+})\equiv\mathscr{D}({\mathbb{R}}^{3n})$. The spaces of distributions $\mathscr{S}^{\prime}(\mathsf{V}^{n}_{m,+})$ and $\mathscr{D}^{\prime}(\mathsf{V}^{n}_{m,+})$ are defined correspondingly. It is easy to see that all these definitions do not depend on the chosen Minkowskian coordinate system.
A pair of dense subspaces is defined according to the previous definitions: the finite-particle Schwartz subspace
${\mathfrak{S}}_{0}:=\{\Psi\in{\mathfrak{F}}_{0}\>|\>\Psi_{n}\in\mathscr{S}(\mathsf{V}_{m,+}^{n}),\>\>n\in{\mathbb{N}}\},$ (11)
and the subspace of finite-particle smooth compactly supported vectors
${\mathfrak{D}}_{0}:=\{\Psi\in{\mathfrak{F}}_{0}\>|\>\Psi_{n}\in\mathscr{D}(\mathsf{V}_{m,+}^{n}),\>\>n\in{\mathbb{N}}\}.$ (12)
Evidently ${\mathfrak{D}}_{0}\subset{\mathfrak{S}}_{0}\subset{\mathfrak{F}}_{0}$.
We leave to the reader the easy proof of the following technical result.
Lemma 3.1:
If $v\in\mathsf{V}$, consider the one-parameter subgroup ${\mathbb{R}}\ni r\mapsto U_{(I,rv)}=e^{irH^{v}}$ of the representation (9). The selfadjoint generator $H^{v}$ satisfies the following.
  • (a)
[…]
  • (b)
    As ${\mathfrak{D}}_{0}$ is dense and made of analytic vectors, and ${\mathfrak{D}}_{0}\subset{\mathfrak{S}}_{0}$, $H^{v}$ is essentially selfadjoint on ${\mathfrak{S}}_{0}$ and ${\mathfrak{D}}_{0}$.
  • (c)
    $H^{v}\Omega=0$ and, for $n=1,2,\ldots$
    $(H^{v}\Psi)(p_{1},\ldots,p_{n})=-\sum_{k=1}^{n}v\cdot p_{k}\Psi_{l}(p_{1},\ldots,p_{n})\quad\mbox{if $\Psi\in{{\cal H}}^{(n)}\cap{\mathfrak{S}}_{0}$.}$ (13)
In the case $v\in\mathsf{V}_{+}$, we call $H^{v}$ the Hamiltonian operator associated with $v$.
3.2 Free Klein-Gordon quantum field in Minkowski spacetime
If $\psi\in{{\cal H}}^{1}$, consider the unique linear continuous extensions of the operators, for every given $n\in{\mathbb{N}}$,
$b_{n}(\psi):{{\cal H}}^{n}\ni\psi_{1}\otimes\cdots\otimes\psi_{n}$ $\mapsto$ $\langle\psi|\psi_{1}\rangle\psi_{2}\otimes\cdots\otimes\psi_{n}\in{{\cal H}}^{n-1}\>,\quad b_{0}:=0$
$b^{\dagger}_{n}(\psi):{{\cal H}}^{n}\ni\psi_{1}\otimes\cdots\otimes\psi_{n}$ $\mapsto$ $\psi\otimes\psi_{1}\otimes\cdots\otimes\psi_{n}\in{{\cal H}}^{n+1}\>.$
Still indicating by $b_{n}(\psi)$ and $b^{\dagger}_{n}(\psi)$ the said extensions, the annihilation and creation operators, respectively $a(\psi):{\mathfrak{F}}_{0}\to{\mathfrak{F}}_{0}$ and $a^{\dagger}(\psi):{\mathfrak{F}}_{0}\to{\mathfrak{F}}_{0}$, are defined as the linear extensions to ${\mathfrak{F}}_{0}$ of the respective maps
$a(\psi)|_{{{\cal H}}^{(n)}}:{{\cal H}}^{(n)}\ni\Psi$ $\mapsto$ $\sqrt{n}S_{n-1}b_{n}(\psi)\Psi\in{{\cal H}}^{(n-1)}\>,\quad a(\psi)|_{{{\cal H}}^{(0)}}:=0$ (14)
$a^{\dagger}(\psi)|_{{{\cal H}}^{(n)}}:{{\cal H}}^{(n)}\ni\Psi$ $\mapsto$ $\sqrt{n+1}S_{n+1}b^{\dagger}_{n}(\psi)\Psi\in{{\cal H}}^{(n+1)}\>.$ (15)
The operators $a(\psi)$ and $a^{\dagger}(\psi)$ thus obtained enjoy some elementary properties (see e.g. [ReSa75].)
Proposition 3.2:
$a(\psi):{\mathfrak{F}}_{0}\to{\mathfrak{F}}_{0}$ and $a^{\dagger}(\psi):{\mathfrak{F}}_{0}\to{\mathfrak{F}}_{0}$, $\psi\in{{\cal H}}_{m}$, satisfy the following.
  • (a)
    ${{\cal H}}^{(1)}\ni\psi\mapsto a(\psi)$ is antilinear and ${{\cal H}}^{(1)}\ni\psi\mapsto a^{\dagger}(\psi)$ is linear; both are $\mathbb{R}$}-linear.
  • (b)
    On their dense and invariant domain ${\mathfrak{F}}_{0}$
[…]
  • (c)
    The bosonic commutation rules hold for every $\psi,\psi^{\prime}\in{{\cal H}}^{(1)}$
    $a(\psi),a^{\dagger}(\psi^{\prime})]=\langle\psi|\psi^{\prime}\rangle I,\quad[a(\psi),a(\psi^{\prime})]=0,\quad[a^{\dagger}(\psi),a^{\dagger}(\psi^{\prime})]=0\>.$
  • (d)
    If $\psi\in{{\cal H}}^{(1)}$ and $\Psi_{n}\in{{\cal H}}^{(n)}$ then
[…]
    where $a^{\#}$ represents either $a$ or $a^{\dagger}$.
  • (e)
    If $\psi\in{{\cal H}}_{m}$ and $g\in IO(1,3)_{+}$ then
    $U_{g}a(\psi)U^{-1}_{g}=a(U_{g}^{(1)}\psi)\>,\quad U_{g}a^{\dagger}(\psi)U^{-1}_{g}=a^{\dagger}(U_{g}^{(1)}\psi)\>.$ (17)
As is well known already from the original formulation of QFT, trying to define field operators $\hat{\phi}(x)$ localized at each point $x$ of spacetime gives rise to insurmountable mathematical difficulties [Haa96]. What one can define is a quantum field operator smeared with test functions $f$ and denoted by $\hat{\phi}[f]$. We therefore move on to define the (free) quantum-field operator smeared with a test function $f\in\mathscr{D}({\mathbb{M}})$, where ${\mathbb{M}}\equiv{\mathbb{R}}^{4}$ according to a Minkowskian coordinate system. We limit ourselves to stating the most relevant elementary technical features of this notion. More information about this classical construction and the physical motivations underpinning this crucial physical tool can be found in the vast literature on the subject (see e.g. [StWi00, Ara09, Haa96]). Regarding the smearing procedure, we have to stress that it is pervasive in rigorous QFT and is the practical procedure used to associate observables with regions of spacetime (where the supports of the smearing functions are localized) in agreement with the basic assumptions of the AHK formulation.
First of all, upon the choice of an origin of $\mathbb{M}$}, the covariant Fourier transform of $f\in\mathscr{S}({\mathbb{M}})$ is
$\hat{f}(p):=\frac{1}{(2\pi)^{2}}\int_{{\mathbb{M}}}e^{-ip\cdot x}f(x)d^{4}x\>.$ (18)
The measure $d^{4}x$ in the integral is the standard Lebesgue measure in every Minkowskian coordinate system, which turns out to be $IO(1,3)_{+}$-invariant.
Definition 3.3:
The real scalar field operator of mass $m>0$ smeared with $f\in\mathscr{D}({\mathbb{M}})$ is the densely defined operator
$\hat{\phi}[f]:=\frac{1}{\sqrt{2}}\left(a({\kappa_{m}\overline{f}})+a^{\dagger}(\kappa_{m}f)\right):{\mathfrak{F}}_{0}\to{\mathfrak{F}}_{s}({{\cal H}}_{m})$ (19)
where we have used the $\mathbb{C}$}-linear map
$\kappa_{m}:\mathscr{S}({\mathbb{M}})\ni f\mapsto\sqrt{2\pi}\hat{f}|_{\mathsf{V}_{+,m}}\in L^{2}(\mathsf{V}_{m,+},d\mu_{m}(p))={{\cal H}}_{m}\>.$ (20)
and the bar in $\overline{f}$ denotes complex conjugation. $\blacksquare$
Remark 3.4:
  • (1)
    Three alternative definitions – all equivalent in Minkowski spacetime – are used in the literature, where the space of smearing functions $\mathscr{D}({\mathbb{R}})$ [Ara09] is respectively replaced by $\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$ or $\mathscr{S}({\mathbb{M}})$ [StWi00] or $\mathscr{S}_{\mathbb{R}}({\mathbb{M}})$. The equivalence is based on two facts. (a) In Minkowski spacetime and with the construction above, $\hat{\phi}[f]=\hat{\phi}[Re(f)]+i\hat{\phi}[Im(f)]\>.$ (b) As is easy to prove, the real subspace $\kappa_{m}(\mathscr{D}_{\mathbb{R}}({\mathbb{M}}))$ satisfies
    $\overline{\kappa_{m}(\mathscr{D}_{\mathbb{R}}({\mathbb{M}}))}=\overline{\kappa_{m}(\mathscr{D}_{\mathbb{R}}({\mathbb{M}}))+i\kappa_{m}(\mathscr{D}_{\mathbb{R}}({\mathbb{M}}))}=\overline{\kappa_{m}(\mathscr{D}({\mathbb{M}}))}={{\cal H}}_{m}\>,$ (21)
    where the bar denotes closure in the topology of ${{\cal H}}_{m}$. (The same properties are valid if one everywhere replaces $\mathscr{D}$ by $\mathscr{S}$.) Identity (21) is equivalent to the fact that $\Omega$ is a pure algebraic state on the abstract *-algebra of the field operators (see, e.g., [KhMo15]).
  • (2)
    The smeared quantum fields constructed as above form an elementary system that satisfies the classical Gårding-Streater-Wightman axioms in Minkowski spacetime [StWi00, Ara09]. $\blacksquare$
To proceed, consider the (non-homogeneous) Klein-Gordon equation in $\mathbb{M}$}
$(\Box-m^{2})\phi=f\>.$ (22)
where the d’Alembert operator $\Box$ is written as $\partial_{\mu}\partial^{\mu}=-\partial^{2}_{x^{0}}+\Delta_{\vec{x}}$ in every Minkowskian coordinate system. That equation admits unique advanced and retarded fundamental solutions, linear maps $A:\mathscr{D}({\mathbb{M}})\to C^{\infty}({\mathbb{M}})\quad\mbox{and}\quad R:\mathscr{D}({\mathbb{M}})\to C^{\infty}({\mathbb{M}})$ respectively, completely defined by the requirement that $supp(Rf)\subset J^{+}(supp(f))$ and $supp(Af)\subset J^{-}(supp(f))$ and, obviously, $(\Box-m^{2})Af=(\Box-m^{2})Rf=f$ for every $f\in\mathscr{D}({\mathbb{M}})$. Their difference, called the causal propagator E:=ARE:=A-R, has the consequent property that $\int_{\mathbb{M}}f(x)(Eg)(x)d^{4}x=0$ if $supp(f)$ and $supp(g)$ are causally separated.
Proposition 3.5:
The field operators defined above satisfy the following properties.
  • (a)
    The ${\mathfrak{F}}_{s}({{\cal H}}_{m})$ subspaces ${\mathfrak{F}}_{0}$, ${\mathfrak{S}}_{0}\subset{\mathfrak{F}}_{0}$, and ${\mathfrak{G}}_{0}\subset{\mathfrak{S}}_{0}$ defined as
    ${\mathfrak{G}}_{0}:=\left\{\Psi\in{\mathfrak{F}}_{0}\>\left|\>\Psi_{n}=span\left\{\prod_{k=1}^{n}\right.\hat{\phi}(f^{(n)}_{k})\Omega,\quad f^{(n)}_{k}\in\mathscr{D}({\mathbb{M}})\>,k=1,\ldots,n\right\}\right\}\>,$ (23)
    are dense, $\hat{\phi}[f]$-invariant, and made of analytic vectors of $\hat{\phi}[f]$ with $f\in\mathscr{D}({\mathbb{M}})$.
  • (b)
    If $f\in\mathscr{D}({\mathbb{R}}^{4})$ then $\hat{\phi}[\overline{f}]\subset\hat{\phi}[f]^{\dagger}$ so that $\hat{\phi}[f]$ is closable. More strongly
    $\overline{\hat{\phi}[f]}=\hat{\phi}[f]^{\dagger}\quad\mbox{if $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$}\>,$
    namely $\hat{\phi}[f]$ is essentially selfadjoint if smeared with real functions. According to (a), ${\mathfrak{F}}_{0}$, ${\mathfrak{S}}_{0}$, and ${\mathfrak{G}}_{0}$ are cores of $\hat{\phi}[f]$ for $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$.
  • (c)
    $\mathscr{D}({\mathbb{M}})\ni f\mapsto\hat{\phi}[f]$ enjoys the following further properties.
    • (c1)
      CCR. The canonical commutation rules hold:
      $\hat{\phi}[f],\hat{\phi}[g]]=-iE(f,g)I:=-i\int_{\mathbb{M}}f(x)(Eg)(x)d^{4}xI\quad\mbox{for}\quad f,g\in\mathscr{D}({\mathbb{M}})\>,$
      so that $\hat{\phi}[f],\hat{\phi}[g]]=0$ if $supp(f)$ and $supp(g)$ are causally separated
    • (c2)
      KG equation. It solves the homogeneous Klein-Gordon equation in the distributional sense:
      $\hat{\phi}[(\Box-m^{2})f]=0\quad\mbox{for}\quad f\in\mathscr{D}({\mathbb{M}})\>.$
    • (c3)
      $IO(1,3)_{+}$-covariance. The representation (9) of $IO(1,3)_{+}$ acts covariantly on the field operator:
      $U_{g}\hat{\phi}[f]U^{-1}_{g}=\hat{\phi}[g_{*}f]\quad\mbox{for $g\in IO(1,3)_{+}$ and $f\in\mathscr{D}({\mathbb{M}})$.}$
      where $(g_{*}f)(x):=f(g^{-1}x)$ for every $x\in{\mathbb{M}}$.
    • (c4)
      The Weyl generators $W(f):=e^{i\overline{\hat{\phi}(f)}}$ satisfy the Weyl relations for every $f,g\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$
      $W(f)W(g)=e^{iE(f,g)/2}W(f+g)\>,\quad W(f)^{\dagger}=W(-f)\>,\quad W(0)=I\>.$ (24)
      In particular $[W(f),W(g)]=0$ if $supp(f)$ and $supp(g)$ are causally separated.
  • (d)
    Reeh-Schlieder property.
    • RS1
      If ${\cal O}\subset{\mathbb{M}}$ is open, bounded, and non-empty, the unital *-algebra ${\cal A}({\cal O})$ generated by the field operators $\hat{\phi}[f]$ for smearing functions such that $supp(f)\subset{\cal O}$ satisfies $\overline{{\cal A}({\cal O})\Omega}={\mathfrak{F}}_{s}({{\cal H}}_{m})$.
    • RS2
      If ${\cal O}\subset{\mathbb{M}}$ is open, bounded, and non-empty, the $C^{*}$-algebra ${\cal W}({\cal O})$ generated by the Weyl operators $W[f]$ for smearing functions such that $supp(f)\subset{\cal O}$ satisfies $\overline{{\cal W}({\cal O})\Omega}={\mathfrak{F}}_{s}({{\cal H}}_{m})$.
Proof.
(a), (b), and (c3) are consequences of Proposition 3.2, see e.g. [ReSa75], paying attention to the use of different notation and taking Remark 3.4 into account. For (c) see [ReSa75] and [KhMo15] for a generic curved globally hyperbolic spacetime and a quasifree state. A proof of the version of the Reeh-Schlieder property presented in (d)RS1 can be obtained by adapting the more general result stated in Theorem 4-2 of [StWi00]; see Theorem 4.14 in [Ara09] for the version RS2 applied to the case of a free scalar field. ∎
Another important feature of the free field operators introduced above and the associated Weyl algebra is known as Haag duality. We shall briefly discuss it in Section 6.3.
Due to Remark 3.4, in Minkowski spacetime the Proposition 3.5 is still valid [^7] CCR need a more delicate readaptation if one smears with non-compactly supported functions, due to the nature of the advanced and retarded solutions of the KG equation. if one replaces $\mathscr{D}({\mathbb{M}})$ (resp. $\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$) by $\mathscr{S}({\mathbb{M}})$ (resp. $\mathscr{S}_{\mathbb{R}}({\mathbb{M}})$) and changes the statements accordingly. $({\mathfrak{F}}_{s}({{\cal H}}_{m}),{\mathfrak{G}}_{0},\pi,\Omega)$ is the GNS structure of the Minkowski-vacuum representation of the unital *-algebra $\cal A$}} called the CCR algebra generated by abstract field operators $\phi[f]$, where $\pi$ is the *-homomorphism induced by $\pi(\phi[h])=\hat{\phi}[f]$ [KhMo15].
3.3 Free Quantum Fields in $\mathbb{M}$} as Quadratic forms
The stress-energy operator is a special case of a Wick polynomial. There are at least two procedures to define them in Minkowski spacetime: one can be extended to general globally hyperbolic spacetimes and refers to Hadamard states, using the powerful machinery of microlocal analysis (see e.g. [KhMo15]). The other, the older one, quite familiar to theoretical physicists, is easier to handle when dealing with Minkowski spacetime and the Poincaré-invariant vacuum state $\Omega$. The language used here is that of quadratic forms. In this pedagogical discussion this older approach is more suitable, also because in the specific case of Minkowski spacetime it permits explicit computations. To this end it is convenient to define a pair of quadratic forms representing the formal operators denoted by $a_{p}$ and $a^{\dagger}_{p}$ in the physical literature, where $p\in\mathsf{V}_{m,+}$.
If $p\in\mathsf{V}_{m,+}$ and $\Psi\in{\mathfrak{S}}_{0}$, the operator $a_{p}:{\mathfrak{S}}_{0}\to{\mathfrak{S}}_{0}$ such that $a_{p}({{\cal H}}^{(n)}\cap{\mathfrak{S}}_{0})\subset{{\cal H}}^{(n-1)}\cap{\mathfrak{S}}_{0}$ is defined as the linear extension of
$(a_{p}\Psi)(k_{1},\ldots,k_{n-1}):=\sqrt{n}\Psi(p,k_{1},\ldots,k_{n})\quad\mbox{for $\Psi\in{{\cal H}}^{(n)}$ and $a_{p}\Omega:=0$}\>.$ (25)
A quadratic form is well defined on ${\mathfrak{S}}_{0}\times{\mathfrak{S}}_{0}$ as
$\langle\Psi^{\prime}|a^{\dagger}_{p}\Psi\rangle=\sum_{n=1}^{+\infty}\int_{{\mathbb{R}}^{3n}}\overline{\Psi^{\prime}_{n}}(k_{1},\ldots,k_{n})(a^{\dagger}_{p}\Psi)_{n}(k_{1},\ldots,k_{n})d\mu_{m}(k_{1})\cdots d\mu_{m}(k_{n})\>.$ (26)
In the integral, $a^{\dagger}_{p}:{{\cal H}}^{(n)}\cap{\mathfrak{S}}_{0}\to\mathscr{S}^{\prime}(\mathsf{V}_{m,+}^{n+1})$ is the map extended by linearity, such that
$(a^{\dagger}_{p}\Psi)(k_{1},\ldots,k_{n+1}):=\sqrt{n+1}\sum_{l=1}^{n+1}\frac{\delta(p,k_{l})\Psi(k_{1},\ldots,k_{l-1},k_{l+1},\ldots,k_{n+1})}{n+1}\>,\quad\Psi\in{{\cal H}}^{(n)}$ (27)
where the Dirac delta $\delta(p,k)$ refers to the mass shell $\mathsf{V}_{m,+}$ and its invariant measure, and the integral in (26) has a distributional meaning, as is appropriate since $\Psi\in{\mathfrak{S}}_{0}$. By direct inspection we have the adjunction relation of quadratic forms:
$\langle\Psi^{\prime}|a^{\dagger}_{p}\Psi\rangle=\langle a_{p}\Psi^{\prime}|\Psi\rangle\quad\mbox{for}\quad\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}\>.$ (28)
According to this identity we can give a general definition.
Definition 3.6:
If $N,M=0,1,2,\ldots$ and the corresponding momenta are $p_{1},\ldots,p_{N}\in\mathsf{V}_{m,+}$, $k_{1},\ldots,k_{M}\in\mathsf{V}_{m,+}$, we define the normally ordered quadratic form,
$\left\langle\Psi^{\prime}\left|\prod_{j=1}^{N}a^{\dagger}_{p_{j}}\prod_{r=1}^{M}a_{k_{r}}\Psi\right.\right\rangle:=\left\langle\prod_{j=1}^{N}a_{p_{j}}\Psi^{\prime}\left|\prod_{r=1}^{M}a_{k_{r}}\Psi\right.\right\rangle\quad\mbox{for}\quad\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}\>,$ (29)
where the right-hand side is a proper inner product and we used definition (25). $\blacksquare$
It is easy to see that, if $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$, then the following map is a function in $\mathscr{S}(\mathsf{V}_{m,+}^{3N+3M})$:
[…]
Proposition A.1 in the appendix states the most important properties of normally ordered quadratic forms. It would not be possible to define, analogously to Definition 3.6, quadratic forms corresponding to a symbol like $\prod_{r=1}^{M}a_{k_{r}}\prod_{j=1}^{N}a^{\dagger}_{p_{j}}$ with $M,N\geq 1$ because, from our perspective, $a^{\dagger}_{p}$ has to be understood in the sense of quadratic forms on ${\mathfrak{S}}_{0}$, and objects like $\langle a_{p^{\prime}}\Psi^{\prime}|a_{p}\Psi\rangle$ are not defined if $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$.
Definition 3.7:
The quantum-field quadratic form is the quadratic form
$\langle\Psi^{\prime}|\hat{\phi}(x)\Psi\rangle:=\frac{1}{\sqrt{2}(2\pi)^{3/2}}\int_{{\mathbb{R}}^{3}}e^{ip\cdot x}\langle\Psi^{\prime}|a_{p}\Psi\rangle+e^{-ip\cdot x}\langle\Psi^{\prime}|a^{\dagger}_{p}\Psi\rangle\>\>d\mu_{m}(p)\>,\quad\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$ (30)
for $x\in{\mathbb{M}}$. $\blacksquare$
Notation 3.8:
The above definition is formally written as
$\hat{\phi}(x):=\frac{1}{\sqrt{2}(2\pi)^{3/2}}\int_{{\mathbb{R}}^{3}}e^{ip\cdot x}a_{p}+e^{-ip\cdot x}a^{\dagger}_{p}\>\>d\mu_{m}(p)\>$ (31)
and we shall use this notation throughout. $\blacksquare$.
The right-hand side of (30) defines a smooth bounded function of $x\in{\mathbb{M}}$ due to (106). This function can therefore be smeared with functions $f\in\mathscr{S}({\mathbb{M}})$ and $\mathscr{D}({\mathbb{M}})$.
Proposition 3.9:
$\mathscr{D}({\mathbb{M}})\ni f\mapsto\langle\Psi^{\prime}|\hat{\phi}[f]\Psi\rangle$ is a distribution in $\mathscr{D}^{\prime}({\mathbb{M}})$ if $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$ and
$\langle\Psi^{\prime}|\hat{\phi}[f]\Psi\rangle=\int_{{\mathbb{R}}^{4}}\langle\Psi^{\prime}|\hat{\phi}(x)\Psi\rangle f(x)d^{4}x\quad\mbox{if $f\in\mathscr{D}({\mathbb{M}})$, $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$.}$ (32)
Proof.
(19), (107), (30) and Fubini-Tonelli’s theorem yield (32). The map $\mathscr{D}_{\mathbb{R}}({\mathbb{R}}^{4})\ni f\mapsto\int_{{\mathbb{R}}^{4}}\langle\Psi^{\prime}|\hat{\phi}(x)\Psi\rangle f(x)d^{4}x\in{\mathbb{C}}$ defines a distribution in $\mathscr{D}^{\prime}({\mathbb{R}}^{4})$, for every $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$, simply because ${\mathbb{R}}^{4}\ni x\mapsto\langle\Psi^{\prime}|\hat{\phi}(x)\Psi\rangle$ is smooth. ∎
4 The stress-energy tensor operator of the Klein-Gordon field in $\mathbb{M}$}
The construction of the previous section can be extended to normally ordered Wick polynomials in Minkowski spacetime and referred to the Poincaré-invariant vacuum $\Omega$. We only consider two special cases, both of second order: the $\phi^{2}$ field and the stress-energy tensor $T_{\mu\nu}$ in Minkowski spacetime. The completely covariant formulation in curved spacetime for Hadamard states has a rather long history (see [KhMo15] and references therein for a review); a specific discussion, together with recent results on normally ordered second-order Wick polynomials in curved spacetime referred to Hadamard states, appears in [San12].
4.1 $\phi^{2}$ and stress-energy-momentum tensor operators
Classically, the stress-energy(-momentum) tensor of a (smooth) real Klein-Gordon scalar field $\phi$ of mass $m>0$ is defined as the symmetric second-order tensor field on $\mathbb{M}$} whose components in an (arbitrary) coordinate representation are
$T_{\mu\nu}(x):=\partial_{\mu}\phi(x)\partial_{\nu}\phi(x)-\frac{1}{2}g_{\mu\nu}(\partial_{\alpha}(x)\phi\partial^{\alpha}(x)\phi+m^{2}\phi(x)^{2})\>.$ (33)
Due to the Klein-Gordon equation $(\Box-m^{2})\phi=0$, the conservation equation
$\partial_{\mu}{T^{\mu\nu}}(x)=0\>$ (34)
is valid. We move on to study the quantized version of the stress-energy tensor. Since it will be useful in Sect. 4.2, it is also convenient to introduce the quantum version of the squared field $\phi^{2}$.
Definition 4.1:
If $x\in{\mathbb{M}}$, the (normally ordered) $\phi^{2}$ quadratic form at $x$ is the quadratic form
${\mathfrak{S}}_{0}\times{\mathfrak{S}}_{0}\ni(\Psi^{\prime},\Psi)\mapsto\langle\Psi^{\prime}|:\thinspace\hat{\phi}^{2}\thinspace:(x)\Psi\rangle\>,$
whereas the stress-energy tensor quadratic form at $x$ is the assignment to every Minkowskian reference frame of a corresponding set of $16$ quadratic forms for $\mu,\nu=0,1,2,3$
${\mathfrak{S}}_{0}\times{\mathfrak{S}}_{0}\ni(\Psi^{\prime},\Psi)\mapsto\langle\Psi^{\prime}|:\thinspace\hat{T}_{\mu\nu}\thinspace:(x)\Psi\rangle\>.$
The right-hand sides of the formulas above are obtained by respectively replacing $\phi(x)$ with $\hat{\phi}(x)$ in $\phi(x)\phi(x)$ and in the expression (33) of $T_{\mu\nu}(x)$, next expanding $\hat{\phi}(x)$ according to (31), next moving the operator $a$ before the operator $a^{\dagger}$ in the resulting products, and finally computing the matrix element $\langle\Psi^{\prime}|\cdot\Psi\rangle$ for $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$:
$\langle\Psi^{\prime}|:\thinspace\hat{\phi}^{2}\thinspace:(x)\Psi\rangle:=$
$\int_{\mathsf{V}_{m,+}^{2}}\frac{e^{i(p+k)\cdot x}\langle\Psi^{\prime}|a_{p}a_{k}\Psi\rangle}{2(2\pi)^{3}}d^{2}\mu_{m}(p,k)+\int_{\mathsf{V}_{m,+}^{2}}\frac{e^{-i(p+k)\cdot x}\langle\Psi^{\prime}|a^{\dagger}_{p}a^{\dagger}_{k}\Psi\rangle}{2(2\pi)^{3}}d^{2}\mu_{m}(p,k)$
$+\int_{\mathsf{V}_{m,+}^{2}}e^{i(k-p)\cdot x}\frac{\langle\Psi^{\prime}|a^{\dagger}_{p}a_{k}\Psi\rangle}{2(2\pi)^{3}}d^{2}\mu_{m}(p,k)+\int_{\mathsf{V}_{m,+}^{2}}e^{i(p-k)\cdot x}\frac{\langle\Psi^{\prime}|a^{\dagger}_{k}a_{p}\Psi\rangle}{2(2\pi)^{3}}d^{2}\mu_{m}(p,k)\>,$ (35)
$\langle\Psi^{\prime}|:\thinspace\hat{T}_{\mu\nu}\thinspace:(x)\Psi\rangle:=$
$\int_{\mathsf{V}_{m,+}^{2}}\thinspace\thinspace\frac{e^{i(p+k)\cdot x}\langle\Psi^{\prime}|a_{p}a_{k}\Psi\rangle}{2(2\pi)^{3}}t_{\mu\nu}(p,-k)d^{2}\mu_{m}(p,k)+\int_{\mathsf{V}_{m,+}^{2}}\thinspace\thinspace\frac{e^{-i(p+k)\cdot x}\langle\Psi^{\prime}|a^{\dagger}_{p}a^{\dagger}_{k}\Psi\rangle}{2(2\pi)^{3}}t_{\mu\nu}(p,-k)d^{2}\mu_{m}(p,k)$
$+\thinspace\int_{\mathsf{V}_{m,+}^{2}}\thinspace\thinspace\thinspace\thinspace e^{i(k-p)\cdot x}\frac{\langle\Psi^{\prime}|a^{\dagger}_{p}a_{k}\Psi\rangle}{2(2\pi)^{3}}t_{\mu\nu}(p,k)d^{2}\mu_{m}(p,k)+\int_{\mathsf{V}_{m,+}^{2}}\thinspace\thinspace\thinspace\thinspace e^{i(p-k)\cdot x}\frac{\langle\Psi^{\prime}|a^{\dagger}_{k}a_{p}\Psi\rangle}{2(2\pi)^{3}}t_{\mu\nu}(p,k)d^{2}\mu_{m}(p,k)$ (36)
where we introduced the symmetric $(0,2)$ tensor
$t_{\mu\nu}(p,k):=\frac{1}{2}[k_{\mu}p_{\nu}+p_{\mu}k_{\nu}-g_{\mu\nu}(p\cdot k+m^{2})]\>.$ (37)
$\blacksquare$
By construction, and taking (106) into account, if $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$, the maps ${\mathbb{M}}\ni x\mapsto\langle\Psi^{\prime}|:\thinspace\hat{\phi}^{2}\thinspace:(x)\Psi\rangle$ and ${\mathbb{M}}\ni x\mapsto\langle\Psi^{\prime}|:\thinspace\hat{T}_{\mu\nu}\thinspace:(x)\Psi\rangle$ are smooth bounded functions so that they can be smeared with $f\in\mathscr{D}({\mathbb{R}}^{4})$, giving rise to distributions in $\mathscr{D}^{\prime}({\mathbb{M}})$. Furthermore, the components $\langle\Psi^{\prime}|:\thinspace\hat{T}_{\mu\nu}\thinspace:(x)\Psi\rangle$ for given $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$ and $x\in{\mathbb{M}}$ define a symmetric $(0,2)$ tensor when the Minkowskian reference frame is changed. Notice that, once defined in Minkowskian coordinates, the quadratic form of the normally ordered stress-energy tensor operator can be defined as a general symmetric $(0,2)$ tensor, independently of the choice of the type of local coordinates, by the standard local tensor transformation law.
Proposition 4.2:
Take $f\in\mathscr{D}({\mathbb{M}})$ and let us refer to a given Minkowski coordinate system concerning the component indices $\mu,\nu=0,1,2,3$.
There exist unique operators $\thinspace\hat{\phi}^{2}\thinspace:[f]:{\mathfrak{S}}_{0}\to{\mathfrak{F}}_{s}({{\cal H}}_{m})$, $\thinspace\hat{T}_{\mu\nu}\thinspace:[f]:{\mathfrak{S}}_{0}\to{\mathfrak{F}}_{s}({{\cal H}}_{m})$ respectively called (smeared normally ordered) $\phi^{2}$
operator and (smeared normally ordered) stress-energy tensor operator such that
$\langle\Psi^{\prime}|:\thinspace\hat{\phi}^{2}\thinspace:[f]\Psi\rangle$ $=\int_{{\mathbb{R}}^{4}}f(x)\langle\Psi^{\prime}|:\thinspace\hat{\phi}^{2}\thinspace:(x)\Psi\rangle d^{4}x\quad\forall\Psi^{\prime},\Psi\in{\mathfrak{S}}_{0}\>,$ (38)
$\langle\Psi^{\prime}|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]\Psi\rangle$ $=\int_{{\mathbb{R}}^{4}}f(x)\langle\Psi^{\prime}|:\thinspace\hat{T}_{\mu\nu}\thinspace:(x)\Psi\rangle d^{4}x\quad\forall\Psi^{\prime},\Psi\in{\mathfrak{S}}_{0}\>.$ (39)
Therefore $\mathscr{D}({\mathbb{M}})\ni f\mapsto\langle\Psi^{\prime}|:\thinspace\hat{\phi}^{2}\thinspace:[f]\Psi\rangle$ and $\mathscr{D}({\mathbb{M}})\ni f\mapsto\langle\Psi^{\prime}|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]\Psi\rangle$ belong to $\mathscr{D}^{\prime}({\mathbb{M}})$.
The following further facts are true.
  • (a)
    ${\mathfrak{S}}_{0}$ is invariant under $\thinspace\hat{\phi}^{2}\thinspace:[f]$ and $\thinspace\hat{T}_{\mu\nu}\thinspace:[f]$.
  • (b)
    If $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$, the said operators admit adjoints and
    $\thinspace\hat{\phi}^{2}\thinspace:[\overline{f}]\subset:\thinspace\hat{\phi}^{2}\thinspace:[f]^{\dagger}\>,\quad:\thinspace\hat{T}_{\mu\nu}\thinspace:[\overline{f}]\subset:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]^{\dagger}$
    so that, in particular, $\thinspace\hat{\phi}^{2}\thinspace:[f]$ and $\thinspace\hat{T}_{\mu\nu}\thinspace:[f]$ are symmetric.
  • (c)
    $\thinspace\hat{T}_{\mu\nu}\thinspace:[f]=\>\>:\thinspace\hat{T}_{\nu\mu}\thinspace:[f]\quad\forall f\in\mathscr{D}({\mathbb{M}})\>.$
  • (d)
    The representation (9) of $IO(1,3)_{+}$ acts covariantly on the stress-energy tensor operator: if $g\equiv(\Lambda,a)\in IO(1,3)_{+}$ and defining $(g_{*}f)(x):=f(g^{-1}x)$ for $x\in{\mathbb{M}}$,
    $U_{g}:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]U_{g}^{\dagger}$ $={(\Lambda^{-1})_{\mu}}^{\alpha}{(\Lambda^{-1})_{\nu}}^{\beta}:\thinspace\hat{T}_{\alpha\beta}\thinspace:[g_{*}f]\quad\mbox{if $f\in\mathscr{D}({\mathbb{M}})$.}$ (40)
    Analogously,
    $U_{g}:\thinspace\hat{\phi}^{2}\thinspace:[f]U_{g}^{\dagger}$ $=\>:\thinspace\hat{\phi}^{2}\thinspace:[g_{*}f]\quad\mbox{if $f\in\mathscr{D}({\mathbb{M}})$.}$ (41)
  • (e)
    The stress-energy tensor operator is conserved in the distributional sense:
    $:\thinspace\hat{T}_{\mu\nu}\thinspace:[\partial^{\mu}f]=0\>,\quad\mbox{for $f\in\mathscr{D}({\mathbb{M}})$}.$ (42)
Proof.
See Appendix B. ∎
4.2 Locality/commutativity properties of $\hat{\phi}[f]$, $\thinspace\hat{\phi}^{2}\thinspace:[f]$, and $\thinspace\hat{T}_{\mu\nu}\thinspace:[f]$
Once we have defined $\hat{\phi}[f]$, $\thinspace\hat{\phi}^{2}\thinspace:[f]$, $\thinspace\hat{T}_{\mu\nu}\thinspace:[f]$ and the associated quadratic forms, we move on to show another technically fruitful way to compute them in terms of a smearing procedure with compactly supported distributions instead of compactly supported smooth functions. First of all, we define the normally ordered product of two field operators smeared with $f,g\in\mathscr{D}({\mathbb{M}})$ as the operator with domain ${\mathfrak{F}}_{0}$
$:\thinspace\hat{\phi}[f]\hat{\phi}[g]\thinspace:\>:=\hat{\phi}[f]\hat{\phi}[g]-\langle\Omega|\hat{\phi}[f]\hat{\phi}[g]\Omega\rangle I\>.$ (43)
This type of definition, when dealing with algebraic states of Gaussian type, can easily be extended to products of many fields by taking advantage of the so-called Wick rule (see e.g. [KhMo15]). We stick to the elementary case above, since it is sufficient for our purposes. Accordingly, if $\Psi\in{\mathfrak{F}}_{0}$, the maps
$\mathscr{D}({\mathbb{M}})\times\mathscr{D}({\mathbb{M}})\ni(f,g)\mapsto\langle\Psi|\hat{\phi}[f]\hat{\phi}[g]\Psi\rangle\>,\quad\mathscr{D}({\mathbb{M}})\times\mathscr{D}({\mathbb{M}})\ni(f,g)\mapsto\langle\Psi|:\thinspace\hat{\phi}[f]\hat{\phi}[g]\thinspace:\Psi\rangle$
are respectively the 2-point function and the normally ordered 2-point function of the state represented by $\Psi$.
Lemma 4.3:
If $f,g\in\mathscr{D}({\mathbb{M}})$, $\thinspace\hat{\phi}[f]\hat{\phi}[g]\thinspace:$ can be obtained from $\hat{\phi}[f]\hat{\phi}[g]$ by expanding the field operators according to (19) and moving $a$ before $a^{\dagger}$:
$:\thinspace\hat{\phi}[f]\hat{\phi}[g]\thinspace:=\frac{1}{2}\left(a(\kappa_{m}\overline{f})a(\kappa_{m}\overline{g})+a^{\dagger}(\kappa_{m}g)a(\kappa_{m}\overline{f})+a^{\dagger}(\kappa_{m}f)a(\kappa_{m}\overline{g})+a^{\dagger}(\kappa_{m}f)a^{\dagger}(\kappa_{m}g)\right)\>,$ (44)
so that, in particular, $\thinspace\hat{\phi}[f]\hat{\phi}[g]\thinspace:=:\thinspace\hat{\phi}[g]\hat{\phi}[f]\thinspace:$.
Proof.
See Appendix B. ∎
The notions introduced allow one to compute $\langle\Psi|:\thinspace\hat{\phi}^{2}\thinspace:[f]\Psi^{\prime}\rangle$ and $\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]\Psi^{\prime}\rangle$ by means of a procedure known as point-splitting which in particular relies upon Schwartz’ kernel theorem.
Proposition 4.4:
If $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$, the map $\mathscr{D}({\mathbb{M}})\times\mathscr{D}({\mathbb{M}})\ni(f,g)\mapsto\langle\Psi|:\thinspace\hat{\phi}[f]\hat{\phi}[g]\thinspace:\Psi^{\prime}\rangle$ is the restriction to $\mathscr{D}({\mathbb{M}})\times\mathscr{D}({\mathbb{M}})$ of a unique distribution in $\mathscr{D}^{\prime}({\mathbb{M}}\times{\mathbb{M}})$ which is a smooth bounded function, denoted by ${\mathbb{M}}\times{\mathbb{M}}\ni(x,y)\mapsto\langle\Psi|:\thinspace\hat{\phi}(x)\hat{\phi}(y)\thinspace:\Psi^{\prime}\rangle$, symmetric under interchange of its arguments. Furthermore,
$\int f(x)\delta(x,y)\langle\Psi|:\thinspace\hat{\phi}(x)\hat{\phi}(y)\thinspace:\Psi^{\prime}\rangle d^{4}xd^{4}y=\langle\Psi|:\thinspace\hat{\phi}^{2}\thinspace:[f]\Psi^{\prime}\rangle\quad\mbox{if $f\in\mathscr{D}({\mathbb{M}})$}.$ (45)
Referring to Minkowskian coordinates,
$\int f(x)\delta(x,y)D_{\mu\nu}(x,y)\langle\Psi|:\thinspace\hat{\phi}(x)\hat{\phi}(y)\thinspace:\Psi^{\prime}\rangle d^{4}xd^{4}y=\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]\Psi^{\prime}\rangle\quad\mbox{if $f\in\mathscr{D}({\mathbb{M}})$,}$ (46)
where we introduced the (formally selfadjoint) second-order differential operator
$D_{\mu\nu}(x,y):=\frac{1}{2}\left[\partial_{x^{\mu}}\partial_{y^{\nu}}+\partial_{x^{\nu}}\partial_{y^{\mu}}-g_{\mu\nu}\left(g^{\alpha\beta}\partial_{x^{\alpha}}\partial_{y^{\beta}}+m^{2}\right)\right].$ (47)
Proof.
See Appendix B. ∎
The results just obtained imply a first crucial fact about relativistic locality in the spirit of the AHK approach. Here the smearing procedure reveals its physical importance.
Proposition 4.5:
If ${\cal O}\subset{\mathbb{M}}$ is open, ${{\cal T}}({\cal O})$ denotes the unital *-algebra of operators on ${\mathfrak{F}}_{s}({{\cal H}}_{m})$ with common invariant domain ${\mathfrak{S}}_{0}$ generated by (a) $\hat{\phi}$, (b) $\thinspace\hat{\phi}^{2}\thinspace:$, (c) $\thinspace\hat{T}_{\mu\nu}\thinspace:$ smeared with test functions supported in $\cal O$}. If the open sets ${\cal O}_{1}\subset{\mathbb{M}}$ and ${\cal O}_{2}\subset{\mathbb{M}}$ are causally separated, then
$A_{1},A_{2}]=0\quad\mbox{for $A_{1}\in{{\cal T}}({\cal O}_{1})$ and $A_{2}\in{{\cal T}}({\cal O}_{2})$. }$
Proof.
See Appendix B. ∎
4.3 Energy inequalities and Quantum Energy Inequalities
Again taking advantage of Proposition 4.4, we move on to consider energy inequalities. We start from the observation that the classical stress-energy tensor (33) enjoys two important properties. As is well known, the four-momentum density in the Minkowskian reference frame $u\in{\mathsf{T}}_{+}$ is defined as $P_{u}^{\mu}(x):=T^{\mu\nu}(x)u_{\nu}=T^{\nu\mu}(x)u_{\nu}$. Some computations based on the explicit expression (33) imply that
$T^{\mu\nu}(x)u_{\mu}u^{\prime}_{\nu}\geq 0\quad\mbox{if $u,u^{\prime}\in{\mathsf{T}}_{+}$.}$ (48)
We observe that, in curved spacetime, $u$ must be taken to be a Killing vector in order to define a conserved four-momentum density $P_{u}$, whereas $u^{\prime}$ may be chosen as any future-directed timelike vector. In general, their roles cannot be interchanged in curved spacetime, although (48) remains valid in either case. The following elementary fact is true.
Proposition 4.6:
A symmetric tensor $T$ in $\mathsf{V}^{*}\otimes\mathsf{V}^{*}$ satisfies $T_{\mu\nu}u^{\mu}u^{\prime\nu}\geq 0$ for every pair $u,u^{\prime}\in{\mathsf{T}}_{+}$ if and only if $J_{u}\in\overline{\mathsf{V}_{+}}$ for every $u\in{\mathsf{T}}_{+}$, where $J^{\nu}_{u}:=-u^{\mu}{T_{\mu}}^{\nu}$.
As a consequence, the four-momentum density $P_{u}(x)$ is causal and future-directed wherever it does not vanish, for every choice of reference frame $u\in{\mathsf{T}}_{+}$ and also, by linearity and taking an obvious limit, for $u\in\overline{\mathsf{V}_{+}}$. Inequality (48) also makes explicit the requirement of positive energy density of $P_{u}$ in the reference frame $u^{\prime}$, which generalizes [^8] This latter condition, even if valid for every $u\in{\mathsf{T}}_{+}$, does not imply that $J_{u}$ is causal and future-directed: $T\equiv diag(0,1,0,0)$ in Minkowskian coordinates is a counterexample. $T^{\mu\nu}(x)u_{\mu}u_{\nu}\geq 0$.
It is known that, in general, energy positivity requirements fail when we pass to the quantum regime and, even in curved spacetime and considering a covariant notion of normally ordered stress-energy operator, only lower bounds for the expectation value of the energy are valid. This is true for Hadamard and adiabatic algebraic states (see [Few12] for an excellent exhaustive review). However, for the Klein-Gordon (real massive) quantum field the inequalities above are still valid in terms of expectation values when one explicitly refers to $n$-particle states $\Psi\in{{\cal H}}^{(n)}\cap{\mathfrak{S}}_{0}=\mathscr{S}(\mathsf{V}^{n}_{m,+})$, as already noted and used in [Ter14, Mor23] – only for $n=1$ – in more elementary versions of the result below.
Proposition 4.7:
Consider a real scalar Klein-Gordon field on $\mathbb{M}$} with mass $m>0$ and the associated normally ordered stress-energy tensor operator $\thinspace\hat{T}^{\mu\nu}[f]\thinspace:$ of Proposition 4.2, whose components are referred to a given Minkowski coordinate system. Consider a closed subspace ${{\cal K}}:=\oplus_{j\in J}{{\cal H}}^{(n_{j})}$ with $|n_{j}-n_{j^{\prime}}|\neq 0,2$ for $j\neq j^{\prime}$ – in particular ${{\cal K}}:={{\cal H}}^{(n)}$ for some $n\in{\mathbb{N}}$. If $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$ satisfies $f\geq 0$ and $\Psi\in{{\cal K}}\cap{\mathfrak{S}}_{0}$, then
$u_{\mu}u^{\prime}_{\nu}\langle\Psi|:\thinspace\hat{T}^{\mu\nu}\thinspace:[f]\Psi\rangle\geq 0,\quad\mbox{for every $u,u^{\prime}\in\overline{\mathsf{V}_{+}}$.}$ (49)
and thus $u_{\mu}\langle\Psi|:\thinspace\hat{T}^{\mu\nu}\thinspace:[f]\Psi\rangle$ is causal and future-directed, or vanishes, if $u\in\overline{\mathsf{V}_{+}}$.
The properties above are also valid if one replaces $\thinspace\hat{T}^{\mu\nu}\thinspace:[f]$ by $\thinspace\hat{T}^{\mu\nu}\thinspace:(x)$ with $x\in{\mathbb{M}}$.
Proof.
According to the structure of the stress-energy operator as presented in the proof of Proposition 4.2, $\langle\Psi|:\thinspace\hat{T}^{\mu\nu}\thinspace:[f]\Psi^{\prime}\rangle=0$ if $\Psi\in{{\cal H}}^{(n)}$, $\Psi^{\prime}\in{{\cal H}}^{(n^{\prime})}$ with $n\neq n^{\prime}$ and $|n-n^{\prime}|\neq 2$. So that, under our hypotheses, $u_{\mu}u^{\prime}_{\nu}\langle\Psi|:\thinspace\hat{T}^{\mu\nu}\thinspace:[f]\Psi^{\prime}\rangle=\sum_{n}u_{\mu}u^{\prime}_{\nu}\langle\Psi_{n}|:\thinspace\hat{T}^{\mu\nu}\thinspace:[f]\Psi_{n}\rangle$. It is therefore sufficient to prove the thesis for ${{\cal K}}:={{\cal H}}^{(n)}$. The case $n=0$ is obvious. If $\Psi\in{{\cal H}}^{(n)}\cap{\mathfrak{S}}_{0}$, $n=1,2,\ldots$, define, where $Q:=(q_{1},\ldots,q_{n-1})\in\mathsf{V}_{m,+}^{n-1}$,
$\psi(x,Q):=\frac{1}{\sqrt{2}(2\pi)^{3/2}}\int_{\mathsf{V}_{m,+}}\Psi(p,Q)e^{-ip\cdot x}d\mu_{m}(p)\>.$ (50)
By direct inspection, taking advantage of (39) and (36), where only the last two addends in the expansion matter, one sees that, if $\Psi,\Psi^{\prime}\in{{\cal H}}^{(n)}\cap{\mathfrak{S}}_{0}$ and $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$, then
$\langle\Psi|:\thinspace\hat{T}_{\mu\nu}[f]\thinspace:\Psi^{\prime}\rangle=\frac{n}{2}\int_{{\mathbb{M}}\times\mathsf{V}_{m,+}^{n-1}}\thinspace\thinspace\thinspace f(x)\left[\partial_{\mu}\overline{\psi}\partial_{\nu}\psi^{\prime}+\partial_{\nu}\overline{\psi}\partial_{\mu}\psi^{\prime}-g_{\mu\nu}\left(\partial_{\alpha}\overline{\psi}\partial^{\alpha}\psi^{\prime}+m^{2}\overline{\psi}\psi^{\prime}\right)\right]d^{4}xdQ\>,$
where $dQ:=d^{n-1}\mu_{m}(Q)$. Now assume $\Psi^{\prime}=\Psi$. Using $f\geq 0$,
$\langle\Psi|:\thinspace\hat{T}_{00}\thinspace:[f]\Psi\rangle$ $=\frac{n}{2}\int f(x)\left(\sum_{\mu=0}^{3}\partial_{\mu}\overline{\psi}\partial_{\mu}\psi+m^{2}\overline{\psi}\psi\right)d^{4}xdQ\geq 0\>,$
$\langle\Psi|:\thinspace\hat{T}_{00}\thinspace:[f]\Psi\rangle-\langle\Psi|:\thinspace\hat{T}_{11}\thinspace:[f]\Psi\rangle$ $=n\int f(x)\left(\sum_{\mu=2}^{3}\partial_{\mu}\overline{\psi}\partial_{\mu}\psi+m^{2}\overline{\psi}\psi\right)d^{4}xdQ\geq 0\>,$
$\langle\Psi|:\thinspace\hat{T}_{00}\thinspace:[f]\Psi\rangle+\langle\Psi|:\thinspace\hat{T}_{11}\thinspace:[f]\Psi\rangle$ $=n\int f(x)\sum_{\mu=0}^{1}\partial_{\mu}\overline{\psi}\partial_{\mu}\psi d^{4}xdQ\geq 0\>.$
so that $\langle\Psi|:\thinspace\hat{T}_{00}\thinspace:[f]\Psi\rangle\geq|\langle\Psi|:\thinspace\hat{T}_{11}\thinspace:[f]\Psi\rangle|$. This bound is obviously valid in every Minkowskian coordinate system. If $u,u^{\prime}\in\mathsf{V}_{+}$, we can choose a Minkowskian coordinate system $x^{\prime 0},x^{\prime 1},x^{\prime 2},x^{\prime 3}$ such that $u=c\partial^{\prime}_{0}+s\partial^{\prime}_{1}$ and $u^{\prime}=c\partial^{\prime}_{0}-s\partial^{\prime}_{1}$, where $s:=\sinh\chi$, $c:=\cosh\chi$ for some $\chi\in{\mathbb{R}}$. Since $\sinh^{2}\chi\leq\cosh^{2}\chi$ and the two bounds above hold, we have $u_{\mu}u^{\prime}_{\nu}\langle\Psi|:\thinspace\hat{T}^{\mu\nu}\thinspace:[f]\Psi\rangle=c^{2}\langle\Psi|:\thinspace\hat{T}^{\prime}_{00}\thinspace:[f]\Psi\rangle-s^{2}\langle\Psi|:\thinspace\hat{T}^{\prime}_{11}\thinspace:[f]\Psi\rangle\geq 0$ (when $f\geq 0$). This result concludes the proof, since the case $u,u^{\prime}\in\overline{\mathsf{V}_{+}}$ is obtained by linearity and continuity; the penultimate statement immediately follows from Proposition 4.6, and the last one is a direct consequence of (39) together with the smoothness of ${\mathbb{M}}\ni x\mapsto\langle\Psi|:\thinspace\hat{T}^{\mu\nu}\thinspace:(x)\Psi\rangle$. ∎
Let us pass to the case $\Psi\in{\mathfrak{S}}_{0}$, where arbitrary superpositions of components with different particle numbers are allowed. Now the proof above fails because in the expansion (36) the first two addends do not vanish in general. Actually, as is in particular suggested in [Few12], this is a general fact which directly follows from the already mentioned Reeh-Schlieder property (d) of Proposition 3.5. We prove this rigorously. Technically this is not completely easy, since it is not known whether all the symmetric operators considered are essentially selfadjoint.
Proposition 4.8:
Consider the *-algebra ${{\cal T}}({\cal O})$ of operators with common invariant dense domain ${\mathfrak{S}}_{0}$, smeared with test functions supported in $\cal O$} where ${\cal O}\subset{\mathbb{M}}$ is a bounded open set defined as in Proposition 4.5. If $A\in{{\cal T}}({\cal O})$ is symmetric, $A\geq 0$, and $\langle\Omega|A\Omega\rangle=0$, then $A=0$.
In particular, $u_{\mu}u^{\prime}_{\nu}:\thinspace\hat{T}^{\mu\nu}\thinspace:\thinspace[f]\geq 0$ and $\thinspace\hat{\phi}^{2}\thinspace:\thinspace[f]\geq 0$ cannot hold for $u,u^{\prime}\in\mathsf{V}$ and $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$ even if $f\geq 0$.
Proof.
[…]
We shall confine our investigation to Minkowski spacetime in the representation of the Poincaré-invariant quasifree state, and we study the quantum version of condition (49), since we shall prove a theorem of crucial relevance for constructing our localization POVMs, which will later be refined in Theorem 6.2 by making the form of the quadratic form $b^{\mu\nu}_{f}$ even more precise through a suitable choice of the smearing function $f$. The results we shall present rely on important known achievements [FeSm08] (and see [San24] for recent extension). In spite of our restriction to flat spacetime and to a special reference vacuum state, the results presented in Proposition 4.10 below should be extendable to Hadamard quasifree states in curved spacetime. We shall take advantage of a special case of the main result achieved in [FeSm08], where the opposite convention to ours concerning the metric signature is adopted. We start from the observation that a normalized vector $\Psi\in{\mathfrak{S}}_{0}$ defines an algebraic state of Hadamard type [FeSm08]. Indeed, the Minkowski vacuum $\Omega$ is Hadamard, as is well known. Every normalized $\Psi\in{\mathfrak{S}}_{0}$ is such that the difference $\Lambda_{\Psi}-\Lambda_{\Omega}$ of the associated two-point functions defines a smooth integral kernel according to Proposition 4.4. Hence $\Psi$ is Hadamard as well, by definition. Take $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$, $\Psi\in{\mathfrak{S}}_{0}$ with $\Psi||=1$, and $Q=A^{\mu}(x)\partial_{x^{\mu}}+B(x)$, an (at most) first-order differential operator with smooth real coefficients. Following the procedure outlined in the proof of Theorem 3.1 of [FeSm08] – which is more generally valid in suitably shaped domains of globally hyperbolic spacetimes and for Hadamard states – taking into account the comments in Section 3 of [Few05], the result of the quoted theorem simplifies to [^9] I am grateful to C. Fewster for clarifying this point to me.
$\int_{\mathbb{M}}f(x)^{2}Q\otimes Q|_{x=y}(\Lambda_{\Psi}(x,y)-\Lambda_{\Omega}(x,y))d^{4}x\geq-b_{Q}(f,f)\>,$ (51)
where, referring to a Minkowskian coordinate system where $p\equiv(p^{0},\vec{p})\in{\mathbb{R}}\times{\mathbb{R}}^{3}$,
$b_{Q}(h,h^{\prime}):=\frac{2}{(2\pi)^{4}}\int_{{\mathbb{R}}_{+}\times{\mathbb{R}}^{3}}\widehat{[(h\otimes h^{\prime})(Q\otimes Q)\Lambda_{\Omega}]}(-p,p)d^{4}p\>,\quad h,h^{\prime}\in\mathscr{D}({\mathbb{M}})$ (52)
the big hat above denoting the Fourier transform in ${\mathbb{R}}^{4}\times{\mathbb{R}}^{4}$. The crucial fact in (51) is that the lower bound found does not depend on the normalized vector $\Psi\in{\mathfrak{S}}_{0}$.
Remark 4.9:
It is interesting to observe that it is not possible to replace the smearing function $f^{2}$ with a non-negative $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$ in (51) by inserting $\sqrt{f}$ in place of $f$ on the right-hand side of (51). That is because $\sqrt{f}\not\in\mathscr{D}({\mathbb{M}})$ for $0\leq f\in\mathscr{D}({\mathbb{M}})$ in general [^10] Not even if $f$ vanishes with all derivatives at its zeros, as implied by a classic counterexample [Gl63].. Proposition 4.10 below is in particular different from Proposition 4.7 precisely because the former adopts a stronger hypothesis on the choice of the positive smearing functions. On the other hand, the main result of [FeSm08] presented above can easily be extended to the case in which $f^{2}$ is replaced by a finite sum $\sum_{i}f_{i}^{2}$ with $f_{i}\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$. This weaker requirement on the smearing functions was used in [San12] for proving (essential)selfadjointenss results about second-order normally ordered Wick polynomials in curved spacetime referred to Hadamard states. However, as discussed therein, there are $f\geq 0$ with $f\in\mathscr{D}({\mathbb{M}})$ which are not of the form $\sum_{i}f_{i}^{2}$. $\blacksquare$
We now come to the announced theorem. The first part of the following proposition concretely shows how some of the states that violate the positivity condition are constructed. They must exist according to the above Proposition 4.8.
Proposition 4.10:
Consider a real scalar Klein-Gordon field on $\mathbb{M}$} with mass $m>0$ and the associated normally ordered stress-energy tensor operator $\thinspace\hat{T}^{\mu\nu}[f]\thinspace:$ of Proposition 4.2, whose components are referred to a given Minkowski coordinate system.
  • (1)
    If $u,u^{\prime}\in\overline{\mathsf{V}_{+}}$, $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$ (in particular $f\geq 0$), there are vectors $\Psi\in{\mathfrak{S}}_{0}$ which are linear combinations of elements in ${{\cal H}}^{(n)}$ and ${{\cal H}}^{(n+2)}$ for $n=0,1,\ldots$ (in particular $\Psi\in{{\cal H}}^{(0)\perp}$) such that
    $u_{\mu}u^{\prime}_{\nu}\langle\Psi|:\thinspace\hat{T}^{\mu\nu}[f]\thinspace:\Psi\rangle<0\>.$
  • (2)
    There exist $\mu,\nu$-symmetric bilinear maps $b^{\mu\nu}:\mathscr{D}({\mathbb{M}})\times\mathscr{D}({\mathbb{M}})\to{\mathbb{C}}$ such that the following facts are valid, where $b_{f}^{\mu\nu}:=b^{\mu\nu}(f,f)$ for $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$.
    • (a)
      If $u,u^{\prime}\in\overline{\mathsf{V}_{+}}$, $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$,
      $u_{\mu}u^{\prime}_{\nu}\langle\Psi|:\thinspace\hat{T}^{\mu\nu}\thinspace:[f^{2}]\Psi\rangle\geq-u_{\mu}u^{\prime}_{\nu}b_{f}^{\mu\nu}||\Psi||^{2}\>,\quad\mbox{for every $\Psi\in{\mathfrak{S}}_{0}$,}$ (53)
      where necessarily $u_{\mu}u^{\prime}_{\nu}b_{f}^{\mu\nu}>0$.
    • (b)
      If $u\in{\mathsf{T}}_{+}$, $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$, $\Psi\in{\mathfrak{S}}_{0}$, the vector
      $u_{\mu}\left(\langle\Psi|:\thinspace\hat{T}^{\mu\nu}\thinspace:[f^{2}]\Psi\rangle+b_{f}^{\mu\nu}||\Psi||^{2}\right)$
      is causal and future-directed, or vanishes.
    The results in (a) and (b) are also valid if $f^{2}$ is replaced by $\sum_{i=1}^{N}f_{i}^{2}$ with $f_{i}\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$, with the obvious redefinition of $b_{f}^{\mu\nu}$.
Proof.
(1) Referring to the structure presented in the proof of Proposition 4.2, consider a vector of the form $\epsilon\Psi_{n+2}+\sqrt{1-\epsilon^{2}}\Psi_{n}$, where $\epsilon\in(0,1)$ and $\Psi_{n+2}\in{{\cal H}}^{(n+2)}\cap{\mathfrak{S}}_{0}$, $\Psi_{n}\in{{\cal H}}^{(n)}\cap{\mathfrak{S}}_{0}$ are normalized. It is easy to prove that, given $f$ and $u,u^{\prime}$, it is possible to choose $\Psi_{n+2}$ in such a way that $u_{\mu}u^{\prime}_{\nu}\langle\Psi_{n}|:\thinspace\hat{T}^{\mu\nu}[f]_{aa}\thinspace:\Psi_{n+2}\rangle\neq 0$. Changing $\Psi_{n+2}$ by a phase factor if necessary, we also have $u_{\mu}u^{\prime}_{\nu}Re(\langle\Psi_{n}|:\thinspace\hat{T}^{\mu\nu}[f]_{aa}\thinspace:\Psi_{n+2}\rangle)<0$. It holds
$u_{\mu}u^{\prime}_{\nu}\langle\Psi|:\thinspace\hat{T}^{\mu\nu}[f]\thinspace:\Psi\rangle=2\epsilon^{2}u_{\mu}u^{\prime}_{\nu}(\langle\Psi_{n+2}|:\thinspace\hat{T}^{\mu\nu}[f]^{+}_{a^{\dagger}a}\thinspace:\Psi_{n+2}\rangle+\langle\Psi_{n}|:\thinspace\hat{T}^{\mu\nu}[f]^{+}_{a^{\dagger}a}\thinspace:\Psi_{n}\rangle)$
$2\epsilon\sqrt{1-\epsilon^{2}}u_{\mu}u^{\prime}_{\nu}Re(\langle\Psi_{n}|:\thinspace\hat{T}^{\mu\nu}[f]_{aa}\thinspace:\Psi_{n+2}\rangle)$
where $u_{\mu}u^{\prime}_{\nu}\langle\Psi_{n}|:\thinspace\hat{T}^{\mu\nu}[f]^{+}_{a^{\dagger}a}\thinspace:\Psi_{n}\rangle\>,u_{\mu}u^{\prime}_{\nu}\langle\Psi_{n+2}|:\thinspace\hat{T}^{\mu\nu}[f]^{+}_{a^{\dagger}a}\thinspace:\Psi_{n+2}\rangle\geq 0$ if $f\geq 0$ (in the other case the proof is already complete). In this case, choosing $\epsilon>0$ sufficiently small, the sign of $u_{\mu}u^{\prime}_{\nu}\langle\Psi|:\thinspace\hat{T}^{\mu\nu}[f]\thinspace:\Psi\rangle$ is nevertheless strictly negative according to the expression above.
(2) Item (b) is a direct consequence of (a) and Proposition 4.6. The idea of the proof of (a) is to show that $u^{\mu}u^{\prime\nu}D_{\mu\nu}$ – where $D_{\mu\nu}$ is defined in (47) – can be written as a sum of operators $\sum_{a=1}^{N}Q^{\prime}_{a}\otimes Q^{\prime}_{a}$ – where the operators may depend on $u,u^{\prime}$ – so that one can apply (51) and (52) to (46). It is sufficient to prove the thesis for $u,u^{\prime}\in\mathsf{V}_{+}$, since the case of lightlike vectors is obtained by an obvious limit. We have already chosen a preferred Minkowskian coordinate system $x^{0},x^{1},x^{2},x^{3}$ in which to compute (52) and in which $D_{\mu\nu}$ is defined. If $u,u^{\prime}\in\mathsf{V}_{+}$, we can choose an auxiliary Minkowskian coordinate system $x^{\prime 0},x^{\prime 1},x^{\prime 2},x^{\prime 3}$, depending on $u,u^{\prime}$, such that $x^{\prime 2}=x^{2}$, $x^{\prime 3}=x^{3}$ and $u=c\partial^{\prime}_{0}+s\partial^{\prime}_{1}$ and $u^{\prime}=c\partial^{\prime}_{0}-s\partial^{\prime}_{1}$, where $s:=\sinh\chi$, $c:=\cosh\chi$ for some $\chi\in{\mathbb{R}}$. We define the first-order differential operators $Q^{\prime}_{\mu}:=\partial^{\prime}_{\mu}$ and $Q^{\prime}_{4}:=mI$. Then
$u_{\mu}u^{\prime}_{\nu}\langle\Psi|:\thinspace\hat{T}^{\mu\nu}\thinspace:[f^{2}]\Psi\rangle=c^{2}\langle\Psi|:\thinspace\hat{T}^{\prime 00}\thinspace:[f^{2}]\Psi\rangle-s^{2}\langle\Psi|:\thinspace\hat{T}^{\prime 11}\thinspace:[f^{2}]\Psi\rangle$
On the other hand, by direct inspection one sees that $u^{\mu}u^{\prime\nu}D_{\mu\nu}=c^{2}D^{\prime}_{00}-s^{2}D^{\prime}_{11}$
$\frac{c^{2}-s^{2}}{2}(Q^{\prime}_{0}\otimes Q^{\prime}_{0}+Q^{\prime}_{1}\otimes Q^{\prime}_{1})+\frac{c^{2}+s^{2}}{2}(Q^{\prime}_{2}\otimes Q^{\prime}_{2}+Q^{\prime}_{3}\otimes Q^{\prime}_{3}+Q^{\prime}_{4}\otimes Q^{\prime}_{4})$
Since both coefficients are non-negative, we can absorb their square roots directly into the definition of the operators $Q^{\prime}_{a}$, and we can apply (51), obtaining
$u_{\mu}u^{\prime}_{\nu}\langle\Psi|:\thinspace\hat{T}^{\mu\nu}\thinspace:[f^{2}]\Psi\rangle\geq-u_{\mu}u^{\prime}_{\nu}b^{\mu\nu}(f,f)\>,$
where, writing the operators in the initial coordinate system and also taking advantage of the linearity of the Fourier transform,
$u_{\mu}u^{\prime}_{\nu}b^{\mu\nu}(h,h^{\prime}):=u_{\mu}u^{\prime}_{\nu}\int_{{\mathbb{R}}_{+}\times{\mathbb{R}}^{3}}\frac{2\widehat{[h\otimes h^{\prime}D^{\mu\nu}\Lambda_{\Omega}]}(-p,p)}{(2\pi)^{4}}d^{4}p.$
Since (53) is valid for all $\Psi\in{\mathfrak{S}}_{0}$, necessarily $u_{\mu}u^{\prime}_{\nu}b^{\mu\nu}(f,f)>0$, otherwise (1) would be false. This ends the proof, since the last statement has an obvious proof. ∎
4.4 Integration of $v^{\mu}:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]$ over rest spaces $\Sigma$ and interplay with $H_{v}$
Our intention is to define the four-momentum operator associated with the whole rest space $\Sigma$, obtained by integrating over $\Sigma$ the stress-energy tensor operator smeared with $f\in\mathscr{D}({\mathbb{M}})$. This notion will be generalized to subregions $\Delta\subset\Sigma$ in Section 6.1. We start by defining a relevant density to be integrated. Take a smearing function $f\in\mathscr{D}({\mathbb{M}})$ and define the unitary strongly continuous representation of the translations $\mathsf{V}$ of Minkowski spacetime, $\mathsf{V}\ni v\mapsto U_{(I,v)}\in{\mathfrak{B}}({\mathfrak{F}}_{c}({{\cal H}}_{m}))$, the latter being the unitary representation (9) of $IO(1,3)_{+}$ on the Fock space and, in all what follows, we have identified the points of $\mathbb{M}$} with the vectors of $\mathsf{V}$ as usual $\mathsf{V}\ni x\mapsto o+x\in{\mathbb{M}}$ through the choice of an origin $o\in{\mathbb{M}}$. What follows does not depend on this choice. Following some ideas in [BuFr82], we define the operator-valued function
$\mathsf{V}\ni x\mapsto U_{(I,x)}:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]U^{-1}_{(I,x)}=:\>:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]\quad\mbox{with}\quad f_{x}(y):=f(y-x)\quad\mbox{for $x,y\in\mathsf{V}$,}$ (54)
according to (d) of Proposition 4.2. This $x$-dependent operator $\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]$ enjoys a number of crucial properties.
Proposition 4.11:
Let $f\in\mathscr{D}({\mathbb{M}})$ and let $\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]$ be defined as in (54). The following facts are valid if $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$.
  • (a)
    The identity holds
    $\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]\Psi^{\prime}\rangle=\int_{\mathbb{M}}f(y-x)\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:(y)\Psi^{\prime}\rangle d^{4}y\>.$
  • (b)
    $\mathsf{V}\ni x\mapsto\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]\Psi^{\prime}\rangle$ is smooth, bounded, and satisfies the following bounds referred to a given Minkowski frame where $x=(x^{0},\vec{x})$.
    (1) If $\alpha$ is a multi-index for the components of $x$, for every $n\in{\mathbb{N}}$ there is a polynomial $P_{|\alpha|+n}$ in the variable $x^{0}$ such that
    $\left|\partial^{\alpha}_{x}\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]|\Psi^{\prime}\rangle\right|\leq\frac{|P_{|\alpha|+n}(x^{0})|}{(1+|\vec{x}|)^{n}}\>,\quad x\in{\mathbb{R}}^{4}\>.$ (55)
    In particular, ${\mathbb{R}}^{3}\ni\vec{x}\mapsto\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{(x^{0},\vec{x})}]|\Psi^{\prime}\rangle$ is a Schwartz function for every $x^{0}\in{\mathbb{R}}$.
    (2) There are finite constants $\epsilon>0$, $C>0$ such that
    $|\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]|\Psi^{\prime}\rangle|\leq\frac{C}{(1+|\vec{x}|)^{3+\epsilon}}\quad\mbox{if $|\vec{x}|>|x^{0}|$.}$
  • (c)
    The conservation equation holds (the derivatives being referred to coordinates of $x$)
    $\partial_{\mu}\langle\Psi|:\thinspace\hat{T}^{\mu\nu}\thinspace:[f_{x}]\Psi^{\prime}\rangle=0\>.$ (56)
  • (d)
    If $v\in\mathsf{V}$, $f\in\mathscr{D}({\mathbb{M}})$, and $\Sigma$ and $\Sigma^{\prime}$ are rest spaces of two Minkowskian reference frames $u_{\Sigma},u_{\Sigma^{\prime}}\in{\mathsf{T}}_{+}$, then
    $\int_{\Sigma}v^{\mu}\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]\Psi^{\prime}\rangle u^{\nu}_{\Sigma}d\Sigma(x)=\int_{\Sigma^{\prime}}v^{\mu}\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]\Psi^{\prime}\rangle u^{\nu}_{\Sigma^{\prime}}d\Sigma^{\prime}(x)\>.$ (57)
    Above, $d\Sigma,d\Sigma^{\prime}$ are the ${\mathbb{R}}^{3}$ Lebesgue measure in spatial Minkowski coordinates adapted to the $3$-planes $\Sigma$ and $\Sigma^{\prime}$, respectively.
  • (e)
    The bound (53) holds uniformly in $x\in{\mathbb{M}}$:
    $u_{\mu}u^{\prime}_{\nu}\langle\Psi|:\thinspace\hat{T}^{\mu\nu}\thinspace:[f_{x}^{2}]\Psi\rangle\geq-u_{\mu}u^{\prime}_{\nu}b_{f}^{\mu\nu}||\Psi||^{2}\>,$ (58)
    for $u,u^{\prime}\in\overline{\mathsf{V}_{+}}$, and $\Psi\in{\mathfrak{S}}_{0}$, where $b_{f}^{\mu\nu}$ is defined in Proposition 4.10.
Proof.
See Appendix B. ∎
We move on to the interplay between the generator $H^{v}$ of spacetime translations along $v\in\mathsf{V}$, defined in Lemma 3.1, and integrals of $\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]$ over rest spaces of Minkowskian reference frames.
Proposition 4.12:
If $v\in\mathsf{V}$, $f\in\mathscr{D}({\mathbb{M}})$, $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$, independently of the choice of the rest space $\Sigma$ of a Minkowskian reference frame $n_{\Sigma}$,
$\int_{\Sigma}\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]\Psi^{\prime}\rangle v^{\mu}u^{\nu}_{\Sigma}\>\>d\Sigma(x)=\left(\int_{\mathbb{M}}fd^{4}x\right)\langle\Psi|H^{v}\Psi^{\prime}\rangle\>,$ (59)
so that, in particular, the left-hand side is positive if $\Psi=\Psi^{\prime}$, $f\geq 0$ and $v\in\mathsf{V}_{+}$.
Proof.
Fix a Minkowskian coordinate system with $\partial_{0}=u_{\Sigma}$ so that $\Sigma$ coincides with the plane $x^{0}=0$. Integrating in $\vec{x}$ over the whole space ${\mathbb{R}}^{3}$ the integrals in (118), with a standard argument based on a sequence of regular distributions weakly tending to $\delta(\vec{p}-\vec{k})$, produces $\int_{\Sigma}v^{\mu}\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]\Psi^{\prime}\rangle u^{\nu}_{\Sigma}d\Sigma(x)=$
$2\pi^{2}\int_{{\mathbb{R}}^{3}}\hat{f}(-2p)\frac{\langle\Psi|a_{p}a_{p}\Psi^{\prime}\rangle}{E(p)^{2}}v^{\mu}t_{\mu 0}(p,-\tilde{p})d^{3}p-2\pi^{2}\int_{{\mathbb{R}}^{3}}\hat{f}(2p)\frac{\langle\Psi|a^{\dagger}_{p}a^{\dagger}_{p}\Psi^{\prime}\rangle}{E(p)^{2}}v^{\mu}t_{\mu 0}(p,-\tilde{p})d^{3}p$
$2\pi^{2}\int_{{\mathbb{R}}^{3}}\hat{f}(0)\frac{\langle\Psi|a^{\dagger}_{p}a_{p}\Psi^{\prime}\rangle}{E(p)^{2}}v^{\mu}t_{\mu 0}(p,p)d^{3}p+2\pi^{2}\int_{{\mathbb{R}}^{6}}\hat{f}(0)\frac{\langle\Psi|a^{\dagger}_{p}a_{p}\Psi^{\prime}\rangle}{E(p)^{2}}v^{\mu}t_{\mu 0}(p,p)d^{3}p\>$
where $\tilde{p}:=(E(p),-\vec{p})$. At this point, direct inspection proves that $t_{\mu 0}(-p,\tilde{p})=t_{\mu 0}(p,-\tilde{p})=0$, whereas $\frac{v^{\mu}t_{\mu 0}(p,p)}{E(p)}=-v\cdot p$, so that
$\int_{\Sigma}v^{\mu}\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]\Psi^{\prime}\rangle u^{\nu}_{\Sigma}d\Sigma(x)=-(2\pi)^{2}\hat{f}(0)\int_{\mathsf{V}_{m,+}}\langle\Psi|a^{\dagger}_{p}a_{p}\Psi^{\prime}\rangle v\cdot p\>d\mu_{m}(p)\>.$
Since $(2\pi)^{2}\hat{f}(0)=\int f(x)d^{4}x$ and taking (109) into account, the right-hand side of the identity found is $(\int fd^{4}x)\langle\Psi|H^{v}\Psi^{\prime}\rangle$. ∎
A surprising fact is that – assuming $\int_{\mathbb{M}}fd^{4}x=1$ – the right-hand side of (59) is independent of the smearing function $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$, which however appears in the left-hand side. We stress that this is not an evident result, since it proves in particular that, in the left-hand side, two integrals can be interchanged, but one is over the whole spacetime, namely ${\mathbb{R}}^{4}$, and the other over a flat Cauchy surface, i.e. ${\mathbb{R}}^{3}$: it is not a direct application of Fubini’s theorem. It is easy to see, at a heuristic level, that this is due to elementary properties of the smearing procedure and the conservation property (57) when $\Sigma$ and $\Sigma^{\prime}$ are parallel. We expect that (59) (or its formulation in terms of quadratic forms) is also valid in curved spacetime in the presence of a Killing vector field $v$. According to (59) it could be convenient to introduce the notation, formally motivated by the replacement $f(y-x)\to\delta(y-x)$,
$H^{v}=\int_{\Sigma}:\thinspace\hat{T}_{\mu\nu}\thinspace:(x)v^{\mu}u^{\nu}_{\Sigma}\>\>d\Sigma(x)$
which is used in theoretical physics textbooks.
5 Relativistic spatial localization observable from the stress-energy operator
We are now in a position to construct a relativistic spatial localization observable for $n$-particle states by integrating a normalized notion of $\thinspace T_{\mu\nu}\thinspace:[f^{2}_{x}]$ over Borel sets $\Delta$ of rest spaces $\Sigma$ of Minkowskian reference frames $u_{\Sigma}$. The construction we are going to present proves that the relativistic spatial localization observables introduced in [Ter14, Mor23] are actually rigorously constructed out of QFT notions: they are restrictions of a more general structure, as conjectured in the conclusions of [CDM26].
5.1 $n$-particle relativistic spatial localization from $\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}]$ satisfying CC
Due to (59), if $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$, $\int_{\mathbb{M}}fd^{4}x=1$, $f\geq 0$, and $u\in{\mathsf{T}}_{+}$, extending a definition given in [Mor23], an expected expression for the desired effects should be
${\mathsf{A}}_{f}^{u}(\Delta)=\frac{1}{\sqrt{H^{u}}}\int_{\Delta}:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]u^{\mu}u^{\nu}_{\Sigma}\>\>d\Sigma(x)\frac{1}{\sqrt{H^{u}}}\>,\quad\Delta\in\mathscr{B}(\Sigma)\quad\mbox{(tentatively!)}$ (60)
where $\mathscr{B}(\Sigma)$ denotes the $\sigma$-algebra of Borel sets in the rest space $\Sigma$ of the Minkowskian reference frame $u_{\Sigma}$. This type of expression was used in [Mor23], and already in [Ter14] for a first version of this localization notion with $u=u_{\Sigma}$. However, the notion of stress-energy tensor operator was introduced in those works only in a heuristic manner, without analyzing the crucial smearing procedure, which makes it a mathematically sound object in rigorous QFT and also allows it to comply with the localization notion of the AHK approach. As a matter of fact, the rigorous version of the constructed POVM was actually defined in [Mor23] in terms of a suitable improvement of Newton-Wigner’s PVM and only in the one-particle space. This type of relation with the Newton-Wigner observable will arise later in a more general form, but we shall not use it to define our notion of localization in our QFT context, contrary to [Mor23].
Evident physical issues with (60) arise immediately in our QFT context. First of all, ${\mathsf{A}}_{f}^{u}(\Delta)$ defined as above is not necessarily positive, since $\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]$ is not positive, not even if we replace $f$ by $f^{2}$! However, some lower bounds hold, as shown in Proposition 4.10 when using $f^{2}$ with $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$. Furthermore, due to Proposition 4.7, if we consider the compressions to $n$-particle spaces ${{\cal H}}^{(n)}$, namely the operators given by $P_{n}:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]P_{n}$ and $P_{n}A_{f}^{u}(\Delta)P_{n}$, they are positive, and the latter is also normalized to $I$ for $\Delta=\Sigma$ because of (59) ($P_{n}{\mathfrak{F}}_{s}({{\cal H}}_{m})\to{\mathfrak{F}}_{s}({{\cal H}}_{m})$ denotes the orthogonal projector onto ${{\cal H}}^{(n)}$). We are particularly interested in the specific case of one particle: $n=1$. From the causality properties proved in Proposition 4.7 and (c) in Proposition 4.11, exploiting a procedure already used in [Mor23, DRM24, CDM26], we expect that the causality condition (CC) in Definition 2.2 is satisfied when (a compression of) the operators ${\mathsf{A}}_{f}^{u}(\Delta)$ is used to define families of POVMs on all rest spaces of $\mathbb{M}$}. Finally, notice that in (60) $\frac{1}{\sqrt{H^{u}}}\Omega$ is not defined. To fix this problem we consider the square root of the resolvent $H^{-1}_{u,\epsilon}$, where henceforth we define the $\epsilon$-regularized Hamiltonian
$H^{u}_{\epsilon}:=(H^{u}+\epsilon I)\quad\mbox{for $\epsilon>0$ (arbitrarily small)}\>,$ (61)
and later we shall consider the limit as $\epsilon\to 0^{+}$. We notice en passant that the $\alpha$-th power of the written operator, defined via functional calculus, satisfies
$(H^{u}_{\epsilon})^{\alpha}({\mathfrak{S}}_{0})={\mathfrak{S}}_{0}\quad\mbox{for $\alpha\in{\mathbb{R}}$ and $\epsilon>0$.}$ (62)
In fact, in every reducing space ${{\cal H}}^{(n)}$ the operator $H^{u}_{\epsilon}$ restricts to a multiplicative operator with a strictly positive, smooth polynomially bounded function. The spectrally defined $\alpha$-power of $H^{u}_{\epsilon}$ therefore coincides with the corresponding power of the said multiplicative operator (e.g. by (f) of Proposition 3.3 in [Mor19]) in each ${{\cal H}}^{(n)}$, giving rise to (62). We start from a technical preliminary result, where we explicitly use $f^{2}$ to smear the stress-energy tensor. Some of the results established below should remain valid even when using $f\geq 0$ in place of $f^{2}$. However, the proof of the proposition below is easier to carry out if one assumes that $\sqrt{f}$ is smooth, and this is not guaranteed by the requirement that $f\geq 0$ be smooth, as observed below the proof of Proposition 4.10.
Proposition 5.1:
Consider $u\in{\mathsf{T}}_{+}$, $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$, $\epsilon>0$ and, for a rest space $\Sigma$ of a Minkowskian reference frame $u_{\Sigma}\in{\mathsf{T}}_{+}$, define the algebra of sets
$\mathscr{B}_{0}(\Sigma):=\{\Delta\in\mathscr{B}(\Sigma)\>|\>\>\mbox{either}\>\>|\Delta|<+\infty\>\>\mbox{or}\>\>|\Sigma\setminus\Delta|<+\infty\}.$
For every $\Delta\in\mathscr{B}_{0}(\Sigma)$ there is an operator ${\mathsf{A}}_{f,\epsilon}^{u}(\Delta):{\mathfrak{F}}_{s}({{\cal H}}_{m})\to{\mathfrak{F}}_{s}({{\cal H}}_{m})$ which is uniquely defined by requiring that
$\left\langle\Psi\left|{\mathsf{A}}_{f,\epsilon}^{u}(\Delta)\right.\Psi^{\prime}\right\rangle=\int_{\Delta}\thinspace\left\langle\frac{1}{\sqrt{H^{u}_{\epsilon}}}\Psi\left|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]u^{\mu}u^{\nu}_{\Sigma}\frac{1}{\sqrt{H^{u}_{\epsilon}}}\Psi^{\prime}\right.\right\rangle d\Sigma(x)\>\>\mbox{if $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$.}$ (63)
(Where the right-hand side more generally exists for a generic $\Delta\in\mathscr{B}(\Sigma)$.) The following facts are true.
  • (a)
    ${\mathsf{A}}_{f,\epsilon}^{u}(\Delta)^{\dagger}={\mathsf{A}}_{f,\epsilon}^{u}(\Delta)\in{\mathfrak{B}}({\mathfrak{F}}_{s}({{\cal H}}_{m}))$.
  • (a1)
    There is $C_{f}>0$ such that $P_{n}{\mathsf{A}}_{f,\epsilon}^{u}(\Delta)P_{n}||\leq C_{f}$ for all $n\in{\mathbb{N}}$, $\epsilon>0$, $\Delta\in\mathscr{B}_{0}(\Sigma)$, where $P_{n}$ is the orthogonal projector onto ${{\cal H}}^{(n)}$.
  • (b)
    If $\Delta=\Sigma$,
    ${\mathsf{A}}^{u}_{f,\epsilon}(\Sigma)=\left(\int_{\mathbb{M}}f^{2}d^{4}x\right)\frac{H^{u}}{{H^{u}+\epsilon I}}\>.$ (64)
  • (c)
    If $g\in IO(1,3)_{+}$ then the covariance relation holds
    $U_{g}{\mathsf{A}}_{f,\epsilon}^{u}(\Delta)U_{g}^{-1}={\mathsf{A}}_{g_{*}f,\epsilon}^{gu}(g\Delta)$ (65)
    where $U$ is the unitary $IO(1,3)_{+}$-representation (9).
  • (d)
    The map $\mathscr{B}_{0}(\Sigma)\ni\Delta\mapsto{\mathsf{A}}_{f,\epsilon}^{u}(\Delta)$ is weakly $\sigma$-additive.
  • (e)
    The operators ${\mathsf{A}}_{f,\epsilon}^{u}(\Delta)$ are bounded from below, but not positive in general: for $x\in{\mathbb{M}}$ and $\Sigma\ni x$, there is a measurable $\Delta\subset\Sigma$ which contains $x$ and $\Psi\in{\mathfrak{S}}_{0}\cap{{\cal H}}^{(0)\perp}$ such that $\langle\Psi|{\mathsf{A}}_{f,\epsilon}^{u}(\Delta)\Psi\rangle<0$.
All the statements remain valid if one replaces $f^{2}$ by $f=\sum_{i=1}^{N}f_{i}^{2}$ with $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$.
Proof.
See Appendix B. ∎
We are now in a position to state and prove one of the main results of this work. We construct a positive-energy relativistic spatial localization observable (Definition 2.1) for states with a definite number of particles in the Fock space and which is causal in the sense that it complies with CC as in Definition 2.2. This family of POVMs uniquely arises from the stress-tensor operator smeared with a test function $f^{2}$ and the choice of a preferred temporal direction $u$. We cannot directly use an identity such as (60) to define an effect because of the states with “negative probability” that would arise from (e) in Proposition 5.1. Even if we shall return later to that issue, what we do now is remove these annoying states by passing to the compressions $P_{n}A_{f}^{u}(\Delta)P_{n}$, as already suggested.
As before, we use the following notation
${\cal R}:=\cup\{\mathscr{B}(\Sigma)\>|\>\Sigma\subset{\mathbb{M}}\>\mbox{spacelike $3$-plane}\}$ (66)
where $\mathscr{B}(\Sigma)$ is the Borel $\sigma$-algebra on $\Sigma$ and $u_{\Sigma}\in{\mathsf{T}}_{+}$ is the future-oriented unit normal vector to $\Sigma$.
Theorem 5.2:
Consider a real scalar Klein-Gordon field on $\mathbb{M}$} with mass $m>0$ and the associated normally ordered stress-energy tensor operator $\thinspace\hat{T}_{\mu\nu}[f]\thinspace:$ defined in the Fock space ${\mathfrak{F}}_{s}({{\cal H}}_{m})$ as in Proposition 4.2. Take $u\in{\mathsf{T}}_{+}$ with associated selfadjoint Hamiltonian $H^{u}$ (13), and $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$ such that $\int_{{\mathbb{M}}}f^{2}d^{4}x=1$. Then there is a unique map
${\cal R}\ni\Delta\mapsto{A}^{u}_{f}(\Delta)\in{\mathfrak{B}}({\mathfrak{F}}_{s}({{\cal H}}_{m}))$ (67)
such that
  • (a)
    $A^{u}_{f}(\Delta)({{\cal H}}^{(n)})\subset{{\cal H}}^{(n)}$ for $\Delta\in{\cal R}$ and $n\in{\mathbb{N}}$;
  • (b)
    ${A}^{u}_{f}(\Delta)({{\cal H}}^{(0)})=0$ for every $\Delta\in{\cal R}$;
  • (c)
    if $\rho\in{\mathsf{S}}({{\cal H}}^{(n)})$ for a given $n=0,1,2,\ldots$ and $\Delta\in\mathscr{B}_{0}(\Sigma)$, for a spacelike $3$-plane $\Sigma$,
    $tr(\rho{A}_{f}^{u}(\Delta))=\lim_{\epsilon\to 0^{+}}tr(\rho{\mathsf{A}}^{u}_{f,\epsilon}(\Delta))\>;$ (68)
  • (d)
[…]
The following further facts are true.
  • (e)
    If $\Delta\in{\cal R}$, $P_{n}:{\mathfrak{F}}_{s}({{\cal H}}_{m})\to{\mathfrak{F}}_{s}({{\cal H}}_{m})$ is the orthogonal projector onto ${{\cal H}}^{(n)}$, and $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$, then
    $\langle\Psi|{A}^{u}_{f}(\Delta)\Psi^{\prime}\rangle=\sum_{n=1}^{+\infty}\int_{\Delta}\left\langle\frac{1}{\sqrt{H^{u}}}P_{n}\Psi\left|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]\frac{1}{\sqrt{H^{u}}}\right.P_{n}\Psi^{\prime}\right\rangle u^{\mu}u^{\nu}_{\Sigma}d\Sigma(x)\>,$ (69)
    where $\frac{1}{\sqrt{H^{u}}}$ is defined by spectral calculus on ${{\cal H}}^{(0)\perp}$.
Furthermore, referring to Definition 2.1, if $U$ is the unitary $IO(1,3)_{+}$-representation (9) and $A_{f}$ denotes the family of maps $\{A_{f}^{u}\}_{u\in{\mathsf{T}}_{+}}$,
  • (f)
[…]
  • (g)
[…]
All the statements remain valid if one replaces $f^{2}$ by $f=\sum_{i=1}^{N}f_{i}^{2}$ with $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$.
Proof.
Referring to the proof of Proposition 5.1 (decomposition (120) in particular), define
${A}^{u}_{f}(\Delta):={\mathsf{A}}_{0,3}(\Delta)(=:{\mathsf{A}}_{\epsilon,3}(\Delta)|_{\epsilon=0})$ (70)
With this definition, (a) and (b) are true, in particular because $P_{n}{\mathsf{A}}_{0,3}(\Delta)={\mathsf{A}}_{0,3}(\Delta)P_{n}$ and $P_{0}{\mathsf{A}}_{0,3}(\Delta)={\mathsf{A}}_{0,3}(\Delta)P_{0}=0$. We stress that ${A}^{u}_{f}(\Delta)\in{\mathfrak{B}}({\mathfrak{F}}_{s}({{\cal H}}_{m}))$ and that this operator is defined for every $\Delta\in\mathscr{B}(\Sigma)$ of every spacelike $3$-plane, as observed in the proof of Proposition 5.1 (see the discussion below (119) and (125)). For future convenience we note that, if $\Psi\in{{\cal H}}^{(n)}\cap{\mathfrak{S}}_{0}$ with $n>0$, we have
$\langle\Psi|{A}^{u}_{f}(\Delta)\Psi\rangle=\langle\Psi|{\mathsf{A}}_{0,3}(\Delta)\Psi\rangle=\int_{\Delta}\thinspace\left\langle\frac{1}{\sqrt{H^{u}}}\Psi\left|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]\frac{1}{\sqrt{H^{u}}}\right.\Psi\right\rangle u^{\mu}u^{\nu}_{\Sigma}d\Sigma(x)\geq 0$ (71)
where $\frac{1}{\sqrt{H^{u}}}$ is defined by spectral calculus on ${{\cal H}}^{(0)\perp}$ and as the zero operator on ${{\cal H}}^{(0)}$, and where we have taken advantage of Proposition 4.7. At this juncture, using $P_{n}{\mathsf{A}}_{0,3}(\Delta)={\mathsf{A}}_{0,3}(\Delta)P_{n}=P_{n}{\mathsf{A}}_{0,3}(\Delta)P_{n}$, linearity and polarization immediately yield (e).
Let us pass to the proof of (c). Since both sides vanish separately if $n=0$, we consider the case $n>0$. First we focus on the more elementary case of $\rho^{\prime}=\sum_{i=1}^{N}p_{i}|\Psi_{i}\rangle\langle\Psi_{i}|$, for $\Psi_{i}\in{{\cal H}}^{(n)}$ with a given $n$. For this type of state it is sufficient to prove that, if $\Delta\in\mathscr{B}(\Sigma)$ (actually it would be sufficient to consider only the case $\Delta\in\mathscr{B}_{0}(\Sigma)$) and $\Psi,\Psi^{\prime}\in{{\cal H}}^{(n)}$, with $n>0$,
$\langle\Psi|{\mathsf{A}}_{\epsilon,3}(\Delta)\Psi^{\prime}\rangle=\int_{\Delta}\thinspace\left\langle\frac{1}{\sqrt{H^{u}+\epsilon I}}\Psi\left|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]\frac{1}{\sqrt{H^{u}+\epsilon I}}\right.\Psi^{\prime}\right\rangle u^{\mu}u^{\nu}_{\Sigma}d\Sigma(x)$
$\to\int_{\Delta}\thinspace\left\langle\frac{1}{\sqrt{H^{u}}}\Psi\left|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]\frac{1}{\sqrt{H^{u}}}\right.\Psi^{\prime}\right\rangle u^{\mu}u^{\nu}_{\Sigma}d\Sigma(x)=\langle\Psi|{\mathsf{A}}_{0,3}(\Delta)\Psi^{\prime}\rangle\quad\mbox{ if $\epsilon\to 0^{+}$. }$ (72)
Let us prove this. Indeed, choosing a Minkowskian coordinate system with $\Sigma$ defined by $x^{0}=0$ and looking at the expression of $\langle(H^{u}+\epsilon I)^{-1/2}\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]|(H^{u}+\epsilon I)^{-1/2}\Psi^{\prime}\rangle$, arising from the expansion (118), where only the last two integrals contribute for $\Psi^{\prime},\Psi\in{\mathfrak{S}}_{0}\cap{{\cal H}}^{(n)}$ with $n>0$, one easily sees that, for every $\vec{x}\in\Sigma$,
$\langle(H^{u}+\epsilon I)^{-1/2}\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]|(H^{u}+\epsilon I)^{-1/2}\Psi^{\prime}\rangle\to\langle(H^{u})^{-1/2}\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]|(H^{u})^{-1/2}\Psi^{\prime}\rangle$
as $\epsilon\to 0$. This is a direct consequence of the dominated convergence theorem, when one expands the scalar products as integrals on $\mathsf{V}_{m,+}^{n}$ using the fact that the function $\Psi$ is of Schwartz type (when viewed as a function of the spatial momenta $(\vec{p}_{1},\ldots,\vec{p}_{n})$), and that $(E^{n}_{\epsilon,u})^{-1/2}$ is a bounded function of $(\vec{p}_{1},\ldots,\vec{p}_{n})$. On the other hand, integration by parts proves that, for every $N=0,1,\ldots$, there are polynomials $Q_{N}(z_{1},\ldots,z_{n})$ of degree $N$ in the variables $z_{k}\in{\mathbb{R}}$, such that
$|\langle(H^{u}+\epsilon I)^{-1/2}\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]|(H^{u}+\epsilon I)^{-1/2}\Psi^{\prime}\rangle|$
$\leq\frac{\int_{{\mathbb{R}}^{3n}}|E^{n}_{\epsilon,u}(P)|^{-1/2}|\Phi(P)||Q_{N}(\nabla_{\vec{p}_{1}})\Phi^{\prime}(P)E^{n}_{\epsilon,u}(P)^{-1/2}|d^{3n}p}{(1+|\vec{x}|)^{N}}$
where $P=(\vec{p}_{1},\ldots,\vec{p}_{n})$ and the Schwartz functions $\Phi,\Phi^{\prime}\in\mathscr{S}({\mathbb{R}}^{3n})$ are obtained from $\Psi$ according to (121). At this point, it is not difficult to see that, due to the special form of the maps $E^{n}_{\epsilon,u}(p_{1},\ldots,p_{n})$ and the fact that the function $\Phi$ is Schwartz, there are constants $C_{N}$ such that
$\int_{{\mathbb{R}}^{3n}}E^{n}_{\epsilon,u}(P)^{-1/2}|\Phi(P)||Q_{N}(\partial_{p_{1}})\Phi^{\prime}(P)E^{n}_{\epsilon,u}(P)^{-1/2}|d^{3n}p\leq C_{N}<+\infty\quad\mbox{uniformly in $\epsilon\in[0,a)$.}$
We are now in a position to apply once more the dominated convergence theorem to the left-hand side of (72) with respect to the $d\Sigma$ integration, proving that (72) holds. This completes the proof of (c) for $n>0$ in the elementary case considered. Let us conclude the proof of (c). The space of density matrices of type $\rho^{\prime}=\sum_{i=1}^{N}p_{i}|\Psi_{i}\rangle\langle\Psi_{i}|$, for $\Psi_{i}\in{{\cal H}}^{(n)}$ with a given common $n$, is dense in ${\mathsf{S}}({{\cal H}}^{(n)})$ in the norm $||\cdot||_{1}$ since ${\mathfrak{S}}_{0}\cap{{\cal H}}^{(n)}$ is dense in ${{\cal H}}^{(n)}$ (we leave to the reader the elementary proof based on the fact that the spectral decomposition of a density matrix is a series of operators $p_{j}|\Psi_{j}\rangle\langle\Psi_{j}|$, $p_{j}\geq 0$, which converges in the norm $||\cdot||_{1}$ and that these operators can be approximated by the previously considered ones in the same topology). As a consequence, if $\rho\in{\mathsf{S}}({{\cal H}}^{(n)})$, and $\rho^{\prime}$ is as above, $|tr(\rho A^{u}_{f}(\Delta))-tr(\rho{\mathsf{A}}^{u}_{f,\epsilon}(\Delta))|$
$|tr(\rho A^{u}_{f}(\Delta))-tr(\rho P_{n}{\mathsf{A}}^{u}_{f,\epsilon}(\Delta)P_{n})|=|tr(\rho A^{u}_{f}(\Delta))-tr(\rho{\mathsf{A}}_{\epsilon,3}(\Delta))|$
$\leq|tr((\rho-\rho^{\prime}){\mathsf{A}}_{\epsilon,3}(\Delta))|+|tr(\rho^{\prime}A^{u}_{f}(\Delta))-tr(\rho^{\prime}{\mathsf{A}}_{\epsilon,3}(\Delta))|+|tr((\rho-\rho^{\prime})A^{u}_{f}(\Delta))|$
[…]
$\langle\Psi|A^{u}_{f}(\Delta)\Psi\rangle=\sum_{n\in{\mathbb{N}}}\int_{\Delta}\thinspace\left\langle\frac{1}{\sqrt{H^{u}}}\Psi_{n}\left|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]u^{\mu}u^{\nu}_{\Sigma}\right.\frac{1}{\sqrt{H^{u}}}\Psi_{n}\right\rangle d\Sigma(x)\>.$
Here, for that given $\Psi$, the right-hand side is evidently $\sigma$-additive (notice that only a finite number of components $\Psi_{n}$ occur and we can use the dominated convergence theorem for (55)). Polarization proves that weak $\sigma$-additivity is valid on ${{\cal H}}^{(0)\perp}\cap{\mathfrak{S}}_{0}$. To conclude, we extend the proof to a generic $\Psi\in{{\cal H}}^{(0)\perp}$. Let $\Delta_{1},\Delta_{2},\ldots$ be a sequence of mutually disjoint sets of $\mathscr{B}(\Sigma)$ and define $S_{N}:=\sum_{j=1}^{N}A^{u}_{f}(\Delta)$. The operators $S_{N}:{{\cal H}}^{(0)\perp}\to{{\cal H}}^{(0)\perp}$ form an increasing sequence of positive operators, so there exists a positive operator $S:=\sum_{j=1}^{+\infty}A^{u}_{f}(\Delta_{j})\in{\mathfrak{B}}({{\cal H}}^{(0)\perp})$, where the sum is understood in the strong sense (see, e.g., [Mor18]). Therefore, in particular, for $\Psi\in{{\cal H}}^{(0)\perp}$, $\langle\Psi|S\Psi\rangle=\sum_{j=1}^{+\infty}\langle\Psi|A^{u}_{f}(\Delta_{j})\Psi\rangle$. In the special case $\Psi\in{{\cal H}}^{(0)\perp}\cap{\mathfrak{S}}_{0}$, $\sigma$-additivity requires $\langle\Psi|S\Psi\rangle=\sum_{j=1}^{+\infty}\langle\Psi|A^{u}_{f}(\Delta_{j})\Psi\rangle=\langle\Psi|A^{u}_{f}(\cup_{j=1}^{+\infty}\Delta_{j})\Psi\rangle$. Since ${{\cal H}}^{(0)\perp}\cap{\mathfrak{S}}_{0}$ is dense in ${{\cal H}}^{(0)^{\perp}}$ and $S,A^{u}_{f}(\cup_{j=1}^{+\infty}\Delta_{j})\in{\mathfrak{B}}({{\cal H}}^{(0)^{\perp}})$, the identity holds for every $\Psi\in{{\cal H}}^{(0)^{\perp}}$. Eventually, the standard argument based on polarization implies $S=A^{u}_{f}(\cup_{j=1}^{+\infty}\Delta_{j})$, so that $\sum_{j=1}^{+\infty}\langle\Psi|A^{u}_{f}(\Delta_{j})\Psi^{\prime}\rangle=\langle\Psi|A^{u}_{f}(\cup_{j=1}^{+\infty}\Delta_{j})\Psi^{\prime}\rangle$ if $\Psi,\Psi^{\prime}\in{{\cal H}}^{(0)\perp}$, concluding the proof of weak $\sigma$-additivity.
[…]
We conclude the proof by establishing (f) and (g). The positive-energy requirment is trivially valid due to the positivity of the Hamiltonian $H^{v}$ for $v\in{\mathsf{T}}_{+}$. Covariance of the constructed families of POVMs easily follows from (c) of Proposition 5.1 and the fact that $U$ leaves the spaces ${{\cal H}}^{(0)}$ invariant. The only non-trivial property is CC. Since ${{\cal H}}^{(0)\perp}\cap{\mathfrak{S}}_{0}$ is dense in ${{\cal H}}^{(0)\perp}$ and $A_{f}^{u}(\Delta)$ is continuous, it is sufficient to prove that the family of non-negative numbers $\langle\Psi|A_{f}^{u}(\Delta)\Psi\rangle$ satisfies CC for $\Psi\in{{\cal H}}^{(0)\perp}\cap{\mathfrak{S}}_{0}$ whose $N$ non-vanishing components are denoted by $\Psi_{n}$. We know that it holds $\langle\Psi|A_{f}^{u}\Psi\rangle=\sum_{n=1}^{N}\int_{\Delta}\langle\frac{1}{\sqrt{H^{u}}}\Psi_{n}|u^{\mu}:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]\frac{1}{\sqrt{H^{u}}}\Psi_{n}\rangle u^{\nu}_{\Sigma}d\Sigma(x)$. The proof relies entirely on the properties of the smooth current which appears in the integrand above ${\mathbb{M}}\ni x\mapsto J_{\Psi}^{\nu}(x):=\sum_{n=1}^{N}\langle\frac{1}{\sqrt{H^{u}}}\Psi_{n}|u^{\mu}:\thinspace{\hat{T}_{\mu}}\>^{\nu}\thinspace:[f^{2}_{x}]\frac{1}{\sqrt{H^{u}}}\Psi_{n}\rangle$. It is causal and future-directed wherever it does not vanish, due to Proposition 4.7, and conserved in view of (c) of Proposition 4.11. The proof of the validity of CC is then the same as for Theorems 35 and 39 in [Mor23]. ∎
5.2 Center of energy, Newton-Wigner position observable, Heisenberg inequality
We suggest a rather direct physical interpretation of the localization observables we have constructed. Results in [Mor23, DRM24] prove that, concerning single particles and localization observables constructed out of the stress-energy tensor (at a formal level in those references), a nice interplay emerges between the first moment of the POVM on $\Sigma$ and the Newton-Wigner selfadjoint position operator for the Minkowskian observer $u_{\Sigma}$ (see below). Irrespective of the choice of the time direction $u$, the three components of the first moment, viewed as symmetric operators, are the three selfadjoint operators representing the components of the Newton-Wigner position observable on $\Sigma$ (restricted to ${\mathsf{S}}({{\cal H}}_{m})$). This fact is of physical relevance, in particular because the Newton-Wigner observable reduces to the standard notion of position in non-relativistic quantum mechanics when the energy content of the quantum state is negligible with respect to the particle mass. This type of result is quite general, since it holds [DRM24] for a wide class of (positive-energy) relativistic position observables studied by Castrigiano [Cas24] for a massive boson.
We recall to the reader that, if $\Sigma$ is a rest space of a Minkowskian reference frame $u_{\Sigma}\in{\mathsf{T}}_{+}$ and $x^{0},x^{1},x^{2},x^{3}$ are Minkowskian coordinates adapted to $u_{\Sigma}$ with $\Sigma$ corresponding to $x^{0}=0$, the three components of the Newton-Wigner position observable on $\Sigma$ for our massive scalar boson [Mor23] are the unique selfadjoint extensions of the three symmetric operators in ${{\cal H}}_{m}=L^{2}(\mathsf{V}_{m,+},d\mu_{m})$,
$(N_{\Sigma}^{a}\psi)(p)=i\sqrt{E_{u_{\Sigma}}(p)}\frac{\partial}{\partial p^{a}}\frac{\psi(p)}{\sqrt{E_{u_{\Sigma}}(p)}}\quad\mbox{for $\psi\in\mathscr{S}(\mathsf{V}_{m,+})$ and $a=1,2,3$,}$ (73)
where, as usual, $E_{u}(p):=-p\cdot u$ for $u\in{\mathsf{T}}_{+}$.
As $N_{\Sigma}^{a}$ is selfadjoint, it has a PVM and, since the PVMs of the three operators commute, one can define a joint PVM $\{Q(\Delta)\}_{\Delta\in\mathscr{B}(\Sigma)}$. It has the structure, for $\psi\in L^{2}(\mathsf{V}_{m,+},d\mu_{m})$,
$(Q_{\Sigma}(\Delta)\psi)(p):=\int_{\Delta}\thinspace d\Sigma(x)\thinspace\int_{\mathsf{V}_{m,+}}\thinspace\thinspace\thinspace\thinspace d\mu_{m}(k)\frac{e^{-i(p-k)\cdot x}}{(2\pi)^{3}}\sqrt{E_{u_{\Sigma}}(p)}\sqrt{E_{u_{\Sigma}}(k)}\psi(k)\>.$ (74)
A technical summary of the properties of the NW localization observable for a scalar particle and the problems it raises with causality is given in [Mor23]. We only stress that, in spite of its appealing properties, the NW localization observable cannot be considered a physically sound notion of spatial localization due to its conflict with elementary causality requirements like CC. However, as said above, it could still coincide with the first moment of several more meaningful unsharp notions of localization which satisfy CC (Proposition 61 in [DRM24]):
$\int_{\Sigma}x^{a}\langle\psi|A^{u}(dx)\psi\rangle=\langle\psi|N^{a}_{\Sigma}\psi\rangle\quad\mbox{for $\psi\in\mathscr{S}(\mathsf{V}_{m,+})$ and $a=1,2,3$.}$
where $A^{u}$ is a $u$-parametrized family of POVMs $A^{u}|_{\mathscr{B}(\Sigma)}$ for every $\Sigma\in{\cal R}$ according to Definition 2.1 of relativistic spatial localization observable.
If $\Sigma$ is the rest space, at some time, of a Minkowskian reference frame $u_{\Sigma}\in{\mathsf{T}}_{+}$, and we use a Minkowskian coordinate system adapted to $u_{\Sigma}$ such that $\Sigma$ is described by $x^{0}=0$, the three operators describing the components $X^{a}_{\Sigma}$ of the center of $u$-energy on $\Sigma$ of our quantum field (the associated particles) are, for $\Psi\in{\mathfrak{S}}_{0}$ with $\Psi=\oplus_{n=0}^{+\infty}\Psi_{n}$,
$X_{u,\Sigma}^{a}\Psi=0$ $\oplus N_{\Sigma}^{a}\Psi_{1}(p)\oplus\sum_{i=1}^{2}\frac{1}{2}\left\{N_{\Sigma,i}^{a},\frac{E_{u}(p_{i})}{E^{2}_{u}(p_{1},p_{2})}\right\}\Psi_{1}(p_{1},p_{2})$
$\oplus\cdots\oplus\sum_{i=1}^{n}\frac{1}{2}\left\{N_{\Sigma,i}^{a},\frac{E_{u}(p_{i})}{E^{n}_{u}(p_{1},\ldots,p_{n})}\right\}\Psi_{1}(p_{1},\ldots,p_{n})\oplus\cdots$ (75)
where, $\{A,B\}:=AB+BA$ is the standard anti symmetrizator and, as before,
$E_{u}^{n}(p_{1},\ldots,p_{n}):=-\sum_{j=1}^{n}u\cdot p_{j}\>.$
In the more implicit form, where $Z_{i}=I\otimes\cdots\otimes I\otimes Z\otimes I\otimes\cdots\otimes I$ and $Z$ occupies the $i$-th slot,
$X_{u,\Sigma}^{a}=\bigoplus_{n=1}^{+\infty}\sum_{i=1}^{n}\frac{1}{2}\left\{N_{\Sigma,i}^{a},\frac{H^{v}_{i}}{H^{v}}\right\}$ (76)
(75) and (76) are nothing but a quantum version of a classical expression of this sort [^11] Notice that the above expression becomes the expression of the coordinates of the standard center of mass in the limit of large mass, when $E^{n}_{u}\to nu^{0}m$.
$X^{a}=\frac{\sum_{n=0}^{+\infty}E^{n}_{u}x_{n}^{a}}{\sum_{n=0}^{+\infty}E^{n}_{u}}$
which, however, takes the non-commutativity of the involved operators into account, as well as the structure of the Fock space.
$X_{\Sigma}^{a}$ is evidently symmetric when defined on ${\mathfrak{S}}_{0}$ and presumably it is also essentially selfadjoint. We only comment that its restriction to the one-particle space ${{\cal H}}^{(1)}={{\cal H}}_{m}$ is essentially selfadjoint because it coincides there with the restriction of the Newton-Wigner operator, which is essentially selfadjoint on $\mathscr{S}(\mathsf{V}_{m,+})$ (see e.g. [Mor23]).
We have the following result.
Proposition 5.3:
Referring to the notions introduced in Theorem 5.2, if $\Sigma$ is any $3$-dimensional rest space, then
$\int_{\Sigma}x^{a}\langle\Psi|A^{u}_{f}(dx)\Psi\rangle=\langle\Psi|X^{a}_{u,\Sigma}\Psi\rangle\quad\mbox{for $\Psi\in{\mathfrak{S}}_{0}$ and $a=1,2,3$}$ (77)
irrespective of $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$ (with $\int_{\mathbb{M}}f^{2}d^{4}x=1$). In particular, for one-particle states
$\int_{\Sigma}x^{a}\langle\psi|A^{u}_{f}(dx)\psi\rangle=\langle\psi|N^{a}_{\Sigma}\psi\rangle\quad\mbox{for $\psi\in\mathscr{S}(\mathsf{V}_{m,+})$ and $a=1,2,3$}$ (78)
irrespective of $u\in{\mathsf{T}}_{+}$ and $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$.
Proof.
Take $\Psi\in{{\cal H}}^{(0)^{\perp}}\cap{\mathfrak{S}}_{0}$ so that $\Psi_{n}=0$ for $n>N_{\Psi}$, which makes meaningful the infinite sum below. According to the Hilbert space isomorphism $J_{n}$ defined in (121), and using $E^{n}_{\epsilon=0,u}=E^{n}_{u}$, we have from (122) and referring to (69),
$\int_{\Sigma}x^{a}\langle\Psi|A^{u}_{f}(dx)\Psi\rangle=$
$i\sum_{n=1}^{+\infty}\frac{n}{2\pi}\int_{\Sigma}\int_{{\mathbb{R}}^{3(n-1)}}\thinspace\int_{{\mathbb{R}}^{6}}\left(\frac{\partial}{\partial k^{a}}e^{i(\vec{k}-\vec{p})\cdot\vec{x}}\right)\widehat{f^{2}}(p-k)\frac{\overline{\Phi(\vec{p},\vec{Q})}\Phi(\vec{k},\vec{Q})}{\sqrt{{E^{n}_{u}(p,Q)}E^{n}_{u}(k,Q)}}u^{\mu}t_{\mu 0}(p,k)\frac{d^{3}pd^{3}kd^{(n-1)}q}{\sqrt{E(p)E(k)}}d^{3}x$
$2(2\pi)^{3}\widehat{f^{2}}(0)\sum_{n=1}^{+\infty}\frac{n}{4\pi}\int_{{\mathbb{R}}^{3n}}\overline{\Phi(\vec{p},\vec{Q})}\left(i\frac{\partial}{\partial p^{a}}\Phi(\vec{p},\vec{Q})\right)\frac{E_{u}(p)}{E_{u}^{n}(p,Q)}d^{3}pd^{(n-1)}q$
$2(2\pi)^{3}\widehat{f^{2}}(0)\sum_{n=1}^{+\infty}\frac{n}{4\pi}\int_{{\mathbb{R}}^{3n}}\overline{\Phi(\vec{p},\vec{Q})}\Phi(\vec{p},\vec{Q})\frac{1}{2}i\frac{\partial}{\partial p^{a}}\frac{E_{u}(p)}{E_{u}^{n}(p,Q)}d^{3}pd^{(n-1)}q$
$\sum_{n=1}^{+\infty}n\left[\int_{{\mathbb{R}}^{3n}}\thinspace\overline{\Phi(\vec{p},\vec{Q})}\left(i\frac{\partial}{\partial p^{a}}\Phi(\vec{p},\vec{Q})\right)\frac{E_{u}(p)}{E_{u}^{n}(p,Q)}+\overline{\Phi(\vec{p},\vec{Q})}\Phi(\vec{p},\vec{Q})\frac{i}{2}\frac{\partial}{\partial p^{a}}\frac{E_{u}(p)}{E_{u}^{n}(p,Q)}\right]d^{3}pd^{(n-1)}q\>.$
The above expression can be rearranged as
$\sum_{n=1}^{+\infty}n\int_{{\mathbb{R}}^{3n}}\thinspace\overline{\Phi(\vec{p},\vec{Q})}\frac{1}{2}\left\{\frac{\partial}{\partial p^{a}},\frac{E_{u}(p)}{E_{u}^{n}(p,Q)}\right\}\Phi(\vec{p},\vec{Q})d^{3}pd^{(n-1)}q\>.$
Since $\Phi(\vec{p},\vec{Q})$ is completely symmetric in its $n$ arguments, the sum above can be rewritten, for every $\Psi\in{\mathfrak{S}}_{0}\cap{{\cal H}}^{(0)\perp}$,
$\int_{\Sigma}x^{a}\langle\Psi|A^{u}_{f}(dx)\Psi\rangle=\sum_{n=1}^{+\infty}\int_{{\mathbb{R}}^{3n}}\thinspace\overline{\Phi(\vec{p}_{1},\ldots,\vec{p}_{n})}\sum_{j=1}^{n}\frac{1}{2}\left\{i\frac{\partial}{\partial p_{j}^{a}},\frac{E_{u}(p_{j})}{E_{u}^{n}(p_{1},\ldots,p_{n})}\right\}\Phi(\vec{p}_{1},\ldots,\vec{p}_{n})d^{3n}p\>.$
This is the thesis, since the action of $N^{a}_{\Sigma}$ on the representation induced by the isomorphism $J_{n}$ on the Schwartz functions $\Phi$ representing the state vectors $\Psi$ is just $i\frac{\partial}{\partial p^{a}}$, whereas the multiplicative operators – like $\frac{E_{u}(p)}{E_{u}^{n}(p,Q)}$ – which are only functions of the momenta $\vec{p}_{i}$, are invariant under the unitary map $J_{n}$. ∎
In spite of the mathematical interest of the proved result, it is difficult to accept the proposed interpretation of the general $A^{u}_{f}$ from a physical perspective for $n\neq 1$. That is because the center of energy of a quantum field does not seem to have the propensity to localize in space! Or, at least, it is really difficult to imagine measurement experiments in which the field is eventually localized at a point in space. By contrast, states of single particles do seem to have this propensity. If we restrict the above result to states $\Psi\in{{\cal H}}^{(1)}={{\cal H}}_{m}$, we once again find the result of [Mor23, DRM24]: the first moment of the POVM on $\Sigma$ is the restriction to ${\mathsf{S}}({{\cal H}}_{m})$ of the Newton-Wigner selfadjoint operator. The novelty with respect to the quoted result is that we have now also found that this result does not depend on the smearing function $f$ used to construct the POVM $A^{u}_{f}$. In the $n=1$ case, the proposed physical interpretation has some chance of being meaningful, since single particles do have the propensity to localize in space under localization experiments. The relevance of this result is that the considered single-particle localization observable is now obtained from standard local and quasi-local observables of QFT in a rigorous way.
$\langle\psi|{A}_{f}^{u}(\Delta)\psi\rangle=\lim_{\epsilon\to 0^{+}}\int_{\Delta}\thinspace\left\langle\frac{1}{\sqrt{H^{u}_{\epsilon}}}\psi\left|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]\frac{1}{\sqrt{H^{u}_{\epsilon}}}\right.\psi\right\rangle u^{\mu}u^{\nu}_{\Sigma}d\Sigma(x)$
defines a POVM on every $\Sigma$ when $\psi\in{{\cal H}}^{(1)}={{\cal H}}_{m}$, and the family of these POVMs is a positive-energy relativistic spatial localization observable for a single particle which satisfies CC.
We finally observe that Heisenberg’s inequality has to be modified according to the new notion of localization (Proposition 61 in [DRM24]). The improved expression reads, for $a=1,2,3$,
$\Delta_{\psi}x^{a}\Delta_{\psi}P_{a}\geq\frac{\hbar}{2}\sqrt{1+4(\Delta_{\psi}P_{a})^{2}\langle\psi|{\mathsf{K}}_{a}\psi\rangle}\>.$
Above, $\Delta_{\psi}x^{a}$ is the standard deviation of the distribution of the coordinate $x^{a}$ in the state represented by the normalized vector $\psi\in\mathscr{S}(\mathsf{V}_{m,+})$, $\Delta_{\psi}P_{a}$ is the analogous quantity for the $a$-th component of the momentum, and ${\mathsf{K}}_{a}$ is a selfadjoint positive operator which is a certain spectral function of $P_{a}$, in principle depending on $u,f,\Sigma$.
See [Mor23, DRM24] for technical discussions on this subject. We only stress that, in the large-mass limit, the above inequality becomes the standard Heisenberg inequality.
5.3 Large-mass/non-relativistic limit: Von Neumann unsharp position measurement
Restricting ourselves to the one-particle case $\psi\in{{\cal H}}_{m}$, we consider, roughly speaking, the large-mass limit. More precisely, we study wave packets for which the relevant values of the momentum $\vec{p}$ satisfy $|\vec{p}|<\thinspace<m$ in natural units. In fact, this limit admits a double interpretation. Restoring the speed of light, the condition becomes $|\vec{p}|<\thinspace<mc$, which may be realized either because $c$ is large, corresponding to the genuinely non-relativistic limit, or not necessarily because the mass is large but relativity holds. For most of the issues discussed here, however, there is no need to distinguish between these two possibilities.
We consider a one-particle state $\psi\in\mathscr{D}(\mathsf{V}_{m,+})$ such that the values of $\vec{p}$ for which $(E(p),\vec{p})\in supp(\psi)$ are negligible with respect to $m$ and the value of $p^{0}=E(p)$ is very close to $m$. Within this approximation we replace $E(p)=\sqrt{\vec{p}^{2}+m^{2}}$ with $m$ and $E_{u}(p)$ with $mu^{0}$, etc., at each occurrence. The probability of finding the particle in a bounded measurable set $\Delta\subset\Sigma$, where $\Sigma$ coincides with $x^{0}=0$, is
$\langle\psi|A_{f}^{u}(\Delta)\psi\rangle=\left\langle\frac{1}{\sqrt{H^{u}}}\psi\left|\int_{\Delta}\thinspace:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]u^{\mu}u^{\nu}_{\Sigma}\frac{1}{\sqrt{H^{u}}}d\Sigma(x)\right.\psi\right\rangle\>.$
Expanding the right-hand side as in (123) and replacing $E(p)$ and $E_{u}(p)$ by $m$ and $u^{0}m$, respectively, and systematically disregarding terms of order $\vec{p}\cdot\vec{k}$, $\vec{k}p_{0}$, $\vec{p}k_{0}$ in $t_{\mu\nu}(p,k)$ in comparison with terms of type $m^{2}$ or $p_{0}k_{0}$ which, in turn, are replaced by $m^{2}$, a lengthy but elementary computation yields
$\langle\psi|A_{f}^{u}(\Delta)\psi\rangle\simeq\int_{\Delta}d\Sigma(x)\int_{\Sigma}\>g_{f}(\vec{y}-\vec{x})|\phi(\vec{y})|^{2}d\Sigma(y)=\int_{\Sigma}(\chi_{\Delta}*g_{f})(\vec{y})|\phi(\vec{y})|^{2}d\Sigma(y)$ (79)
where [^12] The normalization factor, which includes the mass, in the classical limit is automatically embodied in the definition of the wavefunction in momentum representation. It appears here as a relic of the normalization with respect to $\mu_{m}(p)$ on the mass shell and stems from the fact that our discussion obviously does not apply to the massless case.
$\phi(\vec{x}):=\frac{1}{\sqrt{m}}\int_{{\mathbb{R}}^{3}}\frac{e^{i\vec{p}\cdot\vec{x}}}{(2\pi)^{3/2}}\psi(\vec{p})d^{3}p\>,\quad g_{f}(\vec{x}):=\int_{\mathbb{R}}f(x^{0},\vec{x})^{2}dx^{0}\>.$
so that $g_{f}\in\mathscr{D}_{\mathbb{R}}(\Sigma)$, $g_{f}\geq 0$ and $\int_{\Sigma}gd\Sigma=1$. In this non-relativistic approximation the rest space $\Sigma$ can be associated indifferently with $u$ or $u^{\prime}$. Expression (79) is just a Von Neumann model of an unsharp (indirect) position measurement (see e.g. Section 2.3.1 of [Bus09]) where $g$ is related to the wavefunction of the probe particle used to measure the position of a particle of the quantum field of mass $m$ (and this may suggest an indication toward a more concrete interpretation of the smearing procedure in QFT).
Remark 5.4:
We consider, in a given reference frame adapted to $\Sigma$ – indifferently associated with $u$ or $u^{\prime}$ if we are interested in the proper non-relativistic case $c\to+\infty$ as discussed before – smearing functions of the form $f(x)=h(\vec{x})h^{\prime}(x^{0})$, where $h,h^{\prime}$ are real, smooth and compactly supported and which we can always assume to be normalized, $\int_{\mathbb{R}}h^{\prime 2}dx^{0}=1$, and $\int_{\Sigma}h(\vec{x})^{2}d\Sigma=1$. In this case $g_{f}=h^{2}$ by construction. We can consider three regimes.
  • (1)
    As expected from results in [Mor23], the considered localization probability in the large-mass limit tends to become the standard one predicted by non-relativistic quantum mechanics in the limit where $g_{f}(\vec{x})$ tends to $\delta(\vec{x})$.
  • (2)
    If, conversely, the function $g_{f}$ tends to become a constant function (with constant integral $1$), we obtain a more and more imprecise notion of localization in the large-mass limit. That is because the convolution $\chi_{\Delta}*g_{f}$ with an almost constant function makes indistinguishable the characteristic functions of a pair of distinct sets $\Delta$ and $\Delta^{\prime}$.
  • (3)
    Finally, the large-mass/non-relativistic approximation does not depend on the choice of the temporal function $h^{\prime}$, which can be taken to have arbitrarily large support (preserving the constant value of its integral) or to tend to a delta function in time. The former regime is a convenient setup for locally minimizing the negative-energy gap, as we shall discuss in Section 6.1. $\blacksquare$
6 Commutativity of conditional localization POVMs of causally separated laboratories
The families of effects constructed in Theorem 5.2, though arising from (quasi)local observables of QFT, do not satisfy the commutativity requirement for causally separated detection regions $\Delta$ and $\Delta^{\prime}$. This is not a genuine physical problem, as discussed in [Mor26]. The propensity of a particle to localize at a single position does not permit, for instance, invoking the no-signaling principle, which would in turn imply commutativity of the effects. What we intend to investigate here is whether commutativity can be restored by passing to a notion of conditional probability for finite-size laboratories, as discussed in Section 5.5 of [Mor26].
From the purely mathematical viewpoint, non-commutativity of effects $A^{u}_{f}(\Delta),A^{u^{\prime}}_{f}(\Delta^{\prime})$ for $\Delta,\Delta^{\prime}$ causally separated is a consequence of a pair of features.
  • (a)
    First of all, non-commutativity arises from the orthogonal projectors $P_{n}$ – and we are interested in the special case $n=1$ – which appear in (69). They are not local observables in the sense of AHK.
This issue can easily be circumvented by moving the projectors from the operator to the states. In other words, we choose to deal only with one-particle states [^13] The issue is not completely solved from a philosophical viewpoint, since, in order to know whether a field state contains only one particle (or a definite number $n$ of particles), one should collect the whole information about the state from an entire rest space, whereas instruments work locally. However, we shall not address this second-order issue here. and, in order to compute the detection probability of a single particle, we directly use the operators ${\mathsf{A}}_{f,\epsilon}^{u}(\Delta)$ defined in (63), eventually taking the limit as $\epsilon\to 0^{+}$, instead of the effects $A_{f}^{u}(\Delta)$ as in (69). Obviously, when dealing with states in ${\mathsf{S}}({{\cal H}}^{(n)})$ with $n>0$, this is equivalent to the other way around, as stated in (68), since the projectors are already embodied in the states $\rho=P_{n}\rho P_{n}$. However, it is worth stressing that the operators ${\mathsf{A}}_{f,\epsilon}^{u}(\Delta)$, though in ${\mathfrak{B}}({\mathfrak{F}}_{s}({{\cal H}}_{m}))$, are not effects, contrary to $A_{f}^{u}(\Delta)$, since ${\mathsf{A}}_{f,\epsilon}^{u}(\Delta)\not\geq 0$ in general, by (e) of Proposition 5.1.
  • (b)
    The second source of non-commutativity is the appearance of the non-local operators $\frac{1}{\sqrt{H^{u}_{\epsilon}}}$ in the above expression: they do not commute with $\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]$.
In principle, as anticipated, this issue could be addressed by referring to conditional POVMs localized in laboratories, as discussed in the introduction. A laboratory $L(\Delta_{0}):=(\Delta^{\perp\!\!\!\!\perp}_{0})^{\perp\!\!\!\!\perp}\subset{\mathbb{M}}$ is defined by assigning a suitable bounded region (typically a non-empty open set with compact closure) $\Delta_{0}$ in a rest frame $\Sigma$, and one is interested in the (conditional) probability of detecting a particle in subsets $\Delta\subset\Delta_{0}$. We expect that, though the localization effects of a single laboratory do not commute, the localization effects associated with a pair of (sharply) causally separated laboratories based on $\Delta_{0}$ and $\Delta^{\prime}_{0}$ do, if these effects refer to conditional probabilities. That is because these operators are supposed to be constructed in terms of proper local operators, localized in neighborhoods of the laboratories. Causal separation depends not only on the regions $\Delta_{0},\Delta_{0}^{\prime}$ but also on the support of the smearing function $f^{2}$ used in the definition of the stress-energy tensor operator.
According to [Mor26], the idea is therefore to define a “conditioned POVM” in the laboratory based on $\Delta_{0}$, whose effects are labeled by regions $\Delta\subset\Delta_{0}$, and have a form of this type
$V\frac{1}{\sqrt{{\mathsf{A}}^{u}_{f,\epsilon}(\Delta_{0})}}{\mathsf{A}}^{u}_{f,\epsilon}(\Delta)\frac{1}{\sqrt{{\mathsf{A}}^{u}_{f,\epsilon}(\Delta_{0})}}V^{\dagger}\>.$ (80)
(More precisely, these POVMs have the intepretation of conditional POVMs for states whose proability to find the system in $\Delta_{0}$ is close to $1$ according to (5) and the discussion below it.) As discussed in [Mor26], there is the possibility that effects of this sort, referred to different regions $\Delta_{0},\Delta^{\prime}_{0}$, included in causally separated neighborhoods, commute.
The evident problem is that the operators ${\mathsf{A}}^{u}_{f,\epsilon}(\Delta)$ and ${\mathsf{A}}^{u}_{f,\epsilon}(\Delta_{0})$ are not positive! So the above construction seems pointless because it would imply some sort of “negative probabilities” (or even worse, a complex notion of probability, since square roots come into the play). We want to address this issue in the next sections by using energies directly instead of probabilities.
We stress that, if an operator of the form (80), or of a similar kind, is required to be positive and to belong to a local algebra in the AHK sense, then it should be viewed as the restriction to the one-particle space of an operator defined on the full Hilbert space, as in the discussion of (a). It then follows from the Reeh–Schlieder theorem (Proposition 4.8) that positivity on the full Hilbert space is incompatible with the requirement that the vacuum expectation value of the operator vanish. Accordingly, detectors formalized in this manner necessarily exhibit the well-known phenomenon of dark counts. This is a familiar issue in local quantum physics, and it has recently been revisited in a quantitative framework in [FaCo26]. In the scattering-theory literature, by contrast, a different standpoint is usually adopted: detector operators are assumed to annihilate the vacuum, $B\Omega=0$, and are therefore not local. Instead, one works with quasi-local operators, as is done in Haag–Ruelle scattering theory; see in particular [Haa96, Ara09]. In the present paper, we instead consider a genuinely local detector constructed from the local energy density of a quantum field. In this setting, the occurrence of dark counts is unavoidable.
6.1 Bounds on negative energy in finite laboratories
Instead of considering probabilities, let us focus directly on local energies. We can try to define detection probabilities starting from a local notion of Hamiltonian, and compare the energy content in $\Delta$ of a one-particle state with the energy content in $\Delta_{0}\supset\Delta$ of that state. The operators associated with local energy are obviously affected by the problem of negative energy, which in turn would give rise to “negative probabilities”. In some sense these are the same “negative probabilities” measured by the operators ${\mathsf{A}}^{u}_{f,\epsilon}(\Delta_{0})$. From now on, if $\Sigma$ is a rest space, we define the subfamily of bounded Borel sets
$\mathscr{B}_{b}(\Sigma):=\{\Delta\in\mathscr{B}(\Sigma)\>|\>\mbox{$\Delta$ is bounded}\}\>.$ (81)
Proposition 6.1:
Take $u\in{\mathsf{T}}_{+}$, $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$ and suppose that $\Delta\in\mathscr{B}_{b}(\Sigma)$ for a rest space $\Sigma$. There is an operator ${\mathsf{H}}_{f}^{u}(\Delta):{\mathfrak{S}}_{0}\to{\mathfrak{F}}_{s}({{\cal H}}_{m})$, called the local Hamiltonian associated with $\Delta$, that is uniquely defined by
$\langle\Psi|{\mathsf{H}}_{f}^{u}(\Delta)\Psi^{\prime}\rangle=\int_{\Delta}\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]\Psi^{\prime}\rangle u^{\mu}u^{\nu}_{\Sigma}d\Sigma(x)\quad\mbox{for every $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$.}$ (82)
In particular it holds
${\mathsf{H}}_{f}^{u}(\Delta)=:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{\Delta}]\quad\mbox{where}\quad f_{\Delta}(y):=\int_{\Delta}f^{2}(y-x)d\Sigma(x)\>.$ (83)
The following further facts are valid.
  • (a)
    ${\mathsf{H}}_{f}^{u}(\Delta)$ is symmetric, leaves ${\mathfrak{S}}_{0}$ invariant and satisfies the covariance relation
    $U_{g}{\mathsf{H}}_{f}^{u}(\Delta)U_{g}^{-1}={\mathsf{H}}_{g_{*}f}^{gu}(g\Delta)\quad\mbox{for every $g\in IO(1,3)_{+}$.}$
  • (b)
    In general, ${\mathsf{H}}_{f}^{u}(\Delta)$ is not positive, but it is bounded from below.
  • (c)
    If $\Delta\subset\Sigma$ is an open set with compact closure and the globally hyperbolic region $L(\Delta):=(\Delta^{{\perp\!\!\!\!\perp}})^{{\perp\!\!\!\!\perp}}$ is the associated laboratory, then
    $\langle\Psi|{\mathsf{H}}_{f}^{u}(\Delta)\Psi^{\prime}\rangle=\int_{S}\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]\Psi^{\prime}\rangle u^{\mu}u^{\nu}_{S}(x)dS(x)\quad\mbox{for every $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$.}$ (84)
    Above $S\subset L(\Delta)$ is any spacelike smooth Cauchy surface of $L(\Delta)$ and $u_{S}(x)$ its future-oriented unit normal vector at $x\in S$, $dS(x)$ being the metric-induced measure on $S$.
  • (d)
    Taking (62) into account, if $\epsilon>0$, and $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$
    $\langle\Psi|{\mathsf{A}}^{u}_{f,\epsilon}(\Delta)\Psi^{\prime}\rangle=\left\langle\Psi\left|{\frac{1}{\sqrt{H^{u}_{\epsilon}}}{\mathsf{H}}_{f}^{u}(\Delta)\frac{1}{\sqrt{H^{u}_{\epsilon}}}}\right.\Psi^{\prime}\right\rangle\>.$ (85)
Everything asserted remains valid if one replaces $f^{2}$ by $f=\sum_{i=1}^{N}f_{i}^{2}$ with $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$.
Proof.
See Appendix B. ∎
Failure of positivity of ${\mathsf{H}}_{f}^{u}(\Delta)$ arises from the appearance of states with averaged negative energy. Nevertheless, taking advantage of some crucial results in [Few12], we are about to prove that, referring to a Minkowskian coordinate system adapted to $u$, we can make the “negative energy gap” as small as desired, in every reference frame $u^{\prime}$, generally different from $u$, by choosing the support of the temporal part $h^{\prime}=h^{\prime}(x^{0})$ of the smearing function $f(x)=h(\vec{x})h^{\prime}(x^{0})$ suitably large (the coordinates being adapted to $u$). It is worth stressing that this can be done while leaving untouched the spatial part $h(\vec{x})$ of the smearing function. This function is responsible for the precision of the spatial position measurement according to Remark 5.4 in the non-relativistic limit. In practice, we integrate in space a corrected version of the energy density by adding a term to the normally ordered stress-energy tensor operator depending on a small parameter $\eta>0$. Since we deal with spatial regions of finite extent, the amount of added energy remains finite and can be made arbitrarily small by suitably enlarging the support of the temporal smearing function $h^{\prime}$.
Henceforth ${\mathsf{H}}_{f,\eta}^{u}(\Delta)$ is the unique operator such that, for every $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$,
$\langle\Psi|{\mathsf{H}}_{f,\eta}^{u}(\Delta)\Psi^{\prime}\rangle=\int_{\Delta}\int_{{\mathbb{M}}}\left\langle\Psi\left|\left(:\thinspace\hat{T}_{\mu\nu}\thinspace:(y)-\eta g_{\mu\nu}(y)I\right)\right.\Psi^{\prime}\right\rangle f(y-x)^{2}u^{\mu}u^{\nu}_{\Sigma}d\Sigma(x)\>.\quad$ (86)
for every $u\in\mathsf{V}_{+}$, every bounded measurable $\Delta\subset\Sigma$, of every rest space $\Sigma$ adapted to the reference frame $u_{\Sigma}$, $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$, and $\eta\geq 0$. It immediately follows – recalling that $u\cdot u^{\prime}_{\Sigma}<0$! –
${{\mathsf{H}}}^{u}_{f,\eta}(\Delta):={\mathsf{H}}_{f}^{u}(\Delta)+\eta|u\cdot u^{\prime}_{\Sigma}||\Delta|I\>.$ (87)
Notice that the improved stress-energy tensor operator induced by the formal quadratic-form density $\thinspace\hat{T}_{\mu\nu}\thinspace:(y)_{\eta}:=:\thinspace\hat{T}_{\mu\nu}\thinspace:(y)-\eta g_{\mu\nu}I$ is automatically conserved in the usual distributional sense $\thinspace\hat{T}_{\mu\nu}\thinspace:[\partial^{\mu}f]_{\eta}=0$, since the covariant derivative is metric-compatible (the metric is even constant in Minkowskian coordinates!). Finally, (a), (b), and (c) remain valid for ${\mathsf{H}}_{f,\eta}^{u}(\Delta)$, with the latter obtained by replacing the stress-energy tensor by the improved one, and the covariance relation in (a) reads
$U_{g}{\mathsf{H}}_{f,\eta}^{u}(\Delta)U_{g}^{-1}={\mathsf{H}}_{g_{*}f,\eta}^{gu}(g\Delta)\quad\mbox{for every $g\in IO(1,3)_{+}$.}$
Furthermore, as anticipated, (b) can be substantially improved as follows.
Theorem 6.2:
Take $u\in{\mathsf{T}}_{+}$, a Minkowskian coordinate system $x^{0},x^{1},x^{2},x^{3}$ adapted to $u$, and a non-vanishing function $h\in\mathscr{D}_{\mathbb{R}}({\mathbb{R}}^{3})$ with $\int h(\vec{x})^{2}d^{3}x=1$. Finally choose an arbitrarily small $\eta>0$.
There exists $h^{\prime}\in\mathscr{D}_{\mathbb{R}}({\mathbb{R}})$ with $\int h(x^{0})^{\prime 2}dx^{0}=1$ such that, defining $f:=h^{\prime}h$, we have
$\left\langle\Psi\left|\left(:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]-\eta g_{\mu\nu}I\right)\right.\Psi\right\rangle u^{\mu}u^{\prime\nu}\geq c^{u}_{\eta,f}|u\cdot u^{\prime}|||\Psi||^{2}\quad\mbox{if $u^{\prime}\in{\mathsf{T}}_{+}$, $x\in{\mathbb{M}}$, $\Psi\in{\mathfrak{S}}_{0}$}$ (88)
for some finite constant $c^{u}_{\eta,f}>0$ independent of $u^{\prime}$. As a consequence
  • (a)
    if $\Psi\in{\mathfrak{S}}_{0}$, the smooth conserved current $J_{\nu}(x):=-\left\langle\Psi\left|\left(:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]-\eta g_{\mu\nu}I\right)\right.\Psi\right\rangle u^{\mu}$ is causal and future-directed wherever it does not vanish;
  • (b)
    the operators ${{\mathsf{H}}}^{u}_{f,\eta}(\Delta)$ are positive and monotonous:
[…]
  • (c)
    the operators ${{\mathsf{H}}}^{u}_{f,\eta}(\Delta)$ are strictly positive:
    ${{\mathsf{H}}}^{u}_{f,\eta}(\Delta)\geq c^{u}_{\eta,f}|u\cdot u_{\Sigma}||\Delta|I\geq 0\quad\mbox{if $\Delta\in\mathscr{B}_{b}(\Sigma)$, for every rest space $\Sigma$.}$ (90)
Proof.
The proof of (88) relies on the following lemma, which in turn follows from a direct application of general results presented in [Few12].
Lemma 6.3:
With the main hypotheses, if $h^{\prime}\in\mathscr{D}_{\mathbb{R}}({\mathbb{R}})$ and $\Psi\in{\mathfrak{S}}_{0}$ with $\Psi||=1$,
$\int_{\mathbb{R}}\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:(x^{0},\vec{x})\Psi\rangle h^{\prime}(x^{0})^{2}u^{\mu}u^{\prime\nu}dx^{0}\geq-\sqrt{1+\vec{u}^{\prime 2}}\frac{1}{16\pi^{3}}\int_{m}^{+\infty}|\hat{h^{\prime}}(s)|^{2}s^{4}Q_{3}(s)ds$ (91)
where $u^{\nu}=\delta_{0}^{\nu}$, $u^{\prime 0}=\sqrt{1+\vec{u}^{\prime 2}}$, and $Q_{3}:[1,+\infty)\to[0,1)$ is the smooth, strictly increasing, bounded function
$Q_{3}(x):=\left(1-\frac{1}{x^{2}}\right)^{1/2}\left(1-\frac{1}{2x^{2}}\right)-\frac{1}{2x^{4}}\ln\left(x+\sqrt{x^{2}-1}\right)$
which satisfies $Q_{3}(1)=0$ and $Q_{3}(x)\to 1$ as $x\to+\infty$.
Proof.
See Appendix B. ∎
In the following $f:=h\cdot h^{\prime}$. We therefore have
$\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}]\Psi\rangle u^{\mu}u^{\prime\nu}=\int_{{\mathbb{R}}^{4}}\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:(x^{0},\vec{x})\Psi\rangle f(x^{0},\vec{x})^{2}u^{\mu}u^{\prime\nu}dx^{0}d^{3}x$
$\int_{{\mathbb{R}}^{4}}\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:(x^{0},\vec{x})\Psi\rangle h^{\prime}(x^{0})^{2}h(\vec{x})^{2}u^{\mu}u^{\prime\nu}dx^{0}d^{3}x\geq-|u\cdot u^{\prime}|\frac{1}{16\pi^{3}}\int_{m}^{+\infty}|\hat{h^{\prime}}(s)|^{2}s^{4}Q_{3}(s)ds\>.$
As a consequence, since the inequality above is valid for every normalized $\Psi\in{\mathfrak{S}}_{0}$ and this space is invariant under the unitary action of the Poincaré group, we have in particular that
$\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}^{2}]\Psi\rangle u^{\mu}u^{\prime\nu}=\langle V^{-1}_{x}\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}]V^{-1}_{x}\Psi\rangle u^{\mu}u^{\prime\nu}\geq-|u\cdot u^{\prime}|\frac{1}{16\pi^{3}}\int_{m}^{+\infty}|\hat{h^{\prime}}(s)|^{2}s^{4}Q_{3}(s)ds$
where $V_{x}=U_{(I,x)}$ is the unitary representation of spacetime translations (referring to some choice of the origin). This inequality permits us to prove (88). It is sufficient to show that, for every given $\eta>0$, there exists a corresponding $h^{\prime}$ such that
$\eta>\frac{1}{16\pi^{3}}\int_{m}^{+\infty}|\hat{h^{\prime}}(s)|^{2}s^{4}Q_{3}(s)ds\>.$ (92)
If this is true, the required constant $c^{u}_{\eta,f}$ can be defined as $c^{u}_{\eta,f}:=(\eta-\eta^{\prime})$, where $\eta^{\prime}<\eta$ is positive and sufficiently close to $\eta$ so that (92) is still valid with $\eta^{\prime}$ in place of $\eta$. To prove (92), consider a family of smooth compactly supported real functions $h^{\prime}_{\delta}(x^{0})=\sqrt{\delta}\chi(\delta x^{0})$ where $\delta>0$ and $\chi\in\mathscr{D}_{\mathbb{R}}({\mathbb{R}})$ is such that $\int\chi^{2}dx^{0}=1$. $\widehat{h^{\prime}_{\delta}}$ is Schwartz and $\widehat{h^{\prime}_{\delta}}(s)=\frac{1}{\sqrt{\delta}}\widehat{\chi}(s/\delta)$. The dominated convergence theorem proves that $\int_{m}^{+\infty}|\widehat{h^{\prime}_{\delta}}(s)|^{2}s^{4}Q_{3}(s)ds=$
$\delta^{-1}\int_{m}^{+\infty}|\widehat{\chi^{2}}(s/\delta)|^{2}s^{4}Q_{3}(s)ds=\int_{{\mathbb{R}}}\chi_{[m/\delta,+\infty)}(y)\delta^{4}|\widehat{\chi^{2}}(y)|^{2}y^{4}Q_{3}(\delta y)dy\to 0\quad\mbox{as $\delta\to 0^{+}$.}$
Hence (92) is valid for a given $\eta>0$ provided one uses $h^{\prime}=h^{\prime}_{\delta}$ with a sufficiently small $\delta>0$. At this point (a), (b), and (c) follow immediately. ∎
6.2 Conditional localization POVMs in laboratories
A suitable candidate for a conditional localization POVM in a laboratory $L(\Delta_{0})$, for $\Delta_{0}\subset\mathscr{B}_{b}(\Sigma)$ with $|\Delta_{0}|>0$, is expected to be
$B_{\Delta_{0}}(\Delta):=\frac{1}{\sqrt{{\mathsf{H}}^{u}_{f,\eta}(\Delta_{0})}}{\mathsf{H}}^{u}_{f,\eta}(\Delta)\frac{1}{\sqrt{{\mathsf{H}}^{u}_{f,\eta}(\Delta_{0})}}\quad\mbox{for ${\cal R}\ni\Delta\subset\Delta_{0}$}\quad\quad\quad\mbox{(attempt)}$ (93)
where we have chosen $f$ so that, for the given $\eta>0$, (90) is true with $\Delta=\Delta_{0}$. The mathematical problem with this naive definition is that we cannot take advantage of spectral calculus in defining the inverse square root of the indicated operators, since they are not selfadjoint and we do not even know whether they are essentially selfadjoint. Nevertheless, since they are bounded below by (90) (indeed, with a strictly positive lower bound), we can fruitfully use several results from the theory of Krein-von Neuamann extensions, in particular focusing on the Friedrichs extension. We assume that the reader is familiar with the elementary theory of quadratic forms and self-adjoint extensions of symmetric operators [ReSa75]. We also suggest the recent review [GeSu25] on the properties of Friedrichs extensions and Krein-von Neumann extensions. In fact we have the following theorem, whose proof is somewhat technical because it requires several lemmas proved in the appendix.
Notation 6.4:
In the rest of the paper $\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta)$ denotes the Friedrichs extension of ${{\mathsf{H}}}^{u}_{f,\eta}(\Delta):{\mathfrak{S}}_{0}\to{\mathfrak{S}}_{0}\subset{\mathfrak{F}}_{s}({{\cal H}}_{m})$. $\blacksquare$
Notice that this is the only selfadjoint extension if the operator is essentially selfadjoint, in which case it coincides with the closure: $\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta)=\overline{{{\mathsf{H}}}^{u}_{f,\eta}(\Delta)}$.
Theorem 6.5:
Take $u\in{\mathsf{T}}_{+}$, an arbitrarily small $\eta>0$, and $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$ such that (90) is true. The following facts are valid for every rest space $\Sigma$.
  • (a)
    If $\Delta_{0}\subset\mathscr{B}_{b}(\Sigma)$ with $|\Delta_{0}|>0$, then there exists a unique family of operators $B_{\Delta_{0}}(\Delta)^{u}_{f,\eta}\in{\mathfrak{B}}({\mathfrak{F}}_{s}({{\cal H}}_{m}))$ for $\Delta\in\mathscr{B}(\Delta_{0})$ such that
    $B_{\Delta_{0}}(\Delta)^{u}_{f,\eta}\Psi=\frac{1}{\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}{{\mathsf{H}}}^{u}_{f,\eta}(\Delta)\frac{1}{\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}\Psi\quad\mbox{for}\quad\Psi\in{\mathfrak{S}}_{00}$ (94)
    where ${\mathfrak{S}}_{00}:={\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}({\mathfrak{S}}_{0})$ is dense and $\frac{1}{\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}\in{\mathfrak{B}}({\mathfrak{F}}_{s}({{\cal H}}_{m}))$.
  • (b)
    The map $\mathscr{B}(\Delta_{0})\ni\Delta\mapsto B_{\Delta_{0}}(\Delta)^{u}_{f,\eta}$ is a normalized POVM absolutely continuous with respect to the Lebesgue measure on $\Sigma$.
Remark 6.6:
Before proving the theorem, we observe that if $\Delta_{0}$ is sufficiently regular that $L(\Delta_{0})$ is a globally hyperbolic spacetime in its own right with $\Delta_{0}$ as a smooth spacelike Cauchy surface, then the POVM $B_{\Delta_{0}}(\Delta)^{u}_{f,\eta}$ can be extended to a weaker version of relativistic spatial localization observable on the whole spacetime $L(\Delta_{0})$. This can be done with the same technology developed in [DRM24]. That mathematical technology, starting from (a) of Theorem 6.2, should also prove that the constructed relativistic spatial localization observable satisfies a natural generalization of the causal condition CC denoted by GCC therein. If $\Delta_{0}$ is only Borel, a similar generalization should be possible in the corresponding element of the causal logic according to [CDM26], proving therein the validity of an even more general causal condition for achronal sets. These issues will be investigated elsewhere. $\blacksquare$
Proof.
(a) We need some preliminary lemmas. If $A\geq 0$ is a symmetric operator, $A_{F}\geq 0$ henceforth denotes its selfadjoint Friedrichs extension.
Lemma 6.7:
If $A:D(A)\to{{\cal H}}$ is a symmetric operator with $A\geq 0$ and $A_{F}$ is its Friedrichs extension, then
  • (a)
    $\overline{\sqrt{A_{F}}(D(A))}=Ker(\sqrt{A_{F}})^{\perp}=Ker(A_{F})^{\perp}$;
  • (b)
    if $A\geq cI$ with $c>0$, then $A_{F}\geq cI$. In this case: $\overline{\sqrt{A_{F}}(D(A))}=Ran(\sqrt{A_{F}})={{\cal H}}$ and $A_{F}^{-\alpha}\in{\mathfrak{B}}({{\cal H}})$ for $\alpha>0$.
Proof.
See Appendix B. ∎
Lemma 6.8:
Consider a pair of symmetric operators $A^{0}\geq A\geq 0$ with $D(A)=D(A^{0})$. The following facts are true.
  • (a)
    It holds
    $\left\langle\sqrt{A^{0}_{F}}x\left|\sqrt{A^{0}_{F}}\right.x\right\rangle\geq\langle\sqrt{A_{F}}x|\sqrt{A_{F}}x\rangle\quad\mbox{ for $x\in D(A_{F}^{0})$}$
    so that, in particular,
    $\langle x|A^{0}_{F}x\rangle\geq\langle x|A_{F}x\rangle\geq 0\quad\mbox{for $x\in D(A^{0}_{F})$};$
  • (b)
    if $Ker(\sqrt{A^{0}_{F}})(=Ker(A^{0}_{F}))=\{0\}$, then
    $z||^{2}\geq\left\langle\frac{1}{\sqrt{A^{0}_{F}}}z\left|A\frac{1}{\sqrt{A^{0}_{F}}}z\right.\right\rangle=\left\langle\frac{1}{\sqrt{A^{0}_{F}}}z\left|A_{F}\frac{1}{\sqrt{A^{0}_{F}}}z\right.\right\rangle\geq 0\quad\mbox{for $z\in\sqrt{A^{0}_{F}}(D(A^{0}))$;}$
  • (c)
[…]
    $\frac{1}{\sqrt{A^{0}_{F}}}A\frac{1}{\sqrt{A^{0}_{F}}}:\sqrt{A^{0}_{F}}(D(A^{0}))\to{{\cal H}}\>.$
Proof.
See Appendix B. ∎
Returning to the proof of (a), the claim follows immediately from these two lemmas by using $A^{0}={\mathsf{H}}^{u}_{f,\eta}(\Delta_{0})$ and $A={\mathsf{H}}^{u}_{f,\eta}(\Delta)$, both defined on the common dense domain ${\mathfrak{S}}_{0}$, since ${\mathsf{H}}^{u}_{f,\eta}(\Delta_{0})\geq{\mathsf{H}}^{u}_{f,\eta}(\Delta)\geq 0$ and ${\mathsf{H}}^{u}_{f,\eta}(\Delta_{0})\geq cI$ with $c>0$ in view of Theorem 6.2.
Concerning (b), we observe that the operators $B_{\Delta_{0}}(\Delta)^{u}_{f,\eta}$ are effects, as follows from (b) of Lemma 6.8. Furthermore, obviously, $B_{\Delta_{0}}(\Delta_{0})^{u}_{f,\eta}=I$ by construction, since the identity is the unique continuous extension of the said quadratic form. The properties of $\sigma$-additivity and absolute continuity are proved by taking advantage of the Vitali-Hahn-Saks theorem, as in the proof of (a) of Proposition 5.5 in [Mor26]. ∎
Corollary 6.9:
Under the hypotheses of Theorem 6.5, $\frac{1}{\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}{{\mathsf{H}}}^{u}_{f,\eta}(\Delta)\frac{1}{\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}$ is essentially selfadjoint on its natural domanin ${\mathfrak{S}}_{00}={\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}({\mathfrak{S}}_{0})$which therefore is a core:
$B_{\Delta_{0}}(\Delta)^{u}_{f,\eta}=\overline{\frac{1}{\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}{{\mathsf{H}}}^{u}_{f,\eta}(\Delta)\frac{1}{\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}}\>.$ (95)
The same identity is valid if ${{\mathsf{H}}}^{u}_{f,\eta}(\Delta)$ is replaced above by any other symmetric or even selfadjoint extension of ${{\mathsf{H}}}^{u}_{f,\eta}(\Delta)$.
Proof.
Observe that the dense subspace ${\mathfrak{S}}_{00}:={\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}({\mathfrak{S}}_{0})$ is just the domain of $\frac{1}{\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}{{\mathsf{H}}}^{u}_{f,\eta}(\Delta)\frac{1}{\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}$, since $\frac{1}{\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}$ is bijective. The unique everywhere-defined continuous extension of the symmetric operator $\frac{1}{\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}{{\mathsf{H}}}^{u}_{f,\eta}(\Delta)\frac{1}{\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}$ is its closure, which is selfadjoint. Hence the said symmetric operator is essentially selfadjoint. In particular, its domain is a core. If we replace ${{\mathsf{H}}}^{u}_{f,\eta}(\Delta)$ by some symmetric extension $E$ of it, we obtain a symmetric extension $\frac{1}{\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}E\frac{1}{\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}\supset\frac{1}{\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}{{\mathsf{H}}}^{u}_{f,\eta}(\Delta)\frac{1}{\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}$. Since the right-hand side is essentially selfadjoint, its closure is selfadjoint and thus maximally symmetric. As a consequence, the closure of the left-hand side, which is symmetric as well, must coincide with the closure of the right-hand side. This completes the proof. ∎
To conclude this investigation we examine the interplay between the POVM $B_{\Delta_{0}}(\Delta)^{u}_{f,\eta}$ and the global POVM $A_{f}^{u}(\Delta)$ on the given $\Sigma\supset\Delta_{0}$. We prove that, in fact, the map $\mathscr{B}(\Delta_{0})\ni\Delta\mapsto B_{\Delta_{0}}(\Delta)^{u}_{f,\eta}$ can be interpreted as a conditional POVM constructed out of a non-normalized POVM which is an approximation of $A^{u}_{f}$ with the desired precision, in the sense clarified below, and also using suitable unitary operators in accordance with (4).
First of all, we construct positive bounded operators out of ${\mathsf{A}}^{u}_{f,\epsilon}(\Delta)$ (which arbitrarily approximate the effects $A_{f}^{u}(\Delta)$ as stated in (68) but are not positive), exploiting the improved local Hamiltonians ${\mathsf{H}}^{u}_{f,\eta}$. In other words we consider the unique everywhere-defined bounded extension of the operator
$\frac{1}{\sqrt{H^{u}_{\epsilon}}}{\mathsf{H}}^{u}_{f,\eta}(\Delta)\frac{1}{\sqrt{H^{u}_{\epsilon}}}:{\mathfrak{S}}_{0}\to{\mathfrak{S}}_{0}\>.$
which turns out to have the explicit form
${\mathsf{A}}^{u}_{f,\epsilon,\eta}(\Delta):={\mathsf{A}}^{u}_{f,\epsilon}(\Delta)+\eta|u\cdot u_{\Sigma}||\Delta|\frac{1}{H^{u}_{\epsilon}}\in{\mathfrak{B}}({\mathfrak{F}}_{s}({{\cal H}}_{m}))$ (96)
for $\epsilon>0$ and the remaining parameters fixed as before. By construction, for every given arbitrarily small $\eta>0$ we can tune the (temporal part of the) function $f$ in such a way that ${\mathsf{A}}^{u}_{f,\epsilon,\eta}(\Delta)\geq 0$ for all $\Delta\in\mathscr{B}_{b}(\Sigma)$ and every $\epsilon>0$. As asserted, these positive operators, when traced against finite-particle states, approximate the effects $A^{u}_{f}(\Delta)$ with the desired precision in a given laboratory based on $\Delta_{0}\in\mathscr{B}_{b}(\Sigma)$.
Proposition 6.10:
Take $u,\in{\mathsf{T}}_{+}$, $\Delta_{0}\in\mathscr{B}_{b}(\Sigma)$, an arbitrarily small $\eta>0$, and $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$ such that (90) is true. The $\rho,\Delta$-uniform bound holds
$|tr(\rho A^{u}_{f}(\Delta))-\lim_{\epsilon\to 0^{+}}tr(\rho{\mathsf{A}}^{u}_{f,\epsilon,\eta}(\Delta))|\leq\eta\frac{|u\cdot u_{\Sigma}||\Delta_{0}|}{m}$ (97)
for every $\rho\in{\mathsf{S}}({{\cal H}}^{(n)})$ with $n>0$ and every $\Delta\in\mathscr{B}(\Delta_{0})$.
Proof.
It holds $|tr(\rho{{\mathsf{A}}}^{u}_{f,\epsilon}(\Delta))-tr(\rho{{\mathsf{A}}}^{u}_{f,\epsilon,\eta}(\Delta))|\leq\eta|\Delta||u\cdot u_{\Sigma}|tr(\rho\frac{1}{H^{u}_{\epsilon}})\leq\eta|u\cdot u_{\Sigma}|\frac{|\Delta|_{\Sigma}}{m}$ directly from (96). Taking the limit as $\epsilon\to 0^{+}$, the bound remains valid and (68) produces the thesis since $|\Delta|\leq|\Delta_{0}|$. ∎
At this point, we can prove the following result, where the structure of a conditional POVM as in (2) emerges.
Theorem 6.11:
Take $u\in{\mathsf{T}}_{+}$, an arbitrarily small $\eta>0$ and $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$ such that (90) is true. Define the POVM $\mathscr{B}(\Delta_{0})\ni\Delta\mapsto B_{\Delta_{0}}(\Delta)^{u}_{f,\eta}$ for a given $\Delta_{0}\in\mathscr{B}_{b}(\Sigma),\Sigma$ rest space, as in (a) of Theorem 6.5. For every $\epsilon>0$, there is a unitary $V^{u}_{f,\epsilon,\eta,\Delta_{0}}$ such that
$B_{\Delta_{0}}(\Delta)^{u}_{f,\eta}=V^{u}_{f,\epsilon,\eta,\Delta_{0}}\frac{1}{\sqrt{{\mathsf{A}}^{u}_{f,\epsilon,\eta}(\Delta_{0})}}{\mathsf{A}}^{u}_{f,\epsilon,\eta}(\Delta)\frac{1}{\sqrt{{\mathsf{A}}^{u}_{f,\epsilon,\eta}(\Delta_{0})}}V^{u\dagger}_{f,\epsilon,\eta,\Delta_{0}}\>.$ (98)
Proof.
First of all, we observe that (96) and ${{\mathsf{H}}}^{u}_{f,\eta}(\Delta)\geq c^{u}_{\eta,f}|u\cdot u_{\Sigma}||\Delta|I\geq 0$ imply ${\mathsf{A}}^{u}_{f,\epsilon,\eta}(\Delta)\geq\frac{c^{u}_{\eta,f}|u\cdot u_{\Sigma}|}{\epsilon}|\Delta|I\geq 0$ if $\epsilon>0$ and, in particular, ${\mathsf{A}}^{u}_{f,\epsilon,\eta}(\Delta)^{-1}\in{\mathfrak{B}}({\mathfrak{F}}_{s}({{\cal H}}_{m}))$. Furthermore, since $\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})$ is selfadjoint and $1/\sqrt{H^{u}_{\epsilon}}\in{\mathfrak{B}}({\mathfrak{F}}_{s}({{\cal H}}_{m}))$ we have
$\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}\frac{1}{\sqrt{H^{u}_{\epsilon}}}=\left(\frac{1}{\sqrt{H^{u}_{\epsilon}}}\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}\right)^{\dagger}\>\>\mbox{so that}\>\>\left(\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}\frac{1}{\sqrt{H^{u}_{\epsilon}}}\right)^{\dagger}=\overline{\frac{1}{\sqrt{H^{u}_{\epsilon}}}\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}\>,$
where we also used the fact that both sides in the first identity are (closed) densely defined operators. As a consequence, if $\Psi\in\sqrt{H^{u}_{\epsilon}}({\mathfrak{S}}_{0})$ so that $(\sqrt{H^{u}_{\epsilon}})^{-1}\Psi\in{\mathfrak{S}}_{0}=D({{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0}))\subset D(\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0}))\subset D(\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})})$, we find
[…]
$\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta)\frac{1}{\sqrt{H^{u}_{\epsilon}}}\sqrt{H^{u}_{\epsilon}}\frac{1}{\sqrt{{\mathsf{A}}^{u}_{f,\epsilon,\eta}(\Delta_{0})}}V^{\dagger}={\mathsf{A}}^{u}_{f,\epsilon,\eta}(\Delta)\frac{1}{\sqrt{{\mathsf{A}}^{u}_{f,\epsilon,\eta}(\Delta_{0})}}V^{\dagger}\>.$
Using this identity on the right-hand side of the composition (95) and adjusting the left-hand side similarly, we find that the two operators appearing on the two sides of (98) have the same matrix elements on the dense subspace ${\mathfrak{S}}_{00}$. Since both operators are bounded and everywhere defined, the thesis follows. ∎
6.3 Local Weyl and von Neumann algebras and Haag duality
We shall consider the Weyl unitaries $W(f):=e^{i\overline{\hat{\phi}(f)}}$ for $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$ introduced in Proposition 3.5, and we refer to the elementary theory of von Neumann algebras [Tak02, StZs19].
[…]
Definition 6.12:
Let ${\cal O}\subset{\mathbb{M}}$ be an open set.
  • (a)
    The local Weyl algebra ${{\cal W}}({\cal O})\subset{\mathfrak{B}}({\mathfrak{F}}_{s}({{\cal H}}_{m}))$ associated with $\cal O$} is the unital $C^{*}$-algebra generated by the Weyl operators $W(f)$ with $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$ and $supp(f)\subset{\cal O}$. (In other words, it is the operator-norm closure in ${\mathfrak{B}}({\mathfrak{F}}_{s}({{\cal H}}_{m}))$ of linear combinations of products of the aforementioned unitiary operators $W(f)$.)
  • (b)
    ${\mathfrak{W}}({\cal O}):={{\cal W}}({\cal O})^{\prime\prime}$ is the local von Neumann algebra associated with $\cal O$}. $\blacksquare$
By definition, the isotony property holds for both families of algebras:
${{\cal W}}({\cal O})\subset{{\cal W}}({\cal O}_{1})\quad\mbox{and}\quad{\mathfrak{W}}({\cal O})\subset{\mathfrak{W}}({\cal O}_{1})\quad\mbox{if ${\cal O}\subset{\cal O}_{1}$}\>.$ (99)
The Weyl relations (24) imply that the linear subspace
$span\{W(f)\>|\>f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}}),\>supp(f)\subset{\cal O}\}\subset{{\cal W}}({\cal O})$
actually is a unital *-algebra that is dense in ${{\cal W}}({\cal O})$ in the operator norm. Hence it is also dense in the strong operator topology. At this point, von Neumann’s double commutant theorem [Tak02, StZs19] implies that
${\mathfrak{W}}({\cal O})=\overline{span\{W(f)\>|\>f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}}),\>supp(f)\subset{\cal O}\}}^{s}$
$\overline{span\{W(f)\>|\>f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}}),\>supp(f)\subset{\cal O}\}}^{w}$
$=(span\{W(f)\>|\>f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}}),\>supp(f)\subset{\cal O}\})^{\prime\prime}\>.$ (100)
where the closures $\overline{\cdot}^{s}$ and $\overline{\cdot}^{w}$ refer to the strong and weak operator topologies, respectively.
To conclude this short summary, we present a version of Haag duality in the more modern formulation discussed in [Cam07]. This is one of the celebrated mathematical relations in AQFT [Haa96, Ara09]. For the free scalar field, it was first established by Araki in [Ara64].
Definition 6.13:
If $p,q\in{\mathbb{M}}$ are such that $p-q\neq 0$ is future directed, the open double cone generated by them is the set ${\cal O}:=Int(J^{-}(p)\cap J^{+}(q))$.
Referring to the notion of causal complement (3), we define $A^{c}:=Int(A^{\perp\!\!\!\!\perp})$ if $A\subset{\mathbb{M}}$. $\blacksquare$
Observe that every bounded set in $\mathbb{M}$} is contained in a sufficiently large open double cone. In particular, a laboratory $L(\Delta_{0})$, based on a bounded non-empty open set $\Delta_{0}\subset\Sigma$, is always contained in a sufficiently large open double cone $\cal O$} generated by $p,q\in{\mathbb{M}}$ with pqp-q normal to $\Sigma$ and ${\cal O}\cap\Sigma\supset\Delta_{0}$. Open double cones themselves are laboratories based on $\Delta_{0}:={\cal O}\cap\Sigma$ as above. Open double cones are in particular causally complete, ${\cal O}=({\cal O}^{\perp\!\!\!\!\perp})^{\perp\!\!\!\!\perp}$. Finally, the family of open double cones is a topological basis for $\mathbb{M}$}. We leave the elementary proof of these geometric facts to the reader.
Proposition 6.14:
The Haag duality relation holds for the von Neumann algebras generated by the Weyl algebra of a real massive Klein-Gordon field in Minkowski spacetime: if ${\cal O}\subset{\mathbb{M}}$ is an open double cone, then
${\mathfrak{W}}({\cal O})^{\prime}={\mathfrak{W}}({\cal O}^{c})\>.$ (101)
Proof.
See Theorem 4.8 in [Cam07] for $M={\mathbb{M}}$. ∎
6.4 Commutativity of conditional POVMs as local AHK operators
We conclude this work, making particular use of Haag duality, by proving that the conditional POVMs with effects $B_{\Delta_{0}}(\Delta)^{u}_{f,\eta}$ satisfy the commutativity property expected in the AHK approach.
We remind the reader that a closed densely-defined operator $A$ is said to be affiliated with a von Neumann algebra ${\mathfrak{R}}\subset{\mathfrak{B}}({{\cal H}})$ [Tak02, StZs19] if $UA\subset AU$ for every unitary [^14] That is equivalent to the apparently stronger requirement $UAU^{\dagger}=A$ for every $\in{\mathfrak{R}}^{\prime}. UU\in{\mathfrak{R}}^{\prime}$. It turns out, by the very definition of von Neumann algebra and the fact that the unitaries of a von Neumann algebra generate the algebra, that if $A\in{\mathfrak{B}}({{\cal H}})$, affiliation with $\mathfrak{R}$} is equivalent to $A\in{\mathfrak{R}}$. It is not difficult to prove that, if $A:D(A)\to{{\cal H}}$ is selfadjoint, then $A$ is affiliated with $\mathfrak{R}$} if and only the projectors $P^{(A)}(E)$ of the PVM of $A$ are elements of $\mathfrak{R}$}.
We now proceed to prove that the closure of the stress-energy tensor operator is affiliated with every local von Neumann algebra associated with its smearing function. This requires a couple of lemmata.
Lemma 6.15:
If $f,f_{1}\in\mathscr{D}({\mathbb{M}})$, with $f$ real, satisfy $supp(f)\subset{\cal O}$, $supp(f_{1})\subset{\cal O}_{1}$ with the open sets $\cal O$} and ${\cal O}_{1}$ causally separated, then
  • (a)
    $W(f)\Psi\in D\left(\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\right)$ for $\Psi\in{\mathfrak{S}}_{0}$;
  • (b)
    $W(f)\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\subset\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}W(f)$.
The same results are valid if one replaces everywhere $\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]$ by $\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]+cI$ for any $c\in{\mathbb{C}}$.
Proof.
From (a) of Proposition 3.5, the vectors in ${\mathfrak{F}}_{0}$, and thus also those in the subspace ${\mathfrak{S}}_{0}$, are analytic vectors for the selfadjoint operators $\overline{\hat{\phi}(f)}$ when $f$ is real (see Theorem X.41 in [ReSa75] for a detailed proof); in particular, $W(tf)\Psi=\sum_{n=0}^{+\infty}\frac{i^{n}t^{n}\hat{\phi}[f]^{n}}{n!}\Psi$ for every $t\in{\mathbb{R}}$. Now observe that
$\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]\sum_{n=0}^{N}\frac{i^{n}\hat{\phi}[f]^{n}}{n!}\Psi=\sum_{n=0}^{N}\frac{i^{n}\hat{\phi}[f]^{n}}{n!}:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]\Psi$
where we used Proposition 4.5. Since $\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]\Psi\in{\mathfrak{S}}_{0}$ because $\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]({\mathfrak{S}}_{0})\subset{\mathfrak{S}}_{0}$ ((a) of Proposition 4.2), the limits of both sides exist as $N\to+\infty$, and this proves the first assertion because $\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]$ – which is defined on ${\mathfrak{S}}_{0}$ – is closable. In particular, we also find that $\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}W(f)\Psi=W(f)\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\Psi$ if $\Psi\in{\mathfrak{S}}_{0}$. To prove the second assertion, take $\Phi\in D\left(\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\right)$. There is a sequence ${\mathfrak{S}}_{0}\ni\Psi_{n}\to\Phi$ such that $\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]\Psi_{n}\to\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\Phi$ by definition of closure, so that $W(f)\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\Psi_{n}\to W(f)\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\Phi$. On the other hand, $W(f)\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\Psi_{n}=\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}W(f)\Psi_{n}$, and the limit of both sides exists by construction and thus must coincide with $\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}W(f)\Phi$, since the operator under consideration is closed. In summary, if $\Phi\in D\left(\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\right)$, then $\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}W(f)\Phi=W(f)\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{1}]}\Phi$. This is (b). The last statement is obvious if one observes that $\overline{A+cI}=\overline{A}+cI$ and $D(A+cI)=D(A)$, for a closable operator $A$ and any constant $c\in{\mathbb{C}}$. ∎
Lemma 6.16:
Consider an open double cone $\cal O$}. If $f\in\mathscr{D}({\mathbb{M}})$ and $supp(f)\subset{\cal O}$, then $\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]+cI}$ is affiliated with ${\mathfrak{W}}({\cal O})$ for any $c\in{\mathbb{C}}$.
Proof.
Let us start with the case $c=0$. A densely defined closed operator $A$ is affiliated with a von Neumann algebra $\mathfrak{R}$} if $UA\subset AU$ for every unitary $U\in{\mathfrak{R}}^{\prime}$. In our case we can exploit Haag duality, ${\mathfrak{W}}({\cal O})^{\prime}={\mathfrak{W}}({\cal O}^{c})$. A unitary $U\in{\mathfrak{W}}({\cal O})^{\prime}={\mathfrak{W}}({\cal O}^{c})$ is the strong limit of finite linear combinations of Weyl operators converging to $U$, with smearing functions supported in ${\cal O}^{c}$, due to (100). From Lemma 6.15, for $\Psi\in D\left(\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]}\right)$, it holds that $W_{n}\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]}\Psi=\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]}W_{n}\Psi$. Since $\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]}$ is closed and $W_{n}\to U$ strongly, taking the limit as $n\to+\infty$ we find the desired relation $U\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]}\Psi=\overline{:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]}U\Psi$. The case $c\neq 0$ is obvious once one observes that $\overline{A+cI}=\overline{A}+cI$ for a closable operator $A$ and any constant $c\in{\mathbb{C}}$. ∎
It is evident that this affiliation property with local von Neumann Weyl algebras is also valid for $\thinspace\hat{\phi}^{2}\thinspace:[f]$ and all elements of the local algebra ${\cal T}({\cal O})$ defined in Proposition 4.5. In principle, barring unexpected technicalities, the same argument should apply to all Wick polynomials defined through the Wick rule, whether normally ordered or constructed via the locally covariant Hadamard procedure (for the massive scalar field in Minkowski spacetime), including derivatives as well.
The final result on the commutativity of conditional POVMs, stated in (c) below, uses the previously proved lemmata and a crucial result by Kadison [Kad89].
Proposition 6.17 (Kadison):
[…]
Proof.
Corollary 5 in [Kad89]. ∎
We start with a physically obvious definition, which reflects the fact that the identity ${{\mathsf{H}}}^{u}_{f,\eta}(\Delta)=:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{\Delta}]u^{\mu}u^{\nu}_{\Sigma}+\eta|u\cdot u_{\Sigma}||\Delta|I$ holds, as follows from its definition and from (83), where, by construction, $supp(f_{\Delta})\subset\Delta+supp(f^{2})$. Thus the standard localization notion for local observables can also be used for ${{\mathsf{H}}}^{u}_{f,\eta}(\Delta)$.
Definition 6.18:
The operator ${\mathsf{H}}^{u}_{f,\eta}(\Delta)$ defined in (87) is localized in the open double cone $\cal O$} if $\Delta+supp(f^{2})\subset{\cal O}$. $\blacksquare$
Theorem 6.19:
Take $u\in{\mathsf{T}}_{+}$, an arbitrarily small $\eta>0$, and $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$ such that (90) is true. The following facts are true for a rest space $\Sigma$ and an open double cone ${\cal O}\subset{\mathbb{M}}$.
  • (a)
    If ${{\mathsf{H}}}^{u}_{f,\eta}(\Delta)$ is localized in $\cal O$}, then the Friedrichs extension $\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta)$ of ${{\mathsf{H}}}^{u}_{f,\eta}(\Delta)$ is affiliated with ${\mathfrak{W}}({\cal O})$. In particular the spectral projectors of $\hat{{\mathsf{H}}}^{u}_{f,\eta}$ belong to ${\mathfrak{W}}({\cal O})$.
  • (b)
    If ${{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})$ is localized in $\cal O$}, then the effects of the POVM $\mathscr{B}(\Delta_{0})\ni\Delta\mapsto B_{\Delta_{0}}(\Delta)^{u}_{f,\eta}$ defined in (95) satisfy $B_{\Delta_{0}}(\Delta)^{u}_{f,\eta}\in{\mathfrak{W}}({\cal O})$.
  • (c)
[…]
    $[{B}_{\tilde{\Delta}_{0}}(\tilde{\Delta})^{\tilde{u}}_{\tilde{f},\tilde{\eta}},B_{\Delta_{0}}(\Delta)^{u}_{f,\eta}]=0\>,\quad\mbox{for every $\Delta\subset\mathscr{B}(\Delta_{0})$ and $\tilde{\Delta}\subset\mathscr{B}(\tilde{\Delta}_{0})$}.$ (102)
Proof.
(a) First of all observe that the densely defined symmetric positive operator ${{\mathsf{H}}}^{u}_{f,\eta}$ satisfies ${{\mathsf{H}}}^{u}_{f,\eta}=:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{\Delta}]u^{\mu}u^{\nu}_{\Sigma}+\eta|u\cdot u_{\Sigma}||\Delta|I$ as follows from its definition and from (83). By construction, $supp(f_{\Delta})\subset\Delta+supp(f^{2})$. Furthermore, as follows from the definition of the Friedrichs extension, a symmetric positive operator and its closure have the same Friedrichs extension. At this point, Lemma 6.16 proves that $\overline{{{\mathsf{H}}}^{u}_{f,\eta}}$ is affiliated with ${\mathfrak{W}}({\cal O})$. Proposition 6.17 implies that the selfadjoint operator $\hat{{\mathsf{H}}}^{u}_{f,\eta}$ – the Friedrichs extension of (the closure of) ${\mathsf{H}}^{u}_{f,\eta}$ – is affiliated with ${\mathfrak{W}}({\cal O})$ as well and thus, in particular, the spectral projectors of $\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta)$ belong to ${\mathfrak{W}}({\cal O})$.
(b) Since ${\cal O}\supset\Delta_{0}+supp(f^{2})\supset\Delta+supp(f^{2})$ for $\Delta\subset\Delta_{0}$, (a) proves that the spectral projectors of all $\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta)$ and $\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})$ belong to ${\mathfrak{W}}({\cal O})$. Referring to Corollary 6.9, this fact easily implies, via spectral calculus (e.g. Proposition 3.78 in [Mor19]), that if $U\in{\mathfrak{W}}({\cal O})^{\prime}$ and $\Psi\in{\mathfrak{S}}_{00}$,
$U\frac{1}{\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta)\frac{1}{\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}\Psi=\frac{1}{\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta)\frac{1}{\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}U\Psi\>.$
Since the dense subspace ${\mathfrak{S}}_{00}$ is a core for the everywhere-defined bounded operator $B_{\Delta_{0}}(\Delta)^{u}_{f,\eta}=\overline{\frac{1}{\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta)\frac{1}{\sqrt{\hat{{\mathsf{H}}}^{u}_{f,\eta}(\Delta_{0})}}}$, we obtain that $UB_{\Delta_{0}}(\Delta)^{u}_{f,\eta}=B_{\Delta_{0}}(\Delta)^{u}_{f,\eta}U$ if $U\in{\mathfrak{W}}({\cal O})^{\prime}$, so that $B_{\Delta_{0}}(\Delta)^{u}_{f,\eta}\in{\mathfrak{W}}({\cal O})^{\prime\prime}={\mathfrak{W}}({\cal O})$.
(c) If $\cal O$} and $\tilde{\cal O}$ are causally separated, then $\tilde{\cal O}\subset{\cal O}^{c}$ so that ${\mathfrak{W}}(\tilde{\cal O})\subset{\mathfrak{W}}({\cal O}^{c})={\mathfrak{W}}({\cal O})^{\prime}$ by (101). The thesis follows from (b). ∎
Remark 6.20:
As already observed, the fact that $B_{\Delta_{0}}(\Delta)^{u}_{f,\eta}\in{\mathfrak{W}}({\cal O})$ in addition to positivity $B_{\Delta_{0}}(\Delta)^{u}_{f,\eta}\geq 0$, when we use that effect on states in the whole Hilbert space instead of those in ${{\cal H}}^{(1)}$, immediately implies a dark count phenomenon – which is mathematically expressed by $\langle\Omega|B_{\Delta_{0}}(\Delta)^{u}_{f,\eta}\Omega\rangle>0$ – due to the general version of the Reeh-Schlieder theorem (d) RS2 in Proposition 3.5. (See [FaCo26] for a recent physical discussion of the subject.) $\blacksquare$
7 Conclusions and outlook
In this work we have constructed a class of relativistic spatial localization observables $A^{u}_{f}$ within the standard framework of quantum field theory, by exploiting the stress–energy–momentum tensor operator smeared with test functions $f^{2}$. The construction provides, for every fixed timelike direction, a family of POVMs defined on spacelike hypersurfaces, which are well behaved on each $n$-particle sector and satisfy a natural relativistic causality condition of Castrigiano type.
A central outcome of the analysis is that these localization observables arise from local or quasi-local quantum field theoretic quantities, thereby placing previous constructions, originally formulated at a more heuristic level, on a rigorous footing. In particular, in the one-particle sector, the resulting localization scheme reduces to the observable introduced in [Mor23] (which extended and made rigorous an original physical model due to Terno [Ter14]), and its first moment reproduces the Newton–Wigner position operator, independently of the choice of smearing function under suitable normalization conditions.
An important aspect concerns the analysis of the role of energy positivity. While the stress–energy tensor fails to define positive operators on the full Fock space due to the Reeh–Schlieder theorem, we have shown that suitable lower bounds can be established by means of quantum energy inequalities. This analysis makes it possible to control deviations from positivity and to define regularized families of positive operators $A^{u}_{f,\epsilon}$ that approximate the localization effects with arbitrary precision on states with a fixed particle number. The price to pay is, of course, that the operators $A^{u}_{f,\epsilon}$ are non-local; however, they arise as restrictions of non-local and non-positive operators ${\mathsf{A}}^{u}_{f,\epsilon}$ acting on the full Fock space.
The second main result of the paper is the construction of conditional localization observables associated with finite laboratories. Taking energy inequalities into account once again, by introducing suitably modified local energy operators and considering their Friedrichs selfadjoint extensions, we have defined conditional POVMs. These POVMs are related in a natural way to the spatial localization observables constructed above. However, certain intertwining unitary operators $V^{u}_{f,\epsilon,\eta,\Delta_{0}}$ appear in the relation between the local effects and the (approximated and positive) relativistic spatial localization effects ${\mathsf{A}}^{u}_{f,\epsilon,\eta}$, as in (98). The role of these unitary operators, although compatible with the analysis developed in [Mor26], is not yet fully understood and deserves further investigation.
Within this framework, and under appropriate causal separation assumptions on the laboratories, we have shown that the effects of conditional localization observables commute and belong to local von Neumann algebras, in agreement with the Araki–Haag–Kastler description of locality and in accordance with the analysis in [Mor26]. This provides a concrete realization, in a quantum field theoretic setting, of the idea that commutativity of localization observables should be recovered only at the level of conditional measurements performed in spacetime regions of finite extent.
Several directions for further investigation naturally emerge from the present analysis. From a mathematical viewpoint, it would be desirable to extend the construction beyond Minkowski spacetime, in particular to globally hyperbolic curved spacetimes and Hadamard states, where the stress–energy tensor and quantum energy inequalities are available in a suitably generalized form. Another relevant issue concerns the dependence on the smearing function and the extent to which different choices lead to physically equivalent localization schemes.
On the physical side, a more detailed analysis of measurement procedures implementing the proposed observables would be of interest, especially in connection with indirect measurement models and detector-based formulations.
Finally, the interplay between localization, energy conditions, and causality constraints suggests that the framework developed here may be useful in addressing more general questions concerning the operational meaning of localization in relativistic quantum systems and its compatibility with the structure of local quantum field theory.
There remain, however, some open issues that deserve further investigation.
Castrigiano (see especially [Cas17, Cas24]) introduced several relativistic spatial localization observables for bosons and fermions which appear to arise as restrictions to the one-particle space of a more general structure defined on the full Fock space. In that framework, the relevant local observables seem to be given by the electric current rather than the stress–energy tensor. However, it is not clear how a smearing procedure applied to such local observables at the level of the Fock space could reproduce, after spatial integration, the corresponding localization effects. Clarifying this point would be of particular interest and will be addressed elsewhere.
Another related issue concerns the possible existence of connections with the localization observables introduced by Lechner and de Oliveira [LedO26], where modular theory plays a central role. Understanding whether, and in which sense, the present construction can be related to that approach may shed further light on the structural aspects of localization in quantum field theory.
Acknowledgments
The author is grateful to D. Castrigiano, C. De Rosa, C. Fewster, N. Pinamonti, and A. Schenkel for many useful discussions, also over the years, on the issues addressed in this work. This work has been written within the activities of INdAM-GNFM.
Appendix
Appendix A Appendix: Properties of normally ordered quadratic forms
Proposition A.1:
The normally ordered quadratic forms as in Definition 3.6 satisfy the following elementary properties.
  • (1)
    Directly from the definition,
    $\left\langle\Psi^{\prime}\left|\prod_{j=1}^{N}a^{\dagger}_{p_{j}}\prod_{r=1}^{M}a_{k_{r}}\Psi\right.\right\rangle=\overline{\left\langle\Psi\left|\prod_{r=1}^{M}a^{\dagger}_{k_{r}}\prod_{j=1}^{N}a_{p_{j}}\Psi^{\prime}\right.\right\rangle}\>$. (103)
  • (2)
    If $\sigma:\{1,\ldots,N\}\to\{1,\ldots,N\}$ and $\pi:\{1,\ldots,M\}\to\{1,\ldots,M\}$ are arbitrary permutations (i.e., bijective functions):
    $\left\langle\Psi^{\prime}\left|\prod_{j=1}^{N}a^{\dagger}_{p_{j}}\prod_{r=1}^{M}a_{k_{r}}\Psi\right.\right\rangle=\left\langle\Psi^{\prime}\left|\prod_{j=1}^{N}a^{\dagger}_{p_{\sigma(j)}}\prod_{r=1}^{M}a_{k_{\pi(n)}}\Psi\right.\right\rangle\>$. (104)
  • (3)
    Taking (10) and (9) into account: for $(\Lambda,a)\in IO(1,3)_{+}$,
    $\left\langle U^{-1}_{(\Lambda,a)}\Psi^{\prime}\left|\prod_{j=1}^{N}a^{\dagger}_{p_{j}}\prod_{r=1}^{M}a_{k_{r}}U^{-1}_{(\Lambda,a)}\Psi\right.\right\rangle=\left\langle\Psi^{\prime}\left|\prod_{j=1}^{N}e^{-ia\cdot p_{j}}a^{\dagger}_{\Lambda p_{j}}\prod_{r=1}^{M}e^{ia\cdot k_{r}}a_{\Lambda k_{r}}\Psi\right.\right\rangle\>$. (105)
  • (4)
    If $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$, defining $K:=3N+3M$, the map
    $\mathsf{V}_{m,+}^{3N+3M}\ni(p_{1},\ldots,p_{N},k_{1},\ldots,k_{M})\mapsto\left\langle\Psi^{\prime}\left|\prod_{j=1}^{N}a^{\dagger}_{p_{j}}\prod_{r=1}^{M}a_{k_{r}}\Psi\right.\right\rangle\in{\mathbb{C}}\>\>\mbox{is $\mathscr{S}(\mathsf{V}_{m,+}^{K})$}$ (106)
  • (5)
    From (25), for $f_{j},g_{r}\in\mathscr{S}(\mathsf{V}_{m,+})$, $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$:
[…]
    $=\int\prod_{j=1}^{N}f_{j}(p_{j})\prod_{r=1}^{M}\overline{g_{r}(k_{j})}\left\langle\Psi^{\prime}\left|\prod_{j=1}^{N}a^{\dagger}_{p_{j}}\prod_{r=1}^{M}a_{k_{r}}\Psi\right.\right\rangle\;d^{N}\mu_{m}(p)d^{M}\mu_{m}(k)$ (107)
  • (6)
    More generally, if $h:\mathsf{V}_{m,+}^{3N+3M}\to{\mathbb{C}}$ is measurable and polynomially bounded, then the quadratic [^15] Theorem X.44 of [ReSa75] establishes that a quadratic form as in (108) uniquely defines a closable operator if $h\in L^{2}$. However, we shall need a more general type of quadratic form which does not satisfy the hypotheses of that theorem. form on ${\mathfrak{S}}_{0}$
    ${\mathfrak{S}}_{0}\times{\mathfrak{S}}_{0}\ni(\Psi^{\prime},\Psi)\mapsto$
    $\int h(k_{1},\ldots,k_{N},p_{1},\ldots,p_{M})\left\langle\Psi^{\prime}\left|\prod_{j=1}^{N}a^{\dagger}_{p_{j}}\prod_{r=1}^{M}a_{k_{r}}\Psi\right.\right\rangle\;d^{N}\mu_{m}(p)d^{M}\mu_{m}(k)\>.$ (108)
  • (7)
    By direct inspection and referring to Lemma 3.1, if $v\in\mathsf{V}$ and $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$, then
    $\langle\Psi|H^{v}\Psi^{\prime}\rangle=-\int_{\mathsf{V}^{+}_{m}}v\cdot p\langle\Psi|a^{\dagger}_{p}a_{p}\Psi^{\prime}\rangle d\mu_{m}(p)\>.$ (109)
Appendix B Appendix: Proofs of technical propositions
Proof of Proposition 4.2. First statement and (a). We notice that, if operators $\thinspace\hat{\phi}^{2}\thinspace:[f]:{\mathfrak{S}}_{0}\to{\mathfrak{F}}_{s}({{\cal H}}_{m})$, $\thinspace\hat{T}_{\mu\nu}\thinspace:[f]:{\mathfrak{S}}_{0}\to{\mathfrak{F}}_{s}({{\cal H}}_{m})$ exist such that (38) and (39) hold, i.e., they define the corresponding quadratic forms on ${\mathfrak{S}}_{0}$, then they must be unique, since ${\mathfrak{S}}_{0}$ is dense. We choose a Minkowskian reference frame, which we shall use henceforth. We indicate by $\hat{f}=\hat{f}(k^{0},\vec{k})$ the Fourier transform (18) of $f\in\mathscr{S}_{\mathbb{R}}({\mathbb{R}}^{4})$. We want to associate a corresponding operator with each addend in the decomposition (36), whose sum amounts to $\thinspace\hat{T}_{\mu\nu}\thinspace:[f]$. We start with the definition of an operator representing
$\thinspace\hat{T}_{\mu\nu}\thinspace:[f]_{aa}=(\mbox{formally})=\int\left(\int e^{i(p+k)\cdot x}f(x)d^{4}x\right)\frac{a_{p}a_{k}}{2(2\pi)^{3}}t_{\mu\nu}(p,-k)d^{2}\mu_{m}(p,k)$
$\int d^{2}\mu_{m}(p,k)\>\frac{\hat{f}(E(p)+E(k),\vec{p}+\vec{k})}{4\pi}t_{\mu\nu}(p,-k)a_{p}a_{k}\>.$
Notice that the smooth function
$\mathsf{V}_{m,+}^{2}\ni(p,k)\mapsto h_{\mu\nu}(p,k):=\frac{\hat{f}(E(p)+E(k),\vec{p}+\vec{k})}{4\pi}t_{\mu\nu}(p,-k)\in{\mathbb{C}}$
is a function in $\mathscr{S}(\mathsf{V}_{m,+}^{2})$, since, for instance, as $\hat{f}\in\mathscr{S}({\mathbb{R}}^{4})$ and $t_{\mu\nu}(p,k)$ is polynomially bounded,
$|h_{\mu\nu}(p,k)|\leq C_{N}((\vec{p}+\vec{k})^{2}+(E(p)+E(k))^{2})^{-N/2}\leq C^{\prime}_{N}((E(p)+E(k))^{2})^{-N/2}\leq C^{\prime\prime}_{N}(\vec{p}^{2}+\vec{k}^{2})^{-N/2}$
for every $N\in{\mathbb{N}}$. Furthermore, the same argument is valid for every derivative, because $\hat{f}\in\mathscr{S}({\mathbb{R}}^{4})$ and the derivatives of the smooth functions $E(p)$, $E(k)$, $t_{\mu\nu}(p,-k)$ are polynomially bounded. Let us define $\thinspace\hat{T}_{\mu\nu}\thinspace:[f]_{aa}:={\mathfrak{S}}_{0}\to{\mathfrak{S}}_{0}$ as the linear extension of the map ${{\cal H}}^{(n)}\cap{\mathfrak{S}}_{0}\to{{\cal H}}^{(n-2)}\cap{\mathfrak{S}}_{0}$, which vanishes if $n<2$, and where we use the notation $d^{2}\mu_{m}(p):=d^{2}\mu_{m}(p_{1},p_{2})$
$(:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]_{aa}\Psi)({k}_{1},\ldots,{k}_{n-2}):=\sqrt{n(n-1)}\int\thinspace h_{\mu\nu}({p}_{1},{p}_{2})\Psi({p}_{1},{p}_{2},{k}_{1},\ldots,{k}_{n-2})d^{2}\mu_{m}(p).$ (110)
This is a well-defined operator ${\mathfrak{S}}_{0}\to{\mathfrak{S}}_{0}$ with the property that, if $f\in\mathscr{D}({\mathbb{M}})$, a direct use of Fubini-Tonelli’s theorem yields
$\langle\Psi^{\prime}|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]_{aa}\Psi\rangle=\int\left.\left(\int e^{i(p+k)\cdot x}f(x)d^{4}x\right)\right|_{\mathsf{V}_{m,+}^{2}}\thinspace\frac{\langle\Psi^{\prime}|a_{p}a_{k}\Psi\rangle}{2(2\pi)^{3}}t_{\mu\nu}(p,-k)d^{2}\mu_{m}(p,k)\>.$
Analogously, referring to the complex-conjugated function $\overline{h}_{\mu\nu}\in\mathscr{S}(\mathsf{V}_{m,+}^{2})$ we can define $\thinspace\hat{T}_{\mu\nu}\thinspace:[f]_{a^{\dagger}a^{\dagger}}:={\mathfrak{S}}_{0}\to{\mathfrak{S}}_{0}$ as the linear extension of the map ${{\cal H}}^{(n)}\cap{\mathfrak{S}}_{0}\to{{\cal H}}^{(n+2)}\cap{\mathfrak{S}}_{0}$
$(:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]_{a^{\dagger}a^{\dagger}}\Psi)({k}_{1},\ldots,{k}_{n+2}):=\sqrt{(n+1)(n+2)}S_{n+2}\left(\overline{h_{\mu\nu}(k_{1},k_{2})}\Psi(k_{3},\ldots,k_{n+2})\right).$ (111)
$S_{n}:{{\cal H}}^{n}\to{{\cal H}}^{(n)}$ is the symmetrization projector. With this definition we have, as before, that if $f\in\mathscr{D}({\mathbb{M}})$
$\langle\Psi^{\prime}|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]_{a^{\dagger}a^{\dagger}}\Psi\rangle=\int\left.\left(\int e^{-i(p+k)\cdot x}f(x)d^{4}x\right)\right|_{\mathsf{V}_{m,+}^{2}}\thinspace\frac{\langle\Psi^{\prime}|a^{\dagger}_{p}a^{\dagger}_{k}\Psi\rangle}{2(2\pi)^{3}}t_{\mu\nu}(p,-k)d^{2}\mu_{m}(p,k)\>.$
We now have to construct the operators associated with the operators $a_{p}^{\dagger}a_{k}$ and $a^{\dagger}_{k}a_{p}$ in the quadratic form of the stress-energy tensor. Let us start by defining the smooth function
$\mathsf{V}_{m,+}^{2}\ni(p,k)\mapsto s_{\mu\nu}^{+}(p,k):=\frac{\hat{f}(E(p)-E(k),\vec{p}-\vec{k})}{4\pi}t_{\mu\nu}(p,k)\in{\mathbb{C}}\>,$
which, contrary to $h$, is polynomially bounded but not necessarily in $\mathscr{S}(\mathsf{V}_{m,+}^{2})$, due to the minus sign in the temporal entry of $\hat{f}$. Nevertheless the map $(k_{1},\ldots,k_{n})\mapsto\int_{{\mathbb{R}}^{3}}\thinspace s^{+}_{\mu\nu}(k_{1},p)\Psi(p,{k}_{2},\ldots,\vec{k}_{n})d\mu_{m}(p)$ is in $\mathscr{S}(\mathsf{V}_{m,+}^{l})$ as proved in Lemma B.1 below. The wanted operator is therefore the linear extension of the operator ${{\cal H}}^{(n)}\cap{\mathfrak{S}}_{0}\to{{\cal H}}^{(n)}\cap{\mathfrak{S}}_{0}$ such that it vanishes for $n=0$ and
$(:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]^{+}_{a^{\dagger}a}\Psi)({k}_{1},\ldots,{k}_{n}):=nS_{n}\left(\int_{{\mathbb{R}}^{3}}\thinspace s^{+}_{\mu\nu}(k_{1},p)\Psi(p,{k}_{2},\ldots,\vec{k}_{n})d\mu_{m}(p)\right)\>.$ (112)
This is a well-defined operator ${\mathfrak{S}}_{0}\to{\mathfrak{S}}_{0}$ with the property that
$\langle\Psi^{\prime}|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]^{+}_{a^{\dagger}a}\Psi\rangle=\int\left.\left(\int e^{i(k-p)\cdot x}f(x)d^{4}x\right)\right|_{\mathsf{V}_{m,+}^{2}}\thinspace\frac{\langle\Psi^{\prime}|a^{\dagger}_{p}a_{k}\Psi\rangle}{2(2\pi)^{3}}t_{\mu\nu}(p,k)d^{2}\mu_{m}(p,k)\>.$
With the same procedure, defining
$\mathsf{V}_{m,+}^{2}\ni(p,k)\mapsto s^{-}_{\mu\nu}(p,k):=\frac{\hat{f}(E(k)-E(p),\vec{k}-\vec{p})}{4\pi}t_{\mu\nu}(p,k)\in{\mathbb{C}}\>,$
and
$(:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]^{-}_{a^{\dagger}a}\Psi)_{n}({k}_{1},\ldots,{k}_{n}):=nS_{n}\left(\int_{{\mathbb{R}}^{3}}\thinspace s^{-}_{\mu\nu}(k_{1},p)\Psi(p,{k}_{2},\ldots,\vec{k}_{n})d\mu_{m}(p)\right)$ (113)
we find
$\langle\Psi^{\prime}|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]^{-}_{a^{\dagger}a}\Psi\rangle=\int\left.\left(\int e^{-i(k-p)\cdot x}f(x)d^{4}x\right)\right|_{\mathsf{V}_{m,+}^{2}}\thinspace\frac{\langle\Psi^{\prime}|a^{\dagger}_{k}a_{p}\Psi\rangle}{2(2\pi)^{3}}t_{\mu\nu}(p,k)d^{2}\mu_{m}(p,k)$
$\langle\Psi^{\prime}|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]^{+}_{a^{\dagger}a}\Psi\rangle\>.$
The wanted operator, which also satisfies (a) by construction, is therefore
$:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]:=:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]_{aa}+:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]_{a^{\dagger}a^{\dagger}}+2:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]^{+}_{a^{\dagger}a}\>$. (114)
The case of $\phi^{2}$ is essentially identical: it is sufficient to remove the tensor $t_{\mu\nu}$ from each step of the above proof. The final statement is obvious from the fact that the quadratic form of the stress-energy tensor is a smooth function of $x$ and that (38) and (39) hold.
(b) The identities
$\overline{\langle\Psi^{\prime}|:\thinspace\hat{\phi}^{2}\thinspace:(x)\Psi\rangle}=\langle\Psi|:\thinspace\hat{\phi}^{2}\thinspace:(x)\Psi^{\prime}\rangle\>,\quad\overline{\langle\Psi^{\prime}|:\thinspace\hat{T}_{\mu\nu}\thinspace:(x)\Psi\rangle}=\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:(x)\Psi^{\prime}\rangle$
and consequently easily follow from (35) and (36), taking (28) into account. At this juncture, (39) proves that the associated operator satisfies $\thinspace\hat{T}_{\mu\nu}\thinspace:[\overline{f}]\subset:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]^{\dagger}$ since the domain of $\thinspace\hat{T}_{\mu\nu}\thinspace:[f]$ is dense and thus the operator $\thinspace\hat{T}_{\mu\nu}\thinspace:[f]$ admits an adjoint. The case of $\phi^{2}$ is identical. As an immediate consequence of what has been established, $\thinspace\hat{T}_{\mu\nu}\thinspace:[f]$ and $\thinspace\hat{\phi}^{2}\thinspace:[f]$ are symmetric if $f$ is real.
(c) From (36), since $t_{\mu\nu}$ is symmetric we have $\langle\Psi^{\prime}|:\thinspace\hat{T}_{\mu\nu}\thinspace:(x)\Psi\rangle=\langle\Psi^{\prime}|:\thinspace\hat{T}_{\nu\mu}\thinspace:(x)\Psi\rangle$ for every $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$ and $x\in{\mathbb{M}}$. At this juncture, the wanted identity follows from (39) using density of the domain.
(d) Starting from the decomposition (36), taking (105) into account, and using $IO(1,3)_{+}$-invariance of the scalar product and the measure on the mass shell, we find that it holds $\langle U^{-1}_{g}\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:(x)U_{g}^{\dagger}\Psi^{\prime}\rangle={(\Lambda^{-1})_{\mu}}^{\alpha}{(\Lambda^{-1})_{\nu}}^{\beta}\thinspace\langle\Psi|:\thinspace\hat{T}_{\alpha\beta}\thinspace:(gx)\Psi\rangle$ for $x\in{\mathbb{M}}\>,\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$ At this juncture (39) gives (40) as a consequence of domain density and Poincaré invariance of the measure of $\mathbb{M}$}.
(e) It is sufficient to prove that $\partial^{\mu}\langle\Psi^{\prime}|:\thinspace\hat{T}_{\mu\nu}\thinspace:(x)\Psi\rangle=0$ for $x\in{\mathbb{M}},\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$. The thesis then immediately arises from (39) by integrating by parts, using the fact that the quadratic form is a bounded function in spacetime. Concerning the above identity, which is actually equivalent to the thesis, it is an immediate consequence of (36) when passing the derivative under the symbol of integration (which is permitted by the rapid decay of the integrated functions and their derivatives, exploiting the dominated convergence theorem) and noticing that, on shell, $(p+k)^{\mu}t_{\mu\nu}(p,-k)=0$ and $(p-k)^{\mu}t_{\mu\nu}(p,k)=0$. The proof is over. $\Box$
Lemma B.1:
Let $E(q):=\sqrt{\vec{q}^{2}+m^{2}}$ for a fixed constant $m>0$, consider $f\in\mathscr{S}({\mathbb{R}}^{4})$, and define
$h(\vec{k},\vec{p}):=\int_{{\mathbb{R}}^{4}}e^{i(\vec{k}-\vec{p})\cdot\vec{x}}e^{-i(E(k)-E(p))x^{0}}f(x^{0},\vec{x})d^{4}x\>,\quad T_{\mu\nu}(\vec{k},\vec{p}):=t_{\mu\nu}((E(k),\vec{k}),(E(p),\vec{p}))\>,$
where
$t_{\mu\nu}(p,k):=\frac{1}{2}[k_{\mu}p_{\nu}+p_{\mu}k_{\nu}-g_{\mu\nu}(p\cdot k+m^{2})]\>.$
If $\psi=\psi(\vec{k},K)$, with $K:=(\vec{k}_{1},\ldots,\vec{k}_{N})$, is a function in $\mathscr{S}({\mathbb{R}}^{3(N+1)})$, $N\geq 0$, then the function
$\phi(\vec{k},K):=\int_{{\mathbb{R}}^{3}}h(\vec{k},\vec{p})T_{\mu\nu}(\vec{k},\vec{p})\psi(\vec{p},K)\frac{d^{3}p}{E(p)}$
belongs to $\mathscr{S}({\mathbb{R}}^{3(N+1)})$.
Proof.
It is easy to see that $\phi$ is smooth as a consequence of the mean value theorem and the dominated convergence theorem when computing the $\vec{k}$ and $K$ derivatives of any order and passing them under the symbol of integration, since these derivatives are locally uniformly bounded by integrable functions of the variable $\vec{p}$ only. Furthermore, taking advantage of Fubini-Tonelli’s theorem,
$\phi(\vec{k},K)=\int_{{\mathbb{R}}^{3}}\int_{{\mathbb{R}}^{4}}e^{i(\vec{k}-\vec{p})\cdot\vec{x}}e^{-i(E(k)-E(p))x^{0}}f(x^{0},\vec{x}){T}_{\mu\nu}(\vec{k},\vec{p})\frac{\psi(\vec{p},K)}{E(p)}d^{4}xd^{3}p\>.$
For any choice of the indices $\mu,\nu$, $\phi$ can be rewritten as
$\phi(\vec{k},K)=\int_{{\mathbb{R}}^{3}}\int_{{\mathbb{R}}^{4}}e^{i(\vec{k}-\vec{p})\cdot\vec{x}}e^{-i(E(k)-E(p))x^{0}}f^{\prime}(x^{0},\vec{x})\psi^{\prime}(\vec{p},K)d^{4}xd^{3}p\>,$ (115)
where the terms of $\hat{T}_{\mu\nu}$ have been embodied in the new Schwartz functions $f^{\prime}$ and $\psi^{\prime}$ by integrating by parts in the variables $\vec{x}$ and $x^{0}$. These functions depend on the choice of the indices $\mu,\nu$. The new function $\psi^{\prime}$ also embodies the factor $E(p)^{-1}$.
At this juncture the thesis is equivalent to showing that
$|k^{\alpha}\partial^{|\alpha|}_{k^{\alpha}}K^{\beta}\partial^{|\beta|}_{K^{\beta}}\phi(\vec{k},K)|\leq C_{\alpha\beta}\quad\mbox{for}\quad(\vec{k},K)\in{\mathbb{R}}^{3(N+1)}$ (116)
for every choice of the multiindices $\alpha,\beta$ and associated constants $C_{\alpha\beta}<+\infty$. Concerning the operators of type $K^{\beta}\partial^{|\beta|}_{K^{\beta}}$, passing the operator under the integral symbol, their action is simply to change $\psi^{\prime}$ to a different function in $\mathscr{S}({\mathbb{R}}^{3(N+1)})$. Therefore we focus attention on the operators $k^{\alpha}\partial^{|\alpha|}_{k^{\alpha}}$.
The action of an operator $\partial^{|\alpha|}_{k^{\alpha}}$ on $\phi$ produces a linear combination of functions as in (115) with the following changes in the integrand: (a) factors $x^{\beta}$, (b) factors given by products of $k$-derivatives (of order $1$ at least) of the function $E(k)$. The terms of type (a) can be accommodated into a new definition of the Schwartz function $f^{\prime}$. We end up with a linear combination of functions
$Q(\vec{k})\int_{{\mathbb{R}}^{3}}\int_{{\mathbb{R}}^{4}}e^{i(\vec{k}-\vec{p})\cdot\vec{x}}e^{-i(E(k)-E(p))x^{0}}f^{\prime\prime}(x^{0},\vec{x})\psi^{\prime}(\vec{p},K)d^{4}xd^{3}p$
where $Q(\vec{k})$ is a polynomial in $k$-derivatives (of first order at least) of the function $E(k)$. These derivatives, and thus the polynomial itself, are bounded by some $c<+\infty$ as it is easy to prove by direct inspection. The action of the multiplicative operator $k^{\alpha}$ in the components of $\vec{k}$ can be worked out by integrating by parts, obtaining factors $x^{\gamma}$ which can be embodied in a new function $f^{\prime\prime}\in\mathscr{S}({\mathbb{R}}^{4})$ and factors $p^{\omega}$ which can be included in a new function $\psi^{\prime\prime}\in\mathscr{S}({\mathbb{R}}^{3(N+1)})$. In summary
$k^{\alpha}\partial^{|\alpha|}_{k^{\alpha}}K^{\beta}\partial^{|\beta|}_{K^{\beta}}\phi(\vec{k},K)=\sum_{j=1}^{J}Q_{j}(\vec{k})\int_{{\mathbb{R}}^{3}}\int_{{\mathbb{R}}^{4}}e^{i(\vec{k}-\vec{p})\cdot\vec{x}}e^{-i(E(k)-E(p))x^{0}}f_{j}(x^{0},\vec{x})\psi_{j}(\vec{p},K)d^{4}xd^{3}p\>,$
so that, if $|Q_{j}(\vec{k})|\leq c_{j}<+\infty$ as said above,
$|k^{\alpha}\partial^{|\alpha|}_{k^{\alpha}}K^{\beta}\partial^{|\beta|}_{K^{\beta}}\phi(\vec{k},K)|\leq\sum_{j=1}^{J}c_{J}\int_{{\mathbb{R}}^{3}}\int_{{\mathbb{R}}^{4}}|f_{j}(x^{0},\vec{x})\psi_{j}(\vec{p},K)|d^{4}xd^{3}p<+\infty$
concluding the proof. ∎
Proof of Lemma 4.3. It is sufficient to prove that the matrix elements of the two sides of (44) computed on elements $\Psi_{n}\in{{\cal H}}^{(n)}$, $\Psi_{l}\in{{\cal H}}^{(l)}$ are identical. This identity amounts to proving that $\langle\Psi_{n}|a(\kappa_{m}\overline{f})a^{\dagger}(\kappa_{m}g)\Psi_{l}\rangle=\langle\Omega|a(\kappa_{m}\overline{f})a^{\dagger}(\kappa_{m}g)\Omega\rangle\langle\Psi_{n}|\Psi_{l}\rangle$. This identity is true as an easy consequence of the definitions of $a(\psi)$ and $a^{\dagger}(\psi)$. The last identity is now obvious. $\Box$
Proof of Proposition 4.4. According to Propositions 3.5 and 3.9, the map $\mathscr{D}({\mathbb{M}})\ni g\mapsto\langle\Psi|:\thinspace\hat{\phi}[f]\hat{\phi}[g]\thinspace:\Psi^{\prime}\rangle=\langle\hat{\phi}[\overline{f}]\Psi|\hat{\phi}[g]\Psi^{\prime}\rangle-\langle\hat{\phi}[\overline{f}]\Omega|\hat{\phi}[g]\Omega\rangle\langle\Psi|\Psi^{\prime}\rangle$ is continuous for every given $f\in\mathscr{D}({\mathbb{M}})$. With a similar argument, $\langle\Psi|:\thinspace\hat{\phi}[f_{n}]\hat{\phi}[g]\thinspace:\Psi^{\prime}\rangle\to 0$ if $f_{n}\to 0$ in $\mathscr{D}({\mathbb{M}})$ for every given $g\in\mathscr{D}({\mathbb{M}})$. Schwartz’ kernel theorem implies that there exists a unique distribution in $\mathscr{D}^{\prime}({\mathbb{M}}\times{\mathbb{M}})$ which extends the map $\mathscr{D}({\mathbb{M}})\times\mathscr{D}({\mathbb{M}})\ni(f,g)\mapsto\langle\Psi|:\thinspace\hat{\phi}[f]\hat{\phi}[g]\thinspace:\Psi^{\prime}\rangle$. We denote its kernel by $\langle\Psi|:\thinspace\hat{\phi}(x)\hat{\phi}(y)\thinspace:\Psi^{\prime}\rangle$. At this juncture, using the definition (19), (44), and (107) one finds that the bounded smooth $x$-$y$ symmetric function (as the quadratic forms $(p,k)\mapsto\langle\Psi^{\prime}|a^{\#}_{p}a^{\#}_{k}\Psi\rangle$ are Schwartz functions on $\mathsf{V}_{m,+}^{2}$)
$K(x,y):=\int\frac{e^{i(p\cdot x+k\cdot y)}\langle\Psi^{\prime}|a_{p}a_{k}\Psi\rangle}{2(2\pi)^{3}}d^{2}\mu_{m}(p,k)+\int\frac{e^{-i(p\cdot x+k\cdot y)\cdot x}\langle\Psi^{\prime}|a^{\dagger}_{p}a^{\dagger}_{k}\Psi\rangle}{2(2\pi)^{3}}d^{2}\mu_{m}(p,k)$
$+\int e^{i(k\cdot y-p\cdot x)}\frac{\langle\Psi^{\prime}|a^{\dagger}_{p}a_{k}\Psi\rangle}{2(2\pi)^{3}}d^{2}\mu_{m}(p,k)+\int e^{i(p\cdot x-k\cdot y)}\frac{\langle\Psi^{\prime}|a^{\dagger}_{k}a_{p}\Psi\rangle}{2(2\pi)^{3}}d^{2}\mu_{m}(p,k)\>,$ (117)
satisfies
$\int f(x)g(y)K(x,y)d^{4}xd^{4}y=\langle\Psi|:\thinspace\hat{\phi}[f]\hat{\phi}[g]\thinspace:\Psi^{\prime}\rangle\>.$
By uniqueness, $\langle\Psi|:\thinspace\hat{\phi}(x)\hat{\phi}(y)\thinspace:\Psi^{\prime}\rangle=K(x,y)$, which is a smooth bounded function, so that it can be smeared with a distribution with compact support such as $f(x)\delta(x,y)$ for $f\in\mathscr{D}({\mathbb{M}})$. At this juncture (45) immediately arises by comparing the action of this distribution on $K(x,y)$ and (35). A straightforward modification of this argument yields (46) from (36) when using a distribution $D^{*}_{\mu,\nu}(x,y)f(x)\delta(x,y)=D_{\mu,\nu}(x,y)f(x)\delta(x,y)$. $\Box$
Proof of Proposition 4.5. We start by proving the commutativity for two square fields $\thinspace\hat{\phi}^{2}\thinspace:[\overline{h}]$. By direct use of (c3) in Proposition 3.5, we have that
$\hat{\phi}[h_{1}]\hat{\phi}[h_{2}],\hat{\phi}[h^{\prime}_{1}]\hat{\phi}[h^{\prime}_{2}]]=0$
if $supp(h_{i})\subset{\cal O}_{1}$ and $supp(h^{\prime}_{i})\subset{\cal O}_{2}$. Definition (43) immediately implies that also
$\thinspace\hat{\phi}[h_{1}]\hat{\phi}[h_{2}]\thinspace:,:\thinspace\hat{\phi}[h^{\prime}_{1}]\hat{\phi}[h^{\prime}_{2}]\thinspace:]=0$
so that
$\langle:\thinspace\hat{\phi}[\overline{h_{2}}]\hat{\phi}[\overline{h_{1}}]\thinspace:\Psi|:\thinspace\hat{\phi}[h^{\prime}_{1}]\hat{\phi}[h^{\prime}_{2}]\thinspace:\Psi^{\prime}\rangle-\overline{\langle:\thinspace\hat{\phi}[h_{1}]\hat{\phi}[h_{2}]\thinspace:\Psi^{\prime}|:\thinspace\hat{\phi}[\overline{h^{\prime}_{2}}]\hat{\phi}[\overline{h^{\prime}_{1}}]\thinspace:\Psi\rangle}=0$
Proposition 4.4 finally entails
$\int(\langle:\thinspace\hat{\phi}[\overline{h_{2}}]\hat{\phi}[\overline{h_{1}}]\thinspace:\Psi|:\thinspace\hat{\phi}(x)\hat{\phi}(y)\thinspace:\Psi^{\prime}\rangle-\overline{\langle:\thinspace\hat{\phi}[h_{1}]\hat{\phi}[h_{2}]\thinspace:\Psi^{\prime}|:\thinspace\hat{\phi}(y)\hat{\phi}(x)\thinspace:\Psi\rangle})h_{1}^{\prime}(x)h^{\prime}_{2}(y)d^{4}xd^{4}y=0$
A standard argument based on the continuity of the function before $h_{1}^{\prime}(x)h^{\prime}_{2}(y)$ in the integrand, arbitrariness of the functions $h^{\prime}_{1},h^{\prime}_{2}\in\mathscr{D}({\cal O}_{2})$ and the product topology of ${\cal O}_{2}\times{\cal O}_{2}$ yields
$\langle:\thinspace\hat{\phi}[\overline{h_{2}}]\hat{\phi}[\overline{h_{1}}]\thinspace:\Psi|:\thinspace\hat{\phi}(x)\hat{\phi}(y)\thinspace:\Psi^{\prime}\rangle-\overline{\langle:\thinspace\hat{\phi}[h_{1}]\hat{\phi}[h_{2}]\thinspace:\Psi^{\prime}|:\thinspace\hat{\phi}(y)\hat{\phi}(x)\thinspace:\Psi\rangle}=0\quad\mbox{if $(x,y)\in{\cal O}_{2}\times{\cal O}_{2}$.}$
As the function on the left-hand side is smooth, smearing both sides with $h^{\prime}(x)\delta(x,y)$ where $supp(h^{\prime})\subset B$ produces
$\langle:\thinspace\hat{\phi}[\overline{h_{1}}]\hat{\phi}[\overline{h_{2}}]\thinspace:\Psi|:\thinspace\hat{\phi}^{2}\thinspace:[h^{\prime}]\Psi^{\prime}\rangle-\overline{\langle:\thinspace\hat{\phi}[h_{1}]\hat{\phi}[h_{2}]\thinspace:\Psi|:\thinspace\hat{\phi}^{2}\thinspace:[\overline{h^{\prime}}]\Psi^{\prime}\rangle}=0\>.$
Taking the complex conjugate and repeating the argument for the pair $h_{1},h_{2}$, and taking $supp(h)\subset{\cal O}_{1}$, we end up with
$\langle\Psi|[:\thinspace\hat{\phi}^{2}\thinspace:[h],:\thinspace\hat{\phi}^{2}\thinspace:[h^{\prime}]]\Psi^{\prime}\rangle=0$
which is the thesis for $\thinspace\hat{\phi}^{2}\thinspace:$ because $\Psi$ ranges over a dense set. The same procedure applies to the proof of commutativity of $\thinspace\hat{\phi}^{2}\thinspace:$ and $\thinspace\hat{T}_{\mu\nu}\thinspace:$ and to the case of two operators of type $\thinspace\hat{T}_{\mu\nu}\thinspace:$, just by applying the operators $D_{\mu\nu}(x,y)$ and $D_{\alpha\beta}(x^{\prime},y^{\prime})$ at the obvious steps of the proof and taking (46) into account. The cases involving operators $\hat{\phi}$ are simplified versions of the above proof. At this juncture the main proof follows straightforwardly. $\Box$
Proof of Proposition 4.11. The proofs of (a) and (c) are straightforward, by applying the relevant definitions and taking advantage of (39) and (e) in Proposition 4.2. The second bound in (b) will be discussed below. The first bound in (b) easily arises by integrating by parts the explicit expression of ($x$-derivatives of) $\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]|\Psi^{\prime}\rangle$, which reads
$\int_{{\mathbb{R}}^{6}}\thinspace\thinspace e^{i(\vec{p}+\vec{k})\cdot\vec{x}}e^{-i(E(p)+E(k))x^{0}}\hat{f}(-p-k)\frac{\langle\Psi|a_{p}a_{k}\Psi^{\prime}\rangle}{4\pi}t_{\mu\nu}(p,-k)\frac{d^{3}pd^{3}k}{E(p)E(k)}$
$\int_{{\mathbb{R}}^{6}}\thinspace\thinspace e^{-i(\vec{p}+\vec{k})\cdot\vec{x}}e^{i(E(p)+E(k))x^{0}}\hat{f}(p+k)\frac{\langle\Psi|a^{\dagger}_{p}a^{\dagger}_{k}\Psi^{\prime}\rangle}{4\pi}t_{\mu\nu}(p,-k)\frac{d^{3}pd^{3}k}{E(p)E(k)}$
$\thinspace\int_{{\mathbb{R}}^{6}}\thinspace\thinspace e^{i(\vec{k}-\vec{p})\cdot\vec{x}}e^{i(E(p)-E(k))x^{0}}\hat{f}(p-k)\frac{\langle\Psi|a^{\dagger}_{p}a_{k}\Psi^{\prime}\rangle}{4\pi}t_{\mu\nu}(p,k)\frac{d^{3}pd^{3}k}{E(p)E(k)}$
$+\int_{{\mathbb{R}}^{6}}\thinspace\thinspace e^{i(\vec{p}-\vec{k})\cdot\vec{x}}e^{i(E(k)-E(p))x^{0}}\hat{f}(k-p)\frac{\langle\Psi|a^{\dagger}_{k}a_{p}\Psi^{\prime}\rangle}{4\pi}t_{\mu\nu}(p,k)\frac{d^{3}pd^{3}k}{E(p)E(k)}\>,$ (118)
where $p\equiv(E(p),\vec{p})$, $k\equiv(E(k),\vec{k})$ with $E(p)=p^{0}=\sqrt{\vec{p}^{2}+m^{2}}$ and $E(k)=k^{0}=\sqrt{\vec{k}^{2}+m^{2}}$. The integrated functions of $\vec{p}$ and $\vec{k}$ are of Schwartz type in ${\mathbb{R}}^{3}\times{\mathbb{R}}^{3}\equiv\mathsf{V}_{m,+}^{2}$ by construction. Thus, in particular, one can pass the $x$-derivatives under the symbol of integration. Let us discuss the proof of (d). First of all we observe that the integrals on both sides are finite since the integrated functions are Schwartz, in view of the bound (55), which is valid in two Minkowskian coordinate systems comoving with the reference frames $u_{\Sigma}$ and $u_{\Sigma^{\prime}}$ respectively. Define $J^{\nu}:=v_{\mu}\langle\Psi|:\thinspace\hat{T}^{\mu\nu}\thinspace:[f_{x}]\Psi^{\prime}\rangle$. We observe that $J$ is smooth and conserved, $\partial_{\mu}J^{\mu}(x)=0$, by (c). Now consider the case where $\Sigma$ and $\Sigma^{\prime}$ are parallel. In this case we can use a common system of Minkowskian coordinates, where $\Sigma$ is the plane $x^{0}=0$ and $\Sigma^{\prime}$ is the analogous plane at $x^{0}=T>0$ (or vice versa). If $C_{r}\subset{\mathbb{R}}^{4}$ is the cylinder with axis $\vec{x}=0$ and basis $B_{r}:=\{\vec{x}\in{\mathbb{R}}^{3}\>|\>|\vec{x}|<r\}$ in $\Sigma$, we can use the divergence theorem for the compact cylinder $G_{r}\subset C_{r}$ bounded by $B_{r}$ and the analogous basis $B^{\prime}_{r}$ in $\Sigma^{\prime}$. We move on to compute the flux of $J$ across $\partial G_{r}$. This flux must be 0 since $J$ is conserved, taking the divergence theorem into account. The flux of $J$ on the lateral surface of $G_{r}$ is bounded by $cTr^{2}$, so that it tends to 0 as $r\to+\infty$ due to (55). The outward fluxes on the bases respectively tend to $\int_{\Sigma}v^{\mu}\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]\Psi^{\prime}\rangle n^{\nu}_{\Sigma}d\Sigma$ and $\int_{\Sigma^{\prime}}v^{\mu}\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]\Psi^{\prime}\rangle n^{\nu}_{\Sigma^{\prime}}d\mu_{\Sigma^{\prime}}$ in view of the dominated convergence theorem. This proves the thesis for the case under consideration. It remains to consider the case where $\Sigma$ and $\Sigma^{\prime}$ are not parallel and thus meet in a $2$-plane. Let us arrange a Minkowskian reference frame with the origin contained in this intersection and $\Sigma$ described by the plane $x^{0}=0$ in ${\mathbb{R}}^{4}$. With a further spatial rotation of the coordinates if necessary, we can describe the $3$-plane $\Sigma^{\prime}$ as the plane $x^{0}=\gamma x^{1}$ in ${\mathbb{R}}^{4}$ where $\gamma\in(0,1)$, since $\Sigma^{\prime}$ is spacelike. We assume that we deal with vectors $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$ such that their (finite smooth) components $\Psi^{n},\Psi^{\prime n}$ have compact support in the corresponding spaces ${\mathbb{R}}^{3n}\equiv\mathsf{V}_{m,+}^{n}$. At this juncture, using the same proof of (b) in (19) of Theorem [CDM26] for the integrals (118), we have that the second bound in (b) is satisfied:
$|\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]|\Psi^{\prime}\rangle|\leq\frac{C}{(1+|\vec{x}|)^{3+\epsilon}}\quad\mbox{if $|\vec{x}|>|x^{0}|$, for some constants $\epsilon>0$, $C<+\infty$.}$
We can now apply the divergence theorem to the cylindrical compact solid with a basis $B_{r}$ on $\Sigma$ as before, lateral surface normal to $\Sigma$, and a basis on $\Sigma^{\prime}$ made of two parts (one above $\Sigma$ and the other below it). The area of the lateral surface is bounded by $cr^{3}$ for some constant $c>0$. Due to the found estimate, since this lateral surface stays in the region $|\vec{x}|>|x^{0}|$ because $0<\gamma<1$, the flux of $J$ across the lateral surface is bounded by $cr^{3}\frac{C}{(1+r)^{3+\epsilon}}$ which vanishes as $r\to+\infty$. The remaining part of the flux, in the limit as $r\to+\infty$, yields (57) for the compactly supported vectors $\Psi,\Psi^{\prime}\in{\mathfrak{D}}_{0}$. To conclude the proof, we have to remove this compactness requirement. We observe that the space of smooth compactly supported functions $\mathscr{D}({\mathbb{R}}^{3n})$ is dense in the Schwartz topology of $\mathscr{S}({\mathbb{R}}^{3n})\equiv\mathscr{S}(\mathsf{V}_{m,+}^{n})$. Therefore, $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$ admit sequences ${\mathfrak{D}}_{0}\ni\Psi^{n}\to\Psi$, ${\mathfrak{D}}_{0}\ni\Psi^{\prime n}\to\Psi^{\prime}$ (as $n\to+\infty$) such that the convergence, component by component, is in the relevant Schwartz topology. As a consequence
$e^{i(E(p)+E(k))x^{0}}\hat{f}(p+k)\frac{\langle\Psi_{n}|a^{\dagger}_{p}a^{\dagger}_{k}\Psi^{\prime}_{n}\rangle}{2(2\pi)^{3}E(p)E(k)}t_{\mu\nu}(p,-k)$
$\to e^{i(E(p)+E(k))x^{0}}\hat{f}(p+k)\frac{\langle\Psi|a^{\dagger}_{p}a^{\dagger}_{k}\Psi^{\prime}\rangle}{2(2\pi)^{3}E(p)E(k)}t_{\mu\nu}(p,-k)\quad\mbox{as $n\to+\infty$}$
in the Schwartz topology of functions of $(\vec{p},\vec{k})\in{\mathbb{R}}^{6}$, and the same result is valid for the other similar terms in (118). Using (118), and the fact that the Fourier transform (of functions of the variables $\vec{u}:=\vec{k}\pm\vec{p}$ and $\vec{v}:=\vec{k}\mp\vec{p}$) is continuous with respect to the Schwartz topology, we have that, as $n\to+\infty$,
$\int_{\Sigma}n^{\mu}\langle\Psi^{n}|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]\Psi^{\prime n}\rangle n^{\nu}_{\Sigma}d\Sigma(x)\to\int_{\Sigma}n^{\mu}\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]\Psi^{\prime}\rangle n^{\nu}_{\Sigma}d\Sigma(x)\>,$
and the same is valid for $\Sigma^{\prime}$ when referring to a Minkowskian coordinate system adapted to it. This fact concludes the proof of (57) for the general case $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$. Concerning (e), the proof immediately arises from (58), since $u_{\mu}u^{\prime}_{\nu}\langle\Psi|:\thinspace\hat{T}^{\mu\nu}\thinspace:[f_{x}^{2}]\Psi\rangle=$
$u_{\mu}u^{\prime}_{\nu}\langle U_{(I,x)}^{-1}\Psi|:\thinspace\hat{T}^{\mu\nu}\thinspace:[f^{2}]U_{(I,x)}^{-1}\Psi\rangle\geq-u_{\mu}u^{\prime}_{\nu}b_{f}^{\mu\nu}||U_{(I,x)}^{-1}\Psi||^{2}=-u_{\mu}u^{\prime}_{\nu}b_{f}^{\mu\nu}||\Psi||^{2}\>.$
The proof is over. $\Box$
Proof of Proposition 5.1. First of all, we observe that $\frac{1}{\sqrt{H^{u}_{\epsilon}}}:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]\frac{1}{\sqrt{H^{u}_{\epsilon}}}\Psi$ is well defined if $\Psi\in{\mathfrak{S}}_{0}$ because the bounded operators $\frac{1}{\sqrt{H^{u}+\epsilon I}}$ are everywhere defined and leave ${\mathfrak{S}}_{0}$ invariant. Further comments on the proof are in order. Existence of ${\mathsf{A}}_{f,\epsilon}^{u}(\Sigma)$ and (b) are trivial consequences of Proposition 4.12, just referring to state vectors of the form $\frac{1}{\sqrt{H^{u}+\epsilon I}}\Psi$ instead of $\Psi\in{\mathfrak{S}}_{0}$, remembering that $\frac{1}{\sqrt{H^{u}+\epsilon I}}$ leaves ${\mathfrak{S}}_{0}$ invariant. Given that, if $\Delta\in\mathscr{B}_{0}(\Sigma)$ and $|\Delta|:=+\infty$, we define ${\mathsf{A}}_{f,\epsilon}^{u}(\Delta):={\mathsf{A}}_{f,\epsilon}(\Sigma)-{\mathsf{A}}_{f,\epsilon}^{u}(\Sigma\setminus\Delta)$ provided ${\mathsf{A}}_{f,\epsilon}^{u}(E)$ exists when $|E|<+\infty$, as we shall prove below. Notice that with this definition (63) is satisfied in all cases. In summary, it is sufficient to prove that ${\mathsf{A}}_{f,\epsilon}^{u}(\Delta)$ does exist when $\Delta$ has finite measure. Regarding (d), additivity next follows from (63) directly, and weak $\sigma$-additivity then easily follows from the $\sigma$-additivity of the right-hand side of (63). This, in turn, is an obvious consequence of the fact that $\Sigma\ni\vec{x}\mapsto|\langle\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]|\Psi^{\prime}\rangle|$ is integrable in view of (b)(1) of Proposition 4.11. Concerning (a), we observe that ${\mathsf{A}}_{f,\epsilon}^{u}(\Delta)$, if any, is selfadjoint because it is bounded and Hermitian on a dense domain (${\mathfrak{S}}_{0}$). Hermiticity on ${\mathfrak{S}}_{0}$ follows from (63) since the quadratic form in the integral is real for $\Psi=\Psi^{\prime}$, so that the integral and $\langle\Psi|{\mathsf{A}}_{f,\epsilon}^{u}(\Delta)\Psi\rangle$ itself are real as well. An elementary polarization argument implies that ${\mathsf{A}}_{f,\epsilon}^{u}(\Delta)$ is Hermitian thereon. We shall prove (a1) later. (c) is a consequence of (d) of Proposition 4.2, (63), and continuity arguments if $|\Delta|<+\infty$, and it directly follows from (64), additivity, the definition of $H^{u}$ in the remaining cases, and elementary spectral calculus.
To go on with the rest of the proof of existence for the case $|\Delta|<+\infty$ we shall make use of a couple of lemmata whose proofs appear below in this appendix.
Lemma B.2:
[…]
Proof.
See below. ∎
Lemma B.3:
[…]
Proof.
See below. ∎
Let us move on to prove the main thesis. Referring to a Minkowskian coordinate system $x^{0},x^{1},x^{2},x^{3}$ where $\Sigma$ corresponds to $x^{0}=0$, we decompose the quadratic form
$\int_{\Delta}\thinspace u^{\mu}\left\langle\frac{1}{\sqrt{H^{u}+\epsilon I}}\Psi\left|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f^{2}_{x}]\right.\frac{1}{\sqrt{H^{u}+\epsilon I}}\Psi^{\prime}\right\rangle u^{\nu}_{\Sigma}\>\>d\Sigma(x)=\sum_{i=1}^{3}T^{\Delta}_{\epsilon,i}(\Psi,\Psi^{\prime})$ (119)
according to the decomposition (118), where we omitted the indices $f,u$ which are supposed to be given. The quadratic forms $T^{\Delta}_{\epsilon,i}(\Psi,\Psi^{\prime})$ are well defined for the generic case $\Delta\in\mathscr{B}(\Sigma)$, not only for $\Delta\in\mathscr{B}_{0}(\Sigma)$. Furthermore, they are also well defined if $\epsilon=0$ provided $\Psi,\Psi^{\prime}\not\in{{\cal H}}^{(0)}$, and interpreting $\frac{1}{\sqrt{H^{u}}}$ in the spectral sense in ${{\cal H}}^{(0)\perp}$ and as the zero operator in ${{\cal H}}^{(0)}$. At least in the case $\Delta\in\mathscr{B}_{0}(\Sigma)$ and $\epsilon>0$, we expect to find a decomposition corresponding to (119) of the operator ${\mathsf{A}}_{f,\epsilon}^{u}(\Delta)$ we are looking for, where, again, we omit the indices $f,u$ which are supposed to be given,
${\mathsf{A}}_{f,\epsilon}^{u}(\Delta)=\sum_{i=1}^{3}{\mathsf{A}}_{\epsilon,i}(\Delta)\>.$ (120)
We start by focusing on the case $i=3$ where, as we shall see, the operator ${\mathsf{A}}_{\epsilon,3}(\Delta)$ turns out to be defined also for $\epsilon=0$ and $\Delta\in\mathscr{B}(\Sigma)$, possibly unbounded with unbounded complement. Now $T^{\Delta}_{\epsilon,3}(\Psi,\Psi^{\prime}):=T^{\Delta}_{\epsilon,3}(\Psi,\Psi^{\prime})_{+}+T^{\Delta}_{\epsilon,3}(\Psi,\Psi^{\prime})_{-}$ where
$T^{\Delta}_{\epsilon,3}(\Psi,\Psi^{\prime})_{+}:=\int_{\Delta}\int_{{\mathbb{R}}^{6}}\thinspace e^{i(\vec{k}-\vec{p})\cdot\vec{x}}\widehat{f^{2}}(p-k)\frac{\langle W^{\epsilon}_{u}\Psi|a^{\dagger}_{p}a_{k}W^{\epsilon}_{u}\Psi^{\prime}\rangle}{4\pi}u^{\mu}t_{\mu 0}(p,k)\frac{d^{3}pd^{3}k}{E(p)E(k)}d^{3}x\>,$
$T^{\Delta}_{\epsilon,3}(\Psi,\Psi^{\prime})_{-}:=\int_{\Delta}\int_{{\mathbb{R}}^{6}}\thinspace\thinspace e^{i(\vec{p}-\vec{k})\cdot\vec{x}}\hat{f^{2}}(-p+k)\frac{\langle W^{\epsilon}_{u}\Psi|a^{\dagger}_{k}a_{p}W^{\epsilon}_{u}\Psi^{\prime}\rangle}{4\pi}u^{\mu}t_{\mu 0}(p,k)\frac{d^{3}pd^{3}k}{E(p)E(k)}d^{4}x\>,$
and where $W^{\epsilon}_{u}:=(H^{u}+\epsilon I)^{-1/2}\in{\mathfrak{B}}({\mathfrak{F}}_{s}({{\cal H}}_{m}))$. Evidently $T^{\Delta}_{\epsilon,3}(\Psi,\Psi^{\prime})_{+}=T^{\Delta}_{\epsilon,3}(\Psi,\Psi^{\prime})_{-}$. The quadratic form vanishes unless both vectors belong to the same subspace $\Psi,\Psi^{\prime}\in{{\cal H}}^{(n)}\cap{\mathfrak{S}}_{0}$. We henceforth restrict ourselves to the study of the quadratic form on $({{\cal H}}^{(n)}\cap{\mathfrak{S}}_{0})\times({{\cal H}}^{(n)}\cap{\mathfrak{S}}_{0})$ and we want to prove that it is induced by an operator ${\mathsf{A}}_{\epsilon,3,n}(\Delta)\in{\mathfrak{B}}({{\cal H}}^{(n)})$, taking advantage of Lemma B.2. If $n=0$ there is nothing to prove because the quadratic form vanishes: ${\mathsf{A}}_{\epsilon,3,0}(\Delta)=0$. We therefore assume $n>0$ so that $\Psi=\Psi(p,Q)$ and $\Psi^{\prime}=\Psi^{\prime}(k,Q)$ are functions in $\mathscr{S}(\mathsf{V}_{m,+}^{n})$ and $Q:=(q^{2},\ldots,q^{n})$. $T^{\Delta}_{3}(\Psi,\Psi^{\prime})$ can be rearranged to
$\frac{n}{2\pi}\int_{\Delta}\int_{{\mathbb{R}}^{6}}\int_{{\mathbb{R}}^{3(n-1)}}\thinspace e^{i(\vec{k}-\vec{p})\cdot\vec{x}}\widehat{f^{2}}(p-k)\frac{\overline{\Psi(p,Q)}\Psi^{\prime}(k,Q)}{\sqrt{{E^{n}_{\epsilon,u}(p,Q)}E^{n}_{\epsilon,u}(k,Q)}}u^{\mu}t_{\mu 0}(p,k)\frac{d^{3}pd^{3}kd^{(n-1)}q}{E(p)E(k)\prod_{i}E(q_{i})}d^{3}x$
where, for $\epsilon\geq 0$,
$E^{n}_{\epsilon,u}(p,Q):=-u\cdot\left(p+\sum_{r=2}^{n}q_{r}\right)+\epsilon\>.$
Using the Hilbert space isomorphism
$J_{n}:L^{2}(\mathsf{V}^{n}_{m,+},d\mu^{r})\ni\Psi(p_{1},\ldots p_{n})\mapsto\Phi(\vec{p}_{1},\ldots\vec{p}_{n})$
$:=\frac{\Psi(p_{1},\ldots p_{n})}{\sqrt{E(p_{1})\cdots E(p_{r})}}\in L^{2}({\mathbb{R}}^{3n},d^{3n}p)$ (121)
and swapping two integrations of Schwartz functions, the written integral can be rephrased as (where $\vec{Q}:=(\vec{q}_{2},\ldots,\vec{q}_{n})$) $T^{\Delta}_{\epsilon,3}(\Psi,\Psi^{\prime})=S_{\epsilon,3}(\Phi,\Phi^{\prime})$
$:=\frac{n}{2\pi}\int_{\Delta}\int_{{\mathbb{R}}^{3(n-1)}}\thinspace\int_{{\mathbb{R}}^{6}}e^{i(\vec{k}-\vec{p})\cdot\vec{x}}\widehat{f^{2}}(p-k)\frac{\overline{\Phi(\vec{p},\vec{Q})}\Phi^{\prime}(\vec{k},\vec{Q})}{\sqrt{{E^{n}_{\epsilon,u}(p,Q)}E^{n}_{\epsilon,u}(k,Q)}}u^{\mu}t_{\mu 0}(p,k)\frac{d^{3}pd^{3}kd^{(n-1)}q}{\sqrt{E(p)E(k)}}d^{3}x\>.$ (122)
Finally, the convolution theorem and the fact that $f$ is real imply that
$\widehat{f^{2}}(p-k)=\int_{{\mathbb{R}}^{4}}\hat{f}(p-k-u)\hat{f}(u)d^{4}u=\int_{{\mathbb{R}}^{4}}\hat{f}(-u-k)\hat{f}(u+p)d^{4}u=\int_{{\mathbb{R}}^{4}}\overline{\hat{f}(u+k)}\hat{f}(u+p)d^{4}u\>.$
According to this result, if $\Psi,\Psi^{\prime}\in{{\cal H}}^{(n)}\cap{\mathfrak{S}}_{0}$ and $\Phi,\Phi^{\prime}$ are respectively associated by means of the unitary map (121), then $T^{\Delta}_{\epsilon,3}(\Psi,\Psi^{\prime})=S_{\epsilon,3}(\Phi,\Phi^{\prime})$ can be rearranged to
$T^{\Delta}_{\epsilon,3}(\Psi,\Psi^{\prime})=(2\pi)^{2}n\int_{\Delta}\thinspace\int_{{\mathbb{R}}^{3(n-1)}}\int_{{\mathbb{R}}^{6}}\int_{{\mathbb{R}}^{4}}\thinspace\frac{\overline{\hat{f}(u+p)\Phi(\vec{p},\vec{Q})}e^{i(\vec{k}-\vec{p})\cdot\vec{x}}\hat{f}(u+k)\Phi^{\prime}(\vec{k},\vec{Q})}{(2\pi)^{3}\sqrt{{E^{n}_{\epsilon,u}(p,Q)}E^{n}_{\epsilon,u}(k,Q)}}$
$\times\frac{u^{\mu}t_{\mu 0}(p,k)}{\sqrt{E(p)E(k)}}d^{4}ud^{3}pd^{3}kd^{(n-1)}qd^{3}x\>.$ (123)
We now consider the reduced quadratic form
$S^{\prime}_{3}(\Psi,\Psi^{\prime}):=(2\pi)^{2}n\int_{\Delta}\thinspace\int_{{\mathbb{R}}^{3(n-1)}}\int_{{\mathbb{R}}^{6}}\int_{{\mathbb{R}}^{4}}\thinspace\frac{\overline{\hat{f}(u+p)\Phi(\vec{p},\vec{Q})}e^{i(\vec{k}-\vec{p})\cdot\vec{x}}\hat{f}(u+k)\Phi^{\prime}(\vec{k},\vec{Q})}{(2\pi)^{3}}d^{4}ud^{3}pd^{3}kd^{(n-1)}qd^{3}x\>.$
We observe that the internal integrals in $d^{4}u$ and $d^{3}pd^{3}k$ can be swapped by Fubini’s theorem because
$\int_{{\mathbb{R}}^{6}}\int_{{\mathbb{R}}^{4}}\thinspace|\hat{f}(u+p)\Phi(\vec{p},\vec{Q})\hat{f}(u+k)\Phi^{\prime}(\vec{k},\vec{Q})|d^{4}ud^{3}pd^{3}k$
$\int_{{\mathbb{R}}^{6}}|\Phi(\vec{p},\vec{Q})\Phi^{\prime}(\vec{k},\vec{Q})|\int_{{\mathbb{R}}^{4}}\thinspace|\hat{f}(u+p-k)\hat{f}(u)|d^{4}ud^{3}pd^{3}k$
$\leq(\sup|\hat{f}|)\int_{{\mathbb{R}}^{6}}|\Phi(\vec{p},\vec{Q})\Phi^{\prime}(\vec{k},\vec{Q})|d^{3}pd^{3}k\int_{{\mathbb{R}}^{4}}\thinspace|\hat{f}(u)|d^{4}u<+\infty\>.$
The final formula for $S^{\prime}_{3}(\Psi,\Psi^{\prime})$ is
$(2\pi)^{2}n\int_{\Delta}\thinspace\int_{{\mathbb{R}}^{3(n-1)}}\int_{{\mathbb{R}}^{4}}\int_{{\mathbb{R}}^{6}}\thinspace\frac{\overline{\hat{f}(u+p)\Phi(\vec{p},\vec{Q})}e^{i(\vec{k}-\vec{p})\cdot\vec{x}}\hat{f}(u+k)\Phi^{\prime}(\vec{k},\vec{Q})}{(2\pi)^{3}}d^{3}pd^{3}kd^{4}ud^{(n-1)}qd^{3}x\>.$
We can interpret this formula in the Hilbert space $L^{2}({\mathbb{R}}^{3}\times{\mathbb{R}}^{4}\times{\mathbb{R}}^{3n},d^{3}pd^{4}ud^{(n-1)}q)$ as follows. If defining the function $F_{\Phi}:=F_{\Phi}(\vec{p},u,\vec{Q}):=\hat{f}(u^{0}+E(p),\vec{u}+\vec{p})\Phi(\vec{p},\vec{Q})$, we have that, for $s>0$
$\int|F_{\Phi}|^{s}d^{4}ud^{3}pd^{(n-1)}q=\int|\hat{f}(u^{0}+E(p),\vec{u}+\vec{p})|^{s}|\Phi(\vec{p},\vec{Q})|^{s}d^{3}pd^{4}ud^{(n-1)}q$
$\int\left(\int|\hat{f}(u^{0}+E(p),\vec{u}+\vec{p})|^{s}d^{4}u\right)|\Phi(\vec{p},\vec{Q})|^{s}d^{3}pd^{(n-1)}q$
$=\left(\int|\hat{f}(u^{0},\vec{u})|^{s}d^{4}u\right)\int|\Phi(\vec{p},\vec{Q})|^{s}d^{3}pd^{(n-1)}q<+\infty\>.$ (124)
[…]
$\mathscr{B}({\mathbb{R}}^{3})\ni\Delta\mapsto P(\Delta)\otimes I\otimes I\in{\mathfrak{B}}(L^{2}({\mathbb{R}}^{3}\times{\mathbb{R}}^{4}\times{\mathbb{R}}^{3n},d^{3}pd^{4}ud^{(n-1)}q))$
where $P(\Delta)$ is the joint PVM of the three standard position operators in $L^{2}({\mathbb{R}}^{3},d^{3}x)$ in momentum representation (see, e.g., [Mor19]). Notice that this reasoning is valid even if $|\Delta|=+\infty$. Using the fact that the Fourier-Plancherel unitary operator in $L^{2}$ is the standard Fourier transform in $L^{1}$ and ${\mathbb{R}}^{3}\ni\vec{p}\mapsto F_{\Phi}(\vec{p},u,\vec{Q})$ is Schwartz for every choice of $(u,\vec{Q})$,
$S^{\prime}_{3}(\Phi,\Phi^{\prime})=(2\pi)^{2}n\int_{\Delta}d^{3}x\thinspace\int_{{\mathbb{R}}^{3(n-1)}}\thinspace\thinspace\thinspace\thinspace d^{(n-1)}q\int_{{\mathbb{R}}^{4}}d^{4}u\overline{\int_{{\mathbb{R}}^{3}}\thinspace\frac{e^{i\vec{p}\cdot\vec{x}}}{(2\pi)^{3/2}}F_{\Phi}(\vec{p},u,\vec{Q})d^{3}p}\int_{{\mathbb{R}}^{3}}\thinspace\frac{e^{i\vec{k}\cdot\vec{x}}}{(2\pi)^{3/2}}F_{\Phi^{\prime}}(\vec{k},u,\vec{Q})d^{3}k\>.$
$(2\pi)^{2}n\langle F_{\Phi}|(P(\Delta)\otimes I\otimes I)F_{\Phi^{\prime}}\rangle\>.$
As a consequence,
[…]
where we also used the Plancherel theorem. To conclude, we observe that if $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}\cap{{\cal H}}^{(n)}$,
$T^{\Delta}_{\epsilon,3}(\Psi,\Psi^{\prime})=S_{\epsilon,3}(\Phi,\Phi^{\prime})=\sum_{a=0}^{N}S^{\prime}_{3}(M_{\epsilon,a}\Psi,M^{\prime}_{\epsilon,a}\Psi^{\prime})$
where, for some (of first order at most and $r$ independent) polynomials $P_{a}(p)$ in the components of $p\equiv(E(p),\vec{p})$, constructed by decomposing $u^{\mu}t_{\mu\nu}(p,k)u_{\Sigma}^{\nu}$,
$(M_{\epsilon,a}\Phi)(\vec{p},\vec{Q})=\frac{P_{a}(p)}{\sqrt{E^{n}_{\epsilon,u}(p,Q)E(p)}}\Phi(\vec{p},\vec{Q})\>.$
It holds
$M_{\epsilon,a}||=\sup_{(\vec{p},\vec{Q})\in{\mathbb{R}}^{3n}}\left|\frac{P_{a}(p)}{\sqrt{E^{n}_{\epsilon,u}(p,Q)E(p)}}\right|\leq n^{-1/2}K_{a}\>,$
for some constant $K_{a}\geq 0$ independent of $\epsilon$ and $n$ (but depending on $u,u_{\Sigma}$). Indeed, change coordinates in order that the temporal axis coincides with $u$. Observe that the components of $p$ are however bounded by $p^{0}$ (referred to the new basis), and $(u_{\Sigma}^{0}-\frac{\vec{p}\cdot\vec{u}_{\Sigma}}{p^{0}})^{2}$ is bounded below when $p$ varies in $\mathsf{V}_{m,+}$ and $u_{\Sigma}\in{\mathsf{T}}_{+}$ is given [^16] In fact, $(u_{\Sigma}^{0}-\frac{\vec{p}\cdot\vec{u}_{\Sigma}}{p^{0}})^{2}=\left(\cosh\chi-\frac{\vec{p}\cdot\vec{e}}{\sqrt{\vec{p}^{2}+m^{2}}}\sinh\chi\right)^{2}$ where $u_{\Sigma}\equiv(\cosh\chi,\sinh\chi\vec{e})$ for $\chi\geq 0$ given and some unit $\vec{e}$ $g$-normal to $u\equiv(1,\vec{0})$ and where $\vec{p}\in{\mathbb{R}}^{3}$. Varying $\vec{p}$, the function is bounded below by $\cosh\chi-\sinh\chi>0$.. Putting together these facts we have
$\left|\frac{P_{a}(p)}{\sqrt{E^{n}_{\epsilon,u}(p,Q)E(p)}}\right|^{4}\leq\frac{(p^{0})^{4}C_{a}}{(p^{0}+q^{0}_{2}+\cdots+q^{0}_{n})^{2}}\frac{1}{(u_{\Sigma}^{0}p^{0}-\vec{p}\cdot\vec{u}_{\Sigma})^{2}}$
$\leq\frac{(p^{0})^{2}C_{a}}{(p^{0}+q^{0}_{2}+\cdots+q^{0}_{n})^{2}}\frac{1}{(u_{\Sigma}^{0}-\frac{\vec{p}\cdot\vec{u}_{\Sigma}}{p^{0}})^{2}}\leq\frac{(p^{0})^{2}C_{a}}{(p^{0}+q^{0}_{2}+\cdots+q^{0}_{n})^{2}}\>.$
The latter function is defined for $p^{0},q^{0}_{2},\ldots,q_{n}^{0}\geq m$ and reaches its maximum value $C_{a}n^{-2}$ on this boundary as follows by direct inspection. Putting all together, if $\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}\cap{{\cal H}}^{(n)}$, the $\epsilon,n$-uniform bound holds
[…]
We end up with the estimate, for some $C<+\infty$ independent of $\epsilon\geq 0$, $n\in{\mathbb{N}}$, $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{M}})$, and $\Delta$ (but depending on $u,u_{\Sigma}$),
[…]
In view of Lemma B.2 we have that, if $n=0,1,\ldots$, there is a unique ${\mathsf{A}}_{\epsilon,3,n}(\Delta)\in{\mathfrak{B}}({{\cal H}}^{(n)})$ such that $\langle\Psi|{\mathsf{A}}_{\epsilon,3,n}(\Delta)\Psi^{\prime}\rangle=T^{\Delta}_{\epsilon,3}(\Psi,\Psi^{\prime})$ for $\Psi,\Psi^{\prime}\in{{\cal H}}^{(n)}\cap{\mathfrak{S}}_{0}$. ${\mathsf{A}}_{\epsilon,3,0}$ is the zero operator. These maps are defined for $\epsilon\geq 0$ and also if $|\Delta|=+\infty$. We are now allowed to view the found operators as bounded linear maps ${\mathsf{A}}_{\epsilon,3,n}(\Delta):{{\cal H}}^{(n)}\to{\mathfrak{F}}_{s}({{\cal H}}_{m})$, so that ${{\cal H}}^{(n)}$ is an invariant space, and this is consistent with the identity $\langle\Psi|{\mathsf{A}}_{\epsilon,3,n}(\Delta)\Psi^{\prime}\rangle=T^{\Delta}_{\epsilon,3}(\Psi,\Psi^{\prime})$ for $\Psi^{\prime}\in{{\cal H}}^{(n)}\cap{\mathfrak{S}}_{0}$ and $\Psi\in{\mathfrak{S}}_{0}$, because the quadratic form only sees the $n$-th component. Since the norms of the operators ${\mathsf{A}}_{\epsilon,3,n}(\Delta)$ are $\epsilon,n$-uniformly bounded, the spaces ${{\cal H}}^{(n)}$ orthogonally decompose ${\mathfrak{F}}_{s}({{\cal H}}_{m})$, and the images of operators ${\mathsf{A}}_{\epsilon,3,n}(\Delta)$ are mutually orthogonal, Lemma B.3 entails that, for every $\epsilon\geq 0$ and $\Delta\in\mathscr{B}(\Sigma)$, there is a unique ${\mathsf{A}}_{\epsilon,3}(\Delta)\in{\mathfrak{B}}({\mathfrak{F}}_{s}({{\cal H}}_{m}))$ which extends these operators. By construction
$\langle\Psi|{\mathsf{A}}_{\epsilon,3}(\Delta)\Psi^{\prime}\rangle=T^{\Delta}_{\epsilon,3}(\Psi,\Psi^{\prime})\quad\mbox{$\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$},$ (125)
with $T^{\Delta}_{0,3}(\Psi,\Psi^{\prime})$ defined as discussed under (119). In particular, for $\Delta\in\mathscr{B}(\Sigma)$ and $\epsilon\geq 0$,
[…]
Equivalently, there is $C<+\infty$ independent of $f$ such that
${\mathsf{A}}_{\epsilon,3}(\Delta)||\leq C\left(\int_{{\mathbb{M}}}f^{2}d^{4}x\right)\quad\mbox{for every $\epsilon\geq 0$, $\Delta\in\mathscr{B}(\Sigma)$.}$
This inequality proves (a1), when also assuming $\epsilon>0$ and $|\Delta|<+\infty$ to have a well-defined ${\mathsf{A}}^{u}_{f,\epsilon}(\Delta)$, since $P_{n}{\mathsf{A}}^{u}_{f,\epsilon}(\Delta)P_{n}=P_{n}{\mathsf{A}}_{\epsilon,3}(\Delta)P_{n}$. The result trivially extends to the case $|\Sigma\setminus\Delta|<+\infty$.
We consider the further term $T^{\Delta}_{\epsilon,2}$ in the expansion (119), and the corresponding ${\mathsf{A}}_{\epsilon,2}(\Delta)$ in (120) where we explicitly assume that $\Delta$ has finite measure and $\epsilon>0$.
$T^{\Delta}_{\epsilon,2}(\Psi,\Psi^{\prime})=\int_{\Delta}\int_{{\mathbb{R}}^{6}}\thinspace\thinspace e^{i(\vec{p}+\vec{k})\cdot\vec{x}}\widehat{f^{2}}(-p-k)\frac{\langle W^{\epsilon}_{u}\Psi|a_{p}a_{k}W^{\epsilon}_{u}\Psi^{\prime}\rangle}{4\pi}u^{\mu}t_{\mu 0}(p,-k)\frac{d^{3}pd^{3}k}{E(p)E(k)}d^{4}x\>,$
where, as before, $W^{\epsilon}_{u}=(H^{u}+\epsilon I)^{-1/2}\in{\mathfrak{B}}({\mathfrak{F}}_{s}({{\cal H}}_{m}))$. In this case the quadratic form vanishes unless $\Psi^{\prime}\in{{\cal H}}^{(n+2)}\cap{\mathfrak{S}}_{0}$ and $\Psi\in{{\cal H}}^{(n)}\cap{\mathfrak{S}}_{0}$, therefore we will assume this henceforth. With the same procedure as above, taking advantage of the isomorphisms of Hilbert spaces $J_{n}$ defined in (121), $T^{\Delta}_{\epsilon,2}(\Psi,\Psi^{\prime})=S_{\epsilon,2}(\Phi,\Phi^{\prime})$ can be rearranged to
$C_{n}\int_{{\mathbb{R}}^{6}}\int_{{\mathbb{R}}^{3n}}\thinspace\frac{\overline{\Phi(\vec{Q})}}{\sqrt{\overline{E^{n}_{\epsilon,u}(Q)}}}\frac{{\left(\int_{\Delta}e^{i(\vec{p}+\vec{k})\cdot\vec{x}}d^{3}x\right)\widehat{f^{2}}(-p-k)u^{\mu}t_{\mu 0}(p,-k)\Phi^{\prime}(\vec{p},\vec{k},\vec{Q})}}{\sqrt{{E^{n+2}_{\epsilon,u}(p,k,Q)}E(p)E(k)}}d^{4}ud^{3}pd^{3}kd^{n}q\>,$ (127)
where $E^{n}_{\epsilon,u}(Q):=-\sum_{j=1}^{n}u\cdot q_{j}+\epsilon$ and $C_{n}:=(4\pi)^{-1}\sqrt{(n+1)(n+2)}$ and we have also interchanged the integral in $x$ with the others since $|\Delta|<+\infty$ and in the remaining variables the functions are of Schwartz type, including $(\vec{p},\vec{k})\mapsto\hat{f^{2}}(-p-k)$ as already observed in the proof of Proposition 4.2. Defining the function $H_{\Delta}(\vec{p},\vec{k}):=\left(\int_{\Delta}e^{-i(\vec{p}+\vec{k})\cdot\vec{x}}d^{3}x\right)\hat{f^{2}}(-p-k)u^{\mu}t_{\mu 0}(p,-k)$, this (symmetric) element of $L^{2}({\mathbb{R}}^{6},d^{3}pd^{3}k)$ induces a bounded linear map
$L_{n+2,H_{\Delta}}:J_{n+2}({{\cal H}}^{(n+2)})\ni\Phi\mapsto\Phi_{H_{\Delta}}\in J_{n}({{\cal H}}^{(n)})\>,\quad\Phi_{H_{\Delta}}:=\int_{{\mathbb{R}}^{6}}d^{3}pd^{3}qH_{\Delta}(\vec{p},\vec{k})\Phi(\vec{p},\vec{k},\vec{Q})\>.$
The said operator transforms Schwartz functions to Schwartz functions and can be written as $L_{n+2,H_{\Delta}}=\langle H_{\Delta}|\otimes I:L^{2}({\mathbb{R}}^{6},d^{3}pd^{3}k)\otimes L^{2}({\mathbb{R}}^{3n},d^{3n}q)\to L^{2}({\mathbb{R}}^{3n},d^{3n}q)$, so that in view of Riesz’ lemma
[…]
Looking at (127) we can already conclude that, if the Hilbert space isomorphisms $J_{n}$ are defined in (121), a bounded operator which implements the considered quadratic form is ${\mathsf{A}}_{\epsilon,2,n+2}(\Delta):{{\cal H}}^{(n+2)}\to{{\cal H}}^{(n)}$ such that, if $n\geq 0$ and $\epsilon>0$,
${\mathsf{A}}_{\epsilon,2,n+2}(\Delta)=\frac{1}{4\pi}\sqrt{(n+1)(n+2)}J_{n}^{-1}S_{\epsilon,n}L_{n+2,H_{\Delta}}D_{\epsilon,n+2}J_{n+2}$
where $S_{\epsilon,n}$ is the multiplicative operator defined by the function $S_{\epsilon,n}:\vec{Q}\mapsto\frac{1}{\sqrt{E^{n}_{\epsilon,u}(Q)}}$, bounded by $1/\sqrt{nm+\epsilon}$, and $D_{\epsilon,n+2}$ is the analogous operator defined by the function $D_{\epsilon,n+2}:(\vec{p},\vec{k},\vec{Q})\mapsto\frac{1}{\sqrt{{E^{n+2}_{\epsilon,u}(p,k,Q)}E(p)E(k)}}$ bounded by $1/\sqrt{(n+2)m}$. We therefore have the bound for $\Psi\in{{\cal H}}^{(n+2)}$ and $\Psi^{\prime}\in{{\cal H}}^{(n)}$ with $n\geq 0$ and $\epsilon>0$
[…]
In summary, if $\epsilon>0$ and $n=0,1,\ldots$ we have an $n$-uniform bound
${\mathsf{A}}_{\epsilon,2,n+2}(\Delta)||\leq C_{\epsilon,\Delta}\>.$
(Notice that however $C_{\epsilon,\Delta}\to+\infty$ for $\epsilon\to 0^{+}$.) Let us interpret each operator ${\mathsf{A}}_{\epsilon,2,n}(\Delta)$ as defined on the domain ${{\cal H}}^{(n)}$ to the whole ${\mathfrak{F}}_{s}({{\cal H}}_{m})$, and define ${\mathsf{A}}_{\epsilon,2,0}(\Delta)={\mathsf{A}}_{\epsilon,2,1}(\Delta)=0$. Since $Ran({\mathsf{A}}_{\epsilon,2,r})\perp Ran({\mathsf{A}}_{\epsilon,2,s})$ if $r\neq s$, with the same procedure as for the case of $T^{\Delta}_{\epsilon,3}$ based on the two initially proved lemmata, we can conclude that there exists a unique ${\mathsf{A}}_{\epsilon,2}(\Delta)\in{\mathfrak{B}}({\mathfrak{F}}_{s}({{\cal H}}_{m}))$ which extends the family of operators ${\mathsf{A}}_{\epsilon,2,n}(\Delta):{{\cal H}}^{(n)}\to{{\cal H}}^{(n-2)}\subset{\mathfrak{F}}_{s}({{\cal H}}_{m})$, $n=0,1,2,\ldots$ (where ${{\cal H}}^{(n-2)}=\{0\}$ if $n<2$). By construction, if $\epsilon>0$,
[…]
We conclude with the analysis of the term $T^{\Delta}_{\epsilon,1}$ in the expansion (119), i.e. ${\mathsf{A}}_{\epsilon,1}(\Delta)$ in (120), where, again, we explicitly assume that $\Delta$ has finite measure.
$T^{\Delta}_{\epsilon,1}(\Psi,\Psi^{\prime}):=\int_{\Delta}\int_{{\mathbb{R}}^{6}}\thinspace\thinspace e^{-i(\vec{p}+\vec{k})\cdot\vec{x}}\widehat{f^{2}}(p+k)\frac{\langle W^{\epsilon}_{u}\Psi|a^{\dagger}_{p}a^{\dagger}_{k}W^{\epsilon}_{u}\Psi^{\prime}\rangle}{4\pi}u^{\mu}t_{\mu 0}(p,-k)\frac{d^{3}pd^{3}k}{E(p)E(k)}d^{4}x\>,$
where, as before, $W^{\epsilon}_{u}=(H^{u}+\epsilon I)^{-1/2}\in{\mathfrak{B}}({\mathfrak{F}}_{s}({{\cal H}}_{m}))$. An elementary procedure based on the observation that $\overline{\hat{f^{2}}(q)}=\hat{f^{2}}(-q)$ because $f$ is real, and taking the property $\langle W^{\epsilon}_{u}\Psi|a^{\dagger}_{p}a^{\dagger}_{k}W^{\epsilon}_{u}\Psi^{\prime}\rangle=\overline{\langle W^{\epsilon}_{u}\Psi^{\prime}|a_{p}a_{k}W^{\epsilon}_{u}\Psi\rangle}$ into account, proves that the unique wanted operator ${\mathsf{A}}_{\epsilon,1}(\Delta)\in{\mathfrak{B}}({\mathfrak{F}}_{s}({{\cal H}}_{m}))$ such that
$\langle\Psi|{\mathsf{A}}_{\epsilon,1}(\Delta)\Psi^{\prime}\rangle=T^{\Delta}_{\epsilon,1}(\Psi,\Psi^{\prime})\quad\mbox{$\Psi,\Psi^{\prime}\in{\mathfrak{S}}_{0}$}.$ (129)
is exactly ${\mathsf{A}}_{\epsilon,1}(\Delta)={\mathsf{A}}_{\epsilon,2}(\Delta)^{\dagger}$.
Item (e) can be finally proved as follows. The operators are bounded from below as immediate consequence of their definition, the uniform bound from below established in (e) Proposition 4.11 and, obviously, the fact that $(H^{u}_{\epsilon})^{-1}\geq\epsilon^{-1}I$. Regarding the non-positivity statement, fix an origin so that $x$ is a position vector with respect to it and use this origin to represent the action of the translation groups of $\mathbb{M}$}. We know from Proposition 4.10 that there exists $\Psi^{\prime}\in{\mathfrak{S}}_{0}\cap{{\cal H}}^{(0)\perp}$ such that the inequality holds $\langle\Psi^{\prime}|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f]u^{\mu}u^{\nu}_{\Sigma}\Psi^{\prime}\rangle<0$. Defining $\Psi:=U^{-1}_{I,x}\sqrt{H_{u,\epsilon}}\Psi^{\prime}$, we have that the inequality $\langle H_{u,\epsilon}^{-1/2}\Psi|:\thinspace\hat{T}_{\mu\nu}\thinspace:[f_{x}]u^{\mu}u^{\nu}_{\Sigma}H_{u,\epsilon}^{-1/2}\Psi\rangle<-\delta<0$ holds. Since this function of $x$ is continuous, there is a ball $\Delta_{R}$ centered on that $x$ where the function remains bounded above by $\delta/2$. The thesis follows. $\Box$
[…]
Proof of Proposition 6.1. We start with the existence and uniqueness proof and (a). Taking a Minkowskian coordinate system where $\Sigma$ is described by $x^{0}=0$, define $f_{\Delta}(x^{0},\vec{x}):=\int_{\Delta}f^{2}(x^{0},\vec{x}-\vec{y})d^{3}y$ with $f\in\mathscr{D}_{\mathbb{R}}({\mathbb{R}}^{4})$. Since the Borel set $\Delta\subset{\mathbb{R}}^{3}$ is bounded, we have that $f_{\Delta}\in\mathscr{D}_{\mathbb{R}}({\mathbb{R}}^{4})$. At this juncture define ${\mathsf{H}}_{f}^{u}(\Delta)$ directly by (83) so that invariance of ${\mathfrak{S}}_{0}$ is automatic as well as the symmetry property, since $f_{\Delta}$ is real and (b) of Proposition 4.2 holds. Now, (82) immediately follows from (39) by invoking Fubini’s theorem. This concludes the existence part of the proof, the uniqueness being obvious since $\Psi$ in (82) varies in the dense domain ${\mathfrak{S}}_{0}$. The covariance property in (a) immediately follows from the general covariance property of the stress-energy tensor operator.
(b) The found operator, generally speaking, is non-positive as it immediately follows from (e) in Proposition 5.1 choosing $\Delta$ sufficiently narrowed around any given $x$ and renaming $\Psi$ the vector therein replaced by $\frac{1}{\sqrt{H_{\epsilon}^{u}}}\Psi$.
(c) (86) is a consequence of the divergence theorem and the conservation equation in (c) of Proposition 4.11.
(d) First of all notice that the composition $({\mathsf{H}}_{f}^{u}(\Delta)H_{\epsilon}^{u})^{-1/2}$ is well defined on ${\mathfrak{S}}_{0}$ since $({\mathsf{H}}_{f}^{u}(\Delta)H_{\epsilon}^{u})^{-1/2}$ leaves that space invariant. The stated identity is immediate since the two operators in the thesis coincide on ${\mathfrak{S}}_{0}$ by construction. $\Box$
Proof of Lemma 6.3. It is sufficient to apply Eq.(3) in [Few12] referring to the explicit computations presented in Sec.2.5 therein, specialized to the Minkowski spacetime and using as reference state the Poincaré invariant state we indicated by $\Omega$. (The only necessary hypothesis is that the two-point function of the considered normalized state vector $\Psi$ is such that ${\mathbb{M}}\times{\mathbb{M}}\ni(x,y)\mapsto\langle\Psi|\hat{\phi}(x)\hat{\phi}(y)\Psi\rangle-\langle\Omega|\hat{\phi}(x)\hat{\phi}(y)\Omega\rangle\in C^{\infty}({\mathbb{M}}\times{\mathbb{M}})$. For $\Psi\in{\mathfrak{S}}_{0}$ this fact is guaranteed by Proposition 4.4.) In this case the partial differential operator $Q$ is the operator $u^{\mu}u^{\prime\nu}D_{\mu\nu}$ (47) treated as in the proof of Proposition 4.10: $u^{\mu}u^{\prime\nu}D_{\mu\nu}=c^{2}D^{\prime}_{00}-s^{2}D^{\prime}_{11}$
$\frac{c^{2}-s^{2}}{2}(Q^{\prime}_{0}\otimes Q^{\prime}_{0}+Q^{\prime}_{1}\otimes Q^{\prime}_{1})+\frac{c^{2}+s^{2}}{2}(Q^{\prime}_{2}\otimes Q^{\prime}_{2}+Q^{\prime}_{3}\otimes Q^{\prime}_{3}+Q^{\prime}_{4}\otimes Q^{\prime}_{4})$
($c=\cosh\chi$, $s=\sinh\chi$ for some $\chi\geq 0$) so that it is a positive linear combination of products of real first-order (at most) differential operators as requested. In our concrete case $u=\partial_{0}$. With these choices, following the first part of Sec.2.5 in [Few12] we have that, where $k=(\omega,\vec{k})$ with $\omega=\sqrt{\vec{k}^{2}+m^{2}}$,
$\int\langle\Psi|:\thinspace T_{\mu\nu}\thinspace:(x^{0},\vec{x})\Psi\rangle u^{\mu}u^{\prime\nu}h^{\prime}(x^{0})^{2}dx^{0}\geq-\int_{0}^{+\infty}\frac{d\alpha}{\pi}\int_{0}^{+\infty}\frac{d^{3}k}{(2\pi)^{3}}\frac{2\omega(u^{\prime 0}\omega-\vec{k}\cdot\vec{u}^{\prime})}{4\omega}|\hat{h^{\prime}}(\omega+\alpha)|^{2}\>.$
Integration of $\vec{k}\cdot\vec{u}^{\prime}/\omega$ in the angular variables of $d^{3}k=k^{2}\sin\theta dkd\theta d\phi$ (choosing $z$ parallel to $\vec{u}^{\prime}$) gives a vanishing contribution. The remaining integration, following the same route as in the first part of Sec.2.5 in [Few12], furnishes
$\int\langle\Psi|:\thinspace T_{\mu\nu}\thinspace:(x^{0},\vec{x})\Psi\rangle u^{\mu}u^{\prime\nu}h^{\prime}(x^{0})^{2}dx^{0}\geq-u^{\prime 0}\frac{1}{16\pi^{3}}\int_{m}^{+\infty}|\hat{h^{\prime}}(s)|^{2}s^{4}Q_{3}(s)ds\>,$
that is our thesis since $u^{\prime}=(u^{\prime 0},\vec{u}^{\prime})$ with $u^{\prime 0}=\sqrt{1+\vec{u}^{\prime 2}}$. $\Box$
[…]
(b) As is well known, the Friedrichs extension satisfies $A_{F}\geq cI$ if $A\geq cI$ and $c>0$ (see, e.g., [GeSu25]). In this case $\sigma(A_{F})\geq c>0$ by spectral calculus. In particular $Ker(A_{F})=\{0\}$, so that $\overline{\sqrt{A_{F}}(D(A))}=\overline{Ran(\sqrt{A_{F}})}=Ker(\sqrt{A_{F}})^{\perp}=Ker(A_{F})^{\perp}={{\cal H}}$, but also $0\in\rho(A_{F}^{\alpha})$ for $\alpha\geq 0$ and the considered operators are closed since they are selfadjoint, so that $Ran(A_{F}^{\alpha})=\overline{Ran(A_{F}^{\alpha})}=Ker(A_{F}^{\alpha})^{\perp}={{\cal H}}$, in particular, for $\alpha=1/2$ the identity in (b) follows. Again, since $0\in\rho(A_{F}^{\alpha})$ it holds $A_{F}^{-\alpha}\in{\mathfrak{B}}({{\cal H}})$. $\Box$
Proof of Lemma 6.8. Observing that $D(A^{0})=D(A)\subset D(A^{0}_{F})\cap D(A_{F})\subset D(\sqrt{A^{0}_{F}})$, the only item to be proved is the first inequality in (a). Next (b) is a straightforward consequence of (a), defining the operator $1/\sqrt{A^{0}_{F}}$ via functional calculus. (c) follows from Lemma 6.7; in particular the last sentence is a trivial consequence of the previous part. Let us prove the first inequality in (a). Take $c>0$ and define the operators $A^{0}+cI\geq 0$ and $A+cI\geq 0$, and consider the associated forms $a(x):=\langle x|(A+cI)x\rangle$ and $b(x):=\langle x|(A^{0}+cI)x\rangle$. By construction, every Cauchy sequence in the graph norm $||\cdot||_{b}$ is Cauchy in the graph norm $||\cdot||_{a}$. As a consequence, passing to the form completions, which are the forms of respectively $(A+cI)_{F}=A_{F}+cI$ and $(A^{0}+cI)_{F}=A^{0}_{F}+cI$ from known properties of Friedrichs extensions [GeSu25], we have $D(\sqrt{A_{F}^{0}+cI})\subset D(\sqrt{A_{F}+cI})$. It immediately follows from spectral calculus that $x\in D(\sqrt{A_{F}^{0}})$ implies $x\in D(\sqrt{A_{F}^{0}+cI})$. If $x\in D(\sqrt{A_{F}^{0}+cI})$ there is a $||\cdot||_{b}$-Cauchy sequence $D(A)\ni x_{n}\to x$ such that $\langle x_{n}|(A^{0}+cI)x_{n}\rangle\to\langle\sqrt{A_{F}^{0}+cI}x|\sqrt{A_{F}^{0}+cI}x\rangle$. This sequence is also Cauchy for the other norm $||\cdot||_{a}$ and $\langle x_{n}|(A-cI)x_{n}\rangle\to\langle\sqrt{A_{F}+cI}x|\sqrt{A_{F}+cI}x\rangle$. Since $\langle x_{n}|(A^{0}+cI)x_{n}\rangle\geq\langle x_{n}|(A+cI)x_{n}\rangle$ we conclude that $\langle\sqrt{A_{F}^{0}+cI}x|\sqrt{A_{F}^{0}+cI}x\rangle\geq\langle\sqrt{A_{F}+cI}x|\sqrt{A_{F}+cI}x\rangle$. Passing to the spectral representations $\int_{{\mathbb{R}}^{+}}(\lambda+c)\langle x|P^{0}(d\lambda)x\rangle\geq\int_{{\mathbb{R}}^{+}}(\lambda+c)\langle x|P(d\lambda)x\rangle$ is valid for every $c>0$, so that $\int_{{\mathbb{R}}^{+}}\lambda\langle x|P^{0}(d\lambda)x\rangle\geq\int_{{\mathbb{R}}^{+}}\lambda\langle x|P(d\lambda)x\rangle$ which means $\langle\sqrt{A^{0}_{F}}x|\sqrt{A^{0}_{F}}x\rangle\geq\langle\sqrt{A_{F}}x|\sqrt{A_{F}}x\rangle$. By hypothesis this is valid if $x\in D(\sqrt{A_{F}^{0}})$ as wanted. $\Box$